### v1

# Upper Approximation Bounds for Neural Oscillators

Abstract
Neural oscillators, originating from the second-order ordinary differential equations (ODEs), have demonstrated competitive performance in stably learning causal mappings between long-term sequences or continuous temporal functions. However, theoretically quantifying the capacities of their neural network architectures remains a significant challenge. In this study, the neural oscillator consisting of a second-order ODE followed by a multilayer perceptron (MLP) is considered. Its upper approximation bound for approximating causal and uniformly continuous operators between continuous temporal function spaces and that for approximating uniformly asymptotically incrementally stable second-order dynamical systems are derived. The established proof method of the approximation bound for approximating the causal continuous operators can also be directly applied to state-space models consisting of a linear time-continuous complex recurrent neural network followed by an MLP. Theoretical results reveal that the approximation error of the neural oscillator for approximating the second-order dynamical systems scales polynomially with the reciprocals of the widths of two utilized MLPs, thus mitigating the curse of parametric complexity. The decay rates of two established approximation error bounds are validated through two numerical cases. These results provide a robust theoretical foundation for the effective application of the neural oscillator in science and engineering.
keywords:
Neural oscillator , Upper approximation bounds , Causal continuous operator , Stable dynamical system
journal: Neural Networks
[…]
[label1]organization=Institute for Risk and Reliability, Leibniz University Hannover, addressline=Callinstraße 34, city=Hannover, postcode=30167, country=Germany
[…]
[label2]organization=Department of Civil and Environmental Engineering, The Hong Kong Polytechnic University, addressline=Kowloon, city=Hong Kong, country=China
[…]
[label3]organization=Department of Computing and Mathematical Sciences, California Institute of Technology, city=Pasadena, state=California, country=United States
[…]
[label4]organization=Guangdong-Hong Kong Joint Research Laboratory for Marine Infrastructure, The Hong Kong Polytechnic University, addressline=Kowloon, city=Hong Kong, country=China
[…]
[label5]organization=Department of Civil and Environmental Engineering, University of Liverpool, city=Liverpool, postcode=L69 3GH, country=United Kingdom
[…]
[label6]organization=International Joint Research Center for Resilient Infrastructure & International Joint Research Center for Engineering Reliability and Stochastic Mechanics, Tongji University, city=Shanghai, postcode=200092, country=China
1 Introduction
Modeling the mapping relationships of long sequences or continuous temporal functions is a significant challenge in machine learning. A large variety of neural network architectures have been proposed to address sequential and functional mapping tasks, including recurrent neural networks (RNNs) (Rumelhart et al., 1986) and their variants (Hochreiter and Schmidhuber, 1997; Cho et al., 2014), attention-based architectures (Vaswani et al., 2017), state-space (SS) models based on the first-order ordinary differential equations (ODEs) (Gu et al., 2020, 2021, 2022), and neural oscillators based on the second-order ODEs (Rusch and Mishra, 2020, 2021; Lanthaler et al., 2023; Rusch and Rus, 2024; Huang and Xia, 2025). SS models and neural oscillators possess the fast inference capability of RNNs together with the competitive performance of the attention-based architectures in long-range sequence learning (Gu and Dao, 2023; Rusch and Rus, 2024). Furthermore, neural oscillators have rigorously demonstrated to mitigate the problem of vanishing or exploding gradients (Rusch and Mishra, 2020, 2021) and can provide stable dynamics in learning long-range interactions (Rusch and Rus, 2024). A recent overview on SS models and neural oscillators was provided by Tiezzi et al. (2025).
Despite the practical success of SS models and neural oscillators, advancing their theoretical understanding, particularly in quantifying the capacities of their network architectures, remains a significant challenge. For the universal approximation properties of SS models, Li et al. (2022) demonstrated that linear RNNs can universally approximate linear, causal, continuous, regular, and time-homogeneous operators between continuous temporal function spaces along with an approximation bound. Based on the result of Li et al. (2022) and the Volterra series (Boyd et al., 1984), Wang and Xue (2023) proved that given sufficiently large model sizes, an SS model consisting of a time-continuous linear RNN followed by a polynomial nonlinear projection can approximate any time-homogeneous causal continuous operator between continuous temporal function spaces. Similarly, Muca Cirone et al. (2024) demonstrated that a selective SS model comprising a linear controlled differential equation followed by a multilayer perceptron (MLP) can universally approximate causal continuous operators of continuous temporal functions. However, neither result provides a quantitative relationship between the approximation error and the number of parameters required in the SS models. Orvieto et al. (2024) derived an upper approximation bound for an SS model consisting of a time-discrete linear RNN followed by an MLP. However, this result is only limited to the mapping relationships between finite-length sequences and does not apply to infinite-length sequences or continuous temporal functions. In addition to above results, the approximation property of the original RNN for approximating the first-order dynamical systems (Hanson and Raginsky, 2020) and those for approximating fading memory causal operators of time-discrete sequences (Grigoryeva and Ortega, 2018a, b; Gonon et al., 2023) have been studied.
For neural oscillators, a fundamental lemma proposed by Lanthaler et al. (2023) provides a novel insight that the time-varying sine transform coefficients of input continuous temporal functions can be used to construct a universal neural network architecture for the causal continuous operators between continuous temporal function spaces. Based on this fundamental lemma, the neural oscillator architecture consisting of a second-order ODE followed by an MLP was proven to be a universal approximator for such causal continuous operators (Lanthaler et al., 2023; Rusch and Rus, 2024). However, these theoretical results are only qualitative, and the quantitative approximation bounds of neural oscillators remain undeveloped.
In this study, the neural oscillator consisting of a second-order ODE followed by an MLP is considered. Its upper approximation bound for approximating causal and uniformly continuous operators between continuous temporal function spaces and that for approximating uniformly asymptotically incrementally stable second-order dynamical systems are derived. In Section 2, the network architecture of the neural oscillator is briefly introduced, and the newly established theoretical approximation bounds of the neural oscillator are presented informally to facilitate understanding. The formal statements of two theorems that establish the approximation bounds for the neural oscillator, accompanied by the supporting lemmas essential for their proofs, are provided in Section 3. In Section 4, the decay rates of the two established approximation error bounds are validated through two numerical cases. The main contributions of this study are summarized as follows:
  1. 1.
    In Section 3.1, Lemma 1 establishes an error bound for approximating continuous temporal functions using time-varying sine transform coefficients. Theorem 1 derives an upper approximation bound for the neural oscillator in approximating causal and uniformly continuous operators between continuous temporal function spaces. The proof method of Theorem 1 established in this study can also be directly employed to derive the approximation bound for the SS models consisting of a linear time-continuous complex RNN followed by an MLP.
  2. 2.
    In Section 3.2, Theorem 2 establishes an upper approximation bound for the neural oscillator in approximating the uniformly asymptotically incrementally stable second-order dynamical systems. The approximation error scales polynomially with the reciprocals of the widths of two utilized MLPs. This theorem provides a robust theoretical foundation for the effective application of the neural oscillator in science and engineering.
  3. 3.
    During the mathematical derivation in Section 3.2, Lemma 4 enhances the proof of a fundamental Lemma 3.1.2 on differential equation flow by Van Handel (2007). The results in Lemma 6 and D demonstrate that two methods, one based on the differential equation flow and the other one on the Grönwall’s inequality (Ames and Pachpatte, 1997), are equivalent in establishing the bound of the difference between the solutions of two second-order ODEs driven by the same input functions. The proof of Theorem 2 establishes a new method to ensure that all possible solutions of a neural ODE, whose MLP is employed to approximate the derivative function of a target ODE, remain within a domain where the error between the MLP of the neural ODE and the derivative function of the target ODE can be quantified.
Notations: In this study, the $\it{L^{r}}$-norm of an n-dimensional vector $\mathbf{a}=[a_{1},a_{2},...,a_{n}]^{\top},1\leq r\leq+\infty$, is $|\mathbf{a}|_{r}=[\sum_{i=1}^{n}|a_{i}|^{r}]^{1/r}$ for $r<\infty$, or $|\mathbf{a}|_{r}=\max_{1\leq i\leq n}|a_{i}|$ for $r=+\infty$. $|\mathbf{a}|$ is short for $|\mathbf{a}|_{1}$. For a matrix $\mathbf{M}=[m_{ij}]_{p\times q}$, its $L^{r,s}$-norm, $1\leq r,s\leq+\infty$, is $\lvert\mathbf{M}\rvert_{r,s}=\lvert\left[\lvert\mathbf{m}_{1}\rvert_{r},\lvert\mathbf{m}_{2}\rvert_{r},\dots,\lvert\mathbf{m}_{q}\rvert_{r}\right]^{\top}\rvert_{s}$, where $\mathbf{m}_{j}$ is the $j^{\text{th}}$ column vector of the matrix $\mathbf{M}$. The $L^{r}$-norm of $\mathbf{M}$ is $|\mathbf{M}|_{r}=\sup_{|\mathbf{a}|_{r}=1}|\mathbf{Ma}|_{r}/|\mathbf{a}|_{r}$. Re($\cdot$) and Im($\cdot$) represent the real and imaginary parts of a complex number, respectively. $\lfloor\cdot\rfloor$ and $\lceil\cdot\rceil$ represent the floor and ceiling functions, respectively. $\mathbb{Z}$ represents the integer space. $\mathbb{R}$ represents the real number space. $\mathbb{R}^{p{\times}q}$ represents the $p\times q$-dimensional real matrix space. $f^{\prime}(t)=\mathrm{d}f(t)/\mathrm{d}t$, $f^{\prime\prime}(t)=\mathrm{d}^{2}f(t)/\mathrm{d}t^{2}$, and $f^{\prime\prime\prime}(t)=\mathrm{d}^{3}f(t)/\mathrm{d}t^{3}$. $C_{0}([0,T];\mathbb{R}^{p})$ represents a p-dimensional vector-valued real continuous function space defined over $[0,T]$ with respect to $L^{\infty}$-norm ${\lVert\cdot\rVert}_{L^{\infty}}$, e.g., ${\lVert\mathbf{u}(t)\rVert}_{L^{\infty}}$ $\sup_{t\in[0,T]}|\mathbf{u}(t)|$ for $\mathbf{u}(t)=[u_{1}(t),u_{2}(t),...,u_{p}(t)]^{\top}\in C_{0}([0,T];{\mathbb{R}^{p}})$. In addition, $L^{\infty}{[0,t_{0}]}$-norm ${\lVert\cdot\rVert}_{L^{\infty}[0,t_{0}]}$ is ${\lVert\mathbf{u}(t)\rVert}_{L^{\infty}[0,t_{0}]}=\sup_{t\in[0,t_{0}]}|\mathbf{u}(t)|$ and $L^{2}$-norm ${\lVert\cdot\rVert}_{L^{2}}$ is ${\lVert\mathbf{u}(t)\rVert}_{L^{2}}=\sqrt{\int_{0}^{T}|\mathbf{u}(t)|_{2}^{2}\mathrm{d}t}$. Big $O$ notation: $f(n)=O\left[g(n)\right]$ if there exist positive constants $c$ and $n_{0}$ such that $f(n)\leq c\cdot g(n)$ for all $n\geq n_{0}$.
2 Neural Oscillator
Suppose $\mathbf{u}(t)=[u_{1}(t),u_{2}(t),...,u_{p}(t)]^{\top}\in C_{0}([0,T],\mathbb{R}^{p})$, a neural oscillator architecture mapping $\mathbf{u}(t)$ to $\mathbf{y}(t)=[y_{1}(t),y_{2}(t),...,y_{q}(t)]^{\top}$ over $t\in[0,T]$ is
$\left\{\begin{aligned} &\mathbf{x}^{\prime\prime}(t)={\it{\Gamma}}[\mathbf{x}(t),\mathbf{x}^{\prime}(t),\mathbf{u}(t)]\\ &\mathbf{y}(t)={\it{\Pi}}[\mathbf{x}(t),\mathbf{u}(0),t]\end{aligned}\right$. (1)
with $\mathbf{x}(0)=\mathbf{x}^{\prime}(0)=\mathbf{0}$, where ${\it{\Gamma}}(\cdot)$ and ${\it{\Pi}}(\cdot)$ are two MLPs
${\it{\Gamma}}\left[{{\bf{x}}(t),{\bf{x^{\prime}}}(t),{\bf{u}}(t)}\right]={{\bf{W}}_{2}}{\sigma_{\it{\Gamma}}}\left\{{{{\bf{W}}_{1}}{{\left[{{{\bf{x}}^{\top}}(t),{{{\bf{x^{\prime}}}}^{\top}}(t),{{\bf{u}}^{\top}}(t)}\right]}^{\top}}+{{\bf{b}}_{1}}}\right\}+{{\bf{b}}_{2}},$ (2)
${\it{\Pi}}\left[{{\bf{x}}(t),{\bf{u}}(0),t}\right]={{\bf{\tilde{W}}}_{{h_{\it{\Pi}}}}}{\sigma_{\it{\Pi}}}\left({\ldots{\sigma_{\it{\Pi}}}\left\{{{{{\bf{\tilde{W}}}}_{1}}{{\left[{{{\bf{x}}^{\top}}(t),{{\bf{u}}^{\top}}(0),t}\right]}^{\top}}+{{{\bf{\tilde{b}}}}_{1}}}\right\}}\right)+{{\bf{\tilde{b}}}_{{h_{\it{\Pi}}}}},$ (3)
$\mathbf{x}(t)=[x_{1}(t),x_{2}(t),...,x_{r}(t)]^{\top}$ is an intermediate r-dimensional real-valued state function, $\mathbf{W}_{i}$, i = 1 and 2, $\tilde{\mathbf{W}}_{j}$, j = 1, 2,…, $h_{\it{\Pi}}$, are real-valued tunable weight matrices, $\mathbf{b}_{i}$ and $\tilde{\mathbf{b}}_{j}$ are real-valued tunable bias vectors, $h_{\it{\Pi}}$ is the depth of ${\it{\Pi}}(\cdot)$, and $\sigma_{\it{\Gamma}}\left(\cdot\right)$ and $\sigma_{\it{\Pi}}\left(\cdot\right)$ are the activation functions of ${\it{\Gamma}}(\cdot)$ and ${\it{\Pi}}(\cdot)$, respectively. The neural oscillator architecture in Eq. (1) can be denoted as $\mathbf{y}(t)={\it{\Pi}}\circ{\Phi}_{\it{\Gamma}}[\mathbf{u}(\tau)](t)={\it{\Pi}}\left\{{{\Phi}_{\it{\Gamma}}[{\mathbf{u}(\tau)}](t),\mathbf{u}(0),t}\right\}$, where $\circ$ is the composition operator and $\mathbf{x}(t)=\Phi_{\it{\Gamma}}\left[\mathbf{u}(\tau)\right](t)$ represents the casual operator characterized by the second-order ODE in Eq. (1).
When $\sigma_{\it{\Gamma}}(\cdot)$ is the Rectified Linear Unit (ReLU) $\sigma_{\mathrm{ReLU}}(x)=\mathrm{max}(0,x)$ (Nair and Hinton, 2010), Eq. (1) can function as a universal approximator of causal continuous operators by first encoding an input function $\mathbf{u}(t)$ into its time-varying sine transform coefficients through a linear second-order ODE in Eq. (1), and subsequently mapping these coefficients, together with the initial value $\mathbf{u}(0)$ of $\mathbf{u}(t)$ and time t, to an output function, as established in Theorem 1 of this study. This capability is similar to the universal approximation property of neural operators (Kovachki et al., 2023). Moreover, for the second-order dynamical systems in science and engineering (Strogatz, 2024), Eq. (1) can directly approximate these systems using its general ODE with an efficient neural network size, as demonstrated in Theorem 2 of this study.
The initial value u(0) added in ${\it{\Pi}}(\cdot)$ eliminates the need for the assumption of $\mathbf{u}(0)=\mathbf{0}$ in the previous universal approximation theorems for neural oscillators, as presented in Lanthaler et al. (2023) and Rusch and Rus (2024). The influence of time in non-autonomous causal operators is characterized by a term $0.25t^{2}$ in Lanthaler et al. (2023), which arises as the solution to a second-order ODE $f^{\prime\prime}(t)=0.5$. However, this has the possibility of causing unbounded derivatives of the MLP ${\it{\Pi}}(\cdot)$. For example, if ${\it{\Pi}}(t^{2})=g(t)$ and $g^{\prime}(t)$ is continuous and bounded over $t\in[0,T]$ with $g^{\prime}(0)\neq 0$, then $g^{\prime}(t)=2t{\it{\Pi}}^{\prime}(t^{2})$ and ${\it{\Pi}}^{\prime}(x)$ tends to infinity as x approaches zero. The potential unbounded derivatives of ${\it{\Pi}}(\cdot)$ may induce challenges for subsequent theoretical analysis. To address this problem, in this study, a term t is directly added to ${\it{\Pi}}(\cdot)$ to ensure that the derivative of ${\it{\Pi}}(\cdot)$ with respect to t can be bounded.
In the following content of this section, the underlying assumptions and problem setup are outlined. Subsequently, the established theoretical approximation bounds of the neural oscillator in Eq. (1) are presented in an informal manner. The formal statements of the derived theoretical results are presented in Theorems 1 and 2 in Section 3, together with the supporting lemmas required for their proofs.
Assumption 1: $C_{0}([0,T];{\mathbb{R}}^{p})$ with $L^{\infty}$-norm is a Banach space. Its metric is defined as ${\lVert{\mathbf{u}_{1}(t)-\mathbf{u}_{2}(t)}\rVert}_{L^{\infty}}$ with $\mathbf{u}_{1}(t)$ and $\mathbf{u}_{2}(t)$ $\in$ $C_{0}([0,T];\mathbb{R}^{p})$.
Assumption 2: K is a compact subset of $C_{0}([0,T];\mathbb{R}^{p})$} and it is closed. By virtue of the Arzelà-Ascoli theorem (Beattie and Butzmann, 2013), all functions belonging to K are uniformly bounded. There exists a constant $B_{K}$ such that all elements of $K$ are bounded by $B_{K}$, that is ${\lVert{u_{i}(t)}\rVert}_{L^{\infty}}\leq B_{K}$ for all $\mathbf{u}(t)\in K$ and $1\leq i\leq p$, where $u_{i}(t)$ is the $i^{\mathrm{th}}$ element of $\mathbf{u}(t)$. It follows from the Arzelà-Ascoli theorem that the functions belonging to $K$ are equicontinuous, and there exists a continuous modulus of continuity $\mathbf{\upphi}_{K}(t)=[\phi_{K,1}(t),\phi_{K,2}(t),...,\phi_{K,p}(t)]^{\top}$ such that for arbitrary $0\leq t_{1}\leq t_{2}\leq T$ and all $\mathbf{u}(t)\in K$, it satisfies $|u_{i}(t_{1})-u_{i}(t_{2})|\leq\phi_{K,i}(t_{2}-t_{1})$, thereby $|\mathbf{u}(t_{1})-\mathbf{u}(t_{2})|\leq\left|\mathbf{\upphi}_{K}(t_{2}-t_{1})\right|$, where $\phi_{K,i}(t):[0,+\infty)\to[0,+\infty)$ is monotonic with $\phi_{K,i}(t)\to 0$ as $t\to 0$.
Assumption 3: ${{\Phi}}=[\Phi_{1},\Phi_{2}....,\Phi_{q}]:C_{0}([0,T];{\mathbb{R}}^{p})\to C_{0}([0,t];{\mathbb{R}}^{q})$ is a causal and uniformly continuous operator. For $\forall{t}\in[0,T]$, $\Phi[\mathbf{u}(\tau)](t)$ only depends on $\mathbf{u}(\tau)$ over $\tau\in[0,t]$, where $\mathbf{u}(t)\in C_{0}([0,T];{\mathbb{R}}^{p})$. For $\forall{\varepsilon}>0$, there exists a distance $d_{\varepsilon}>0$, such that ${\lVert{\mathbf{u}_{1}(t)-\mathbf{u}_{2}(t)}\rVert}_{L^{\infty}}<d_{\varepsilon}\Rightarrow{\lVert{\Phi[\mathbf{u}_{1}(\tau)](t)-\Phi[\mathbf{u}_{2}(\tau)](t)}\rVert}_{L^{\infty}}<\varepsilon$ for arbitrary $\mathbf{u}_{1}(t)$ and $\mathbf{u}_{2}(t)$ $\in C_{0}([0,t];{\mathbb{R}}^{p})$.
Assumption 4: Since ${{\Phi}}=[\Phi_{1},\Phi_{2}....,\Phi_{q}]$ is a causal and uniformly continuous operator mapping from $C_{0}([0,T];{\mathbb{R}}^{p})$ to $C_{0}([0,T];{\mathbb{R}}^{q})$ and $K$ is a compact subset of $C_{0}([0,t];{\mathbb{R}}^{p})$, the image set $\Phi(K)$ of $K$ by $\Phi$ is a compact subset of $C_{0}([0,T];{\mathbb{R}}^{q})$. There exists a constant $B_{\Phi(K)}$ such that all elements of $\Phi(K)$ are bounded by $B_{\Phi(K)}$, that is ${\lVert{v_{j}(t)}\rVert}_{L^{\infty}}\leq B_{\Phi(K)}$ for all $\mathbf{v}(t)=[v_{1}(t),v_{2}(t),...,v_{q}(t)]^{\top}\in\Phi(K)$ and $1\leq j\leq q$.
Assumption 5: For the causal and uniformly continuous operator ${{\Phi}}=[\Phi_{1},\Phi_{2}....,\Phi_{q}]$ mapping from $C_{0}([0,T];{\mathbb{R}}^{p})$ to $C_{0}([0,T];{\mathbb{R}}^{q})$, for each $\Phi_{j}$, it is assumed that the first-order derivatives of all temporal functions in the image set $\Phi_{j}\left\{C_{0}([0,T];{\mathbb{R}^{p}})\right\}$ of $C_{0}([0,T];{\mathbb{R}^{p}})$ by $\Phi_{j}$ are bounded by a finite constant $D_{\Phi_{j}}$.
Assumption 6: Since ${{\Phi}}=[\Phi_{1},\Phi_{2}....,\Phi_{q}]$ is a causal and uniformly continuous operator mapping from $C_{0}([0,T];{\mathbb{R}}^{p})$ to $C_{0}([0,T];{\mathbb{R}}^{q})$, its $j^{\mathrm{th}}$ element $\Phi_{j}$ has a monotonic continuous modulus of continuity $\phi_{\Phi_{j}}(\cdot):[0,+\infty)\to[0,+\infty)$ with $\phi_{\Phi_{j}}(r)\to 0$ as $r\to 0$, such that ${\left\lVert\Phi_{j}[\mathbf{u}_{1}(\tau)](t)-\Phi_{j}[\mathbf{u}_{2}(\tau)](t)\right\rVert}_{L^{\infty}}$ $\leq\phi_{\Phi_{j}}[{\left\lVert\mathbf{u}_{1}(t)-\mathbf{u}_{2}(t)\right\rVert}_{L^{\infty}}]$ for arbitrary $\mathbf{u}_{1}(t)$ and $\mathbf{u}_{2}(t)$ $\in C_{0}([0,T];{\mathbb{R}^{p}})$. $\phi_{\Phi}(\cdot)$ $\sum_{j=1}^{q}{\phi_{\Phi_{j}}(\cdot)}:[0,+\infty)\to[0,+\infty)$ with $\phi_{\Phi}(r)\to 0$ as $r\to 0$, is a monotonic modulus of continuity of $\Phi$, such that ${\left\lVert\Phi[\mathbf{u}_{1}(\tau)](t)-\Phi[\mathbf{u}_{2}(\tau)](t)\right\rVert}_{L^{\infty}}\leq\sum_{j=1}^{q}{\left\lVert\Phi_{j}[\mathbf{u}_{1}(\tau)](t)-\Phi_{j}[\mathbf{u}_{2}(\tau)](t)\right\rVert}_{L^{\infty}}\leq\phi_{\Phi}[{\left\lVert\mathbf{u}_{1}(t)-\mathbf{u}_{2}(t)\right\rVert}_{L^{\infty}}]$ for arbitrary $\mathbf{u}_{1}(t)$ and $\mathbf{u}_{2}(t)$ $\in C_{0}([0,T];{\mathbb{R}^{p}})$.
Setting 1: Given two real continuous function spaces $C_{0}([0,T];{\mathbb{R}^{p}})$ and $C_{0}([0,T];{\mathbb{R}^{q}})$ with $L^{\infty}$-norm ${\left\lVert\cdot\right\rVert}_{L^{\infty}}$, a compact set $K\subset C_{0}([0,T];{\mathbb{R}^{p}})$, and a causal and uniformly continuous operator $\Phi:C_{0}([0,T];{\mathbb{R}^{p}})\to C_{0}([0,T];{\mathbb{R}^{q}})$, for each $\mathbf{u}(t)\in K$, its corresponding image function $\mathbf{v}(t)\in C_{0}([0,T];{\mathbb{R}^{q}})$ is calculated by $\Phi$
${\bf{v}}(t)=\Phi\left[{{\bf{u}}(\tau)}\right](t)={\left\{{{\Phi_{1}}\left[{{\bf{u}}(\tau)}\right](t),{\Phi_{2}}\left[{{\bf{u}}(\tau)}\right](t),\ldots,{\Phi_{q}}\left[{{\bf{u}}(\tau)}\right](t)}\right\}^{\top}}.$ (4)
Informal Statement of Theorem 1: Given a compact subset $K\subset C_{0}([0,T];{\mathbb{R}^{p}})$ and a causal and uniformly continuous operator $\Phi:C_{0}([0,T];{\mathbb{R}^{p}})$ $\to C_{0}([0,T];{\mathbb{R}^{q}})$ satisfying Assumptions 1 to 6 with $\mathrm{\upphi}_{K}(t)=L_{K}[t,t,\dots,t]^{\top}$ and $L_{K}>0$, for every integer $M_{\it{\Gamma}}$ larger than an independent deterministic threshold and every positive integer $H_{\it{\Pi}}$, there exist two MLPs $\it{\Gamma}(\cdot)$ and $\it{\Pi}(\cdot)$ employing $\sigma_{\mathrm{ReLU}}(\cdot)$, such that for an arbitrary $\mathbf{u}(t)\in K$, the corresponding solution $\mathbf{y}(t)=[y_{1}(t),y_{2}(t),...,y_{q}(t)]^{\top}$ from Eq. (1) subject to $\mathbf{u}(t)$ with the initial conditions $\mathbf{x}(0)={\mathbf{x}}^{\prime}(0)=\mathbf{0}$ satisfies
$\left|{\Phi\left[{{\bf{u}}(\tau)}\right](t)-{\bf{y}}(t)}\right|\leq{\varepsilon_{\bf{y}}}$ (5)
for all $t\in[0,T]$, where
[…]
and ${\phi_{\Phi}}(\cdot)$ is the modulus of continuity of $\Phi$ in Assumption 6. $\it{\Gamma}(\cdot)$ has one hidden layer and the widths of its input, hidden, and output layers are $w_{{\it{\Gamma},{\mathrm{in}}}}=O\left(pM_{{\it{\Gamma}}}\right)$, $w_{\it{\Gamma}}=O\left(pM_{{\it{\Gamma}}}\right)$, and $w_{{\it{\Gamma},{\mathrm{out}}}}=O\left(pM_{{\it{\Gamma}}}\right)$, respectively. For ${\it{\Pi}}(\cdot)$, the widths of its input, hidden, and output layers are $w_{{\it{\Pi},{\mathrm{in}}}}=O\left(pM_{{\it{\Gamma}}}\right)$, $w_{\it{\Pi}}=O\left(qpM_{{\it{\Gamma}}}\right)$, and $w_{{\it{\Pi},{\mathrm{out}}}}=q$, respectively, and its depth is $h_{\it{\Pi}}=H_{\it{\Pi}}+1$.
Informal Statement of Theorem 2: Let K be a compact subset of $C_{0}([0,T];{\mathbb{R}^{p}})$, a second-order dynamical system
$\left\{\begin{aligned} &{\hat{\bf x}}^{\prime\prime}(t)=g\left[{{\hat{\bf x}}(t),{\hat{\bf x}}^{\prime}(t),{\bf{u}}(t)}\right]\\ &{\hat{\bf y}}(t)=h\left[{{\hat{\bf x}}(t)}\right]\end{aligned}\right$., (7)
with ${\hat{\mathbf{x}}}(0)={{\hat{\mathbf{x}}}}^{\prime}(0)=\mathbf{0}$, is uniformly asymptotically incrementally stable for all $\mathbf{u}(t)\in K$ on a domain $\left[-B_{\mathbf{x}},B_{\mathbf{x}}\right]^{r}\times\left[-B_{{\mathbf{x}}^{\prime}},B_{{\mathbf{x}}^{\prime}}\right]^{r}\times\left[-B_{K},B_{K}\right]^{p}\subset\mathbb{R}^{2r+p}$, where ${\hat{\mathbf{x}}}(t)=\left[{\hat{x}}_{1}(t),{\hat{x}}_{2}(t),...,{\hat{x}}_{r}(t)\right]^{\top}$, ${\hat{\mathbf{y}}}(t)=\left[{\hat{y}}_{1}(t),{\hat{y}}_{2}(t),...,{\hat{y}}_{q}(t)\right]^{\top}$, $B_{{\mathbf{x}}}=\alpha B_{\beta{g}}+B_{\hat{\mathbf{x}}}$, $B_{{\mathbf{x}}^{\prime}}=\alpha B_{\beta{g}}+B_{\hat{\mathbf{x}}^{\prime}}$, $\alpha$ is a positive coefficient, $B_{\hat{\mathbf{x}}}$ and $B_{\hat{\mathbf{x}}^{\prime}}$ are the bounds of ${\hat{\mathbf{x}}}(t)$ and ${\hat{\mathbf{x}}}^{\prime}(t)$, respectively, $B_{K}$ is the bound of all $\mathbf{u}(t)\in K$ in Assumption 2, and $B_{\beta{g}}$ is the stability bound of the system in Eq. (7). The functions $g(\cdot)=\left[g_{1}(\cdot),g_{2}(\cdot),...g_{r}(\cdot)\right]^{\top}$ and $h(\cdot)=\left[h_{1}(\cdot),h_{2}(\cdot),...h_{q}(\cdot)\right]^{\top}$ are Lipschitz continuous over ${\mathbb{R}}^{2r+p}$ and $\mathbb{R}^{r}$, respectively, with arbitrary finite-order derivatives. Then, for arbitrary positive errors $\varepsilon_{1}\leq\alpha/r$ and $\varepsilon_{2}$, there exist two one-hidden-layer MLPs ${\it{\Gamma}}(\cdot)$ and ${\it{\Pi}}(\cdot)$ employing $\sigma_{\mathrm{ReLU}}(\cdot)$, such that for all $\mathbf{u}(t)\in K$, the corresponding solutions $\mathbf{y}(t)$ to Eq. (1) and ${\hat{\mathbf{y}}}(t)$ to Eq. (7) satisfy
$\left|{{\bf{y}}(t)-{\hat{\bf y}}(t)}\right|\leq{L_{h}}{B_{{\beta_{g}}}}\sqrt{r}{\varepsilon_{1}}+q{\varepsilon_{2}}$ (8)
for all $t\in[0,T]$, where $L_{h}$ is the Lipschitz constant of $h(\cdot)$. For ${\it{\Gamma}}(\cdot)$, the widths of its input, hidden, and output layers are $w_{{\it{\Gamma},{\mathrm{in}}}}=2r+p$, $w_{\it{\Gamma}}=O\left(r\varepsilon_{1}^{-2}\right)$, and $w_{{\it{\Gamma},{\mathrm{out}}}}=r$, respectively. For ${\it{\Pi}}(\cdot)$, the widths of its input, hidden, and output layers are $w_{{\it{\Pi},{\mathrm{in}}}}=r+p+1$, $w_{\it{\Pi}}=O\left(q\varepsilon_{2}^{-2}\right)$, and $w_{{\it{\Pi},{\mathrm{out}}}}=q$, respectively.
3 Upper Approximation Bounds of the Neural Oscillator
In this section, an upper approximation bound of the neural oscillator in Eq. (1) for approximating causal and uniformly continuous operators between continuous temporal function spaces and that for approximating uniformly asymptotically incrementally stable second-order dynamical systems are established in Subsections 3.1 and 3.2, respectively.
3.1 Causal Continuous Operators
In this subsection, Lemma 1 derives an error bound for approximating all $\mathbf{u}(t)\in K$ using their initial values $\mathbf{u}(0)$ and a finite number of time-varying sine transform coefficients, combined with a set of corresponding deterministic base functions. Lemma 2 proves the existence of an MLP ${\it{\Gamma}}(\cdot)$ with $\sigma_{\mathrm{ReLU}}(\cdot)$, such that the solution $\mathbf{x}(t)$ to the first ODE in Eq. (1) corresponds to the sine transform coefficient vector of the input function $\mathbf{u}(t)$. Lemma 3 demonstrates that the target causal continuous operator $\Phi:C_{0}([0,T];{\mathbb{R}^{p}})\to C_{0}([0,T];{\mathbb{R}^{q}})$ in Setting 1 can be approximated by a continuous function $\Psi\left[\mathbf{x}(t),\mathbf{u}(0),t\right]$, which depends on time $t$, $\mathbf{u}(0)$, and the solution $\mathbf{x}(t)$, and establishes an associated error bound. Lemma 4 determines the Lipschitz constants, linear moduli of continuity, and a compact argument domain of $\Psi\left[\mathbf{x}(t),\mathbf{u}(0),t\right]$. Finally, by approximating $\Psi\left[\mathbf{x}(t),\mathbf{u}(0),t\right]$ with an MLP ${\it{\Pi}}(\cdot)$ in Eq. (3) and combing its approximation error with the error bound from Lemma 3, Theorem 1 establishes an approximation bound for the neural oscillator in Eq. (1). The proofs for all lemmas and Theorem 1 in this subsection are presented in A.
Lemma 1: Given a compact subset $K\subset C_{0}([0,T];{\mathbb{R}^{p}})$ satisfying Assumptions 1 and 2, for an arbitrary $v$ satisfying $0<v<T$ and every positive integer M larger than one, there exist M frequencies $\omega_{1},\omega_{2},...,\omega_{M}\in\mathbb{R}$, M phases $\theta_{1},\theta_{2},...,\theta_{M}\in\mathbb{R}$, and M weights $\alpha_{1},\alpha_{2},...,\alpha_{M}\in\mathbb{R}$, such that
$\mathop{\sup}\limits_{\tau\in[0,t]}\left|{{\bf{u}}(t-\tau)-\left({{\bf{u}}(0)+\sum\limits_{n=1}^{M}{{\alpha_{n}}\left\{{{\mathcal{L}_{t}}{\bf{u}}({\omega_{n}})+\frac{{{\bf{u}}(0)}}{{{\omega_{n}}}}\left[{\cos({\omega_{n}}t)-1}\right]}\right\}\sin({\omega_{n}}\tau-{\theta_{n}})}}\right)}\right|\leq{\varepsilon_{K}}$ (9)
for all $t\in[0,T]$ and all $\mathbf{u}(t)\in K$, where
${\mathcal{L}_{t}}{\bf{u}}(\omega)=\int_{0}^{t}{{\bf{u}}(t-\tau)\sin(\omega\tau){\rm{d}}\tau},$ (10)
${\varepsilon_{K}}=\left|{\mathbf{\upphi}_{K}(v)}\right|+\frac{{8T^{2}\left|{\mathbf{\upphi}_{K}(T)}\right|{c_{\rho,2}}}}{{{\pi^{2}}{v^{2}}M}},$ (11)
$c_{\rho,2}$ is an independent constant, and $\mathbf{\upphi}_{K}(t)$ is a modulus of continuity for all $\mathbf{u}(t)\in K$ in Assumption 2.
Remark 1: Lemma 1 is an enhanced version of Lemma B.1 in Lanthaler et al. (2023) by providing a precise error bound and incorporating the influence of the initial value $\mathbf{u}(0)$. The initial value $\mathbf{u}(0)$ in the original Lemma B.1 is assumed to be zero. To address this limitation, a warm-up phase was introduced in Lanthaler et al. (2023) to linearly extend $\mathbf{u}(t)$ from $t=0$ to a negative time instant $t=-t_{0}$, $t_{0}>0$. However, this approach may introduce some practical considerations in selecting an appropriate value for $t_{0}$.
Remark 2: In Lemma 1, in addition to the positive integer M, there exists an additional free parameter v arising from the discontinuities of $\tilde{\mathbf{f}}_{t}(\tau)$ in Eq. (39) in A.1 and its derivatives at $\tau=0$ and $\tau=\pm t$. Since $\mathbf{\upphi}_{K}(t)$ increases monotonically, v can be eliminated by minimizing $\varepsilon_{K}$. For example, if $\mathbf{\upphi}_{K}(t)=[L_{K}t,L_{K}t,...,L_{K}t]^{\top}$, where $L_{K}$ is a positive Lipschitz constant of all $\mathbf{u}(t)\in K$, then, $\varepsilon_{K}$ is calculated as
${\varepsilon_{K}}=p{L_{K}}v+\frac{{8{T^{3}}p{L_{K}}{c_{\rho,2}}}}{{{\pi^{2}}{v^{2}}M}}.$ (12)
The minimum of $\varepsilon_{K}$ in Eq. (12) is
${\varepsilon_{K}}=p{L_{K}}2T{\left({\frac{{{2}{c_{\rho,2}}}}{{{\pi^{2}}M}}}\right)^{{1\mathord{\left/{\vphantom{13}}\right.\kern-1.2pt}3}}}+\frac{{8{T^{3}}p{L_{K}}{c_{\rho,2}}}}{{{\pi^{2}}M}}\frac{1}{4T^{2}}{\left({\frac{{{\pi^{2}}M}}{{{2}{c_{\rho,2}}}}}\right)^{{2\mathord{\left/{\vphantom{23}}\right.\kern-1.2pt}3}}}=\frac{{54^{{1\mathord{\left/{\vphantom{13}}\right.\kern-1.2pt}3}}p{L_{K}}Tc_{\rho,2}^{{1\mathord{\left/{\vphantom{13}}\right.\kern-1.2pt}3}}}}{{{\pi^{{2\mathord{\left/{\vphantom{23}}\right.\kern-1.2pt}3}}}{M^{{1\mathord{\left/{\vphantom{13}}\right.\kern-1.2pt}3}}}}},$ (13)
which is obtained by
$\frac{{{\rm{d}}{\varepsilon_{K}}}}{{{\rm{d}}v}}=p{L_{K}}-\frac{{16{T^{3}}p{L_{K}}{c_{\rho,2}}}}{{{\pi^{2}}{v^{3}}M}}=0\Rightarrow v=2T{\left({\frac{{{2}{c_{\rho,2}}}}{{{\pi^{2}}M}}}\right)^{{1\mathord{\left/{\vphantom{13}}\right.\kern-1.2pt}3}}}.$ (14)
Thus, for each positive integer M satisfying
$M>16{c_{\rho,2}}{\pi^{-2}},$ (15)
the approximation error $\varepsilon_{K}$ in Eq. (9) can be expressed in Eq. (13). The condition in Eq. (15) ensures that the underlying parameter $v$ is smaller than $T$.
Lemma 2: Given a compact subset K $\subset C_{0}([0,T];{\mathbb{R}^{p}})$ satisfying Assumptions 1 and 2, for every positive integer M larger than one and arbitrary M frequencies $\omega_{1}$, $\omega_{2}$,…, $\omega_{M}$ $\in\mathbb{R}$, there exists a one-hidden-layer MLP ${\it{\Gamma}}(\cdot)$ with $\sigma_{\mathrm{ReLU}}(\cdot)$, the widths of its input, hidden, and output layers are $w_{{\it{\Gamma},{\mathrm{in}}}}=p(2M+1)$, $w_{\it{\Gamma}}=2pM$, and $w_{{\it{\Gamma},{\mathrm{out}}}}=pM$, respectively, such that for all $t\in[0,T]$ and all $\mathbf{u}(t)\in K$, the solution $\mathbf{x}(t)=[\mathbf{x}_{1}^{\top}(t),\mathbf{x}_{2}^{\top}(t),...,\mathbf{x}_{M}^{\top}(t)]^{\top}$ from the second-order ODE in Eq. (1) subject to $\mathbf{u}(t)$ with the initial conditions $\mathbf{x}(0)=\mathbf{x}^{\prime}(0)=\mathbf{0}$ satisfies $\mathbf{x}(t)=[\mathcal{L}_{t}\mathbf{u}^{\top}(\omega_{1}),\mathcal{L}_{t}\mathbf{u}^{\top}(\omega_{2}),...,\mathcal{L}_{t}\mathbf{u}^{\top}(\omega_{M})]^{\top}$, where $\mathcal{L}_{t}\mathbf{u}(\omega)$ is defined in Eq. (10) in Lemma 1.
Lemma 3: Given a compact subset $K\subset C_{0}([0,T];{\mathbb{R}^{p}})$ and a causal and uniformly continuous operator $\Phi:C_{0}([0,T];{\mathbb{R}^{p}})\to C_{0}([0,T];{\mathbb{R}^{q}})$ satisfying Assumptions 1, 2, 3, and 6, for an arbitrary positive $v<T$ and every positive integer M larger than one, there exists a continuous mapping $\Psi=[\psi_{1},\psi_{2},...,\psi_{q}]:{\mathbb{R}}^{p(M+1)}\times{[0,T]}\to{\mathbb{R}}^{q}$ and a one-hidden-layer MLP ${\it{\Gamma}}(\cdot)$ specified in Lemma 2, such that for all $\mathbf{u}(t)\in K$, the solution $\mathbf{x}(t)=[\mathbf{x}_{1}^{\top}(t),\mathbf{x}_{2}^{\top}(t),...,\mathbf{x}_{M}^{\top}(t)]^{\top}$ from the second-order ODE in Eq. (1) subject to $\mathbf{u}(t)$ with the initial conditions $\mathbf{x}(0)=\mathbf{x}^{\prime}(0)=\mathbf{0}$ satisfies
$\left|{\Phi\left[{{\bf{u}}(\tau)}\right](t)-\Psi({\bf{x}},{{\bf{u}}_{0}},t)}\right|\leq{\phi_{\Phi}}({\varepsilon_{K}})$ (16)
for all $t\in[0,T]$, where $\mathbf{x}=[\mathbf{x}_{1}^{\top},\mathbf{x}_{2}^{\top},...,\mathbf{x}_{M}^{\top}]^{\top}$, $\mathbf{x}_{n}$ represents the value of $\mathbf{x}_{n}(t)$, $\mathbf{u}_{0}$ represents the value of $\mathbf{u}(0)$, $\phi_{\Phi}(\cdot)$ is a modulus of continuity of $\Phi$ in Assumption 6, and $\varepsilon_{K}$} is defined in Eq. (11) in Lemma 1.
Remark 3: Lemma 3 is an enhanced version of the novel fundamental lemma in Lanthaler et al. (2023) by providing a precise error bound and incorporating the influence of the initial value $\mathbf{u}(0)$, while adding one constraint that the target operator $\Phi$ is uniformly continuous. This constraint can ensure that the error term $\phi_{\Phi}(\varepsilon_{K})$ in Eq. (16) is independent of $\mathbf{u}(t)$. Under the additional condition $\mathbf{\upphi}_{K}(t)=[L_{K}t,L_{K}t,...,L_{K}t]^{\top}$, $\varepsilon_{K}$ in Eq. (16) can be calculated by Eq. (13) for $M$ satisfying Eq. (15).
Remark 4: Even if two additional conditions that $\mathbf{u}(0)$ = $\mathbf{0}$ for all $\mathbf{u}(t)\in K$ and $\Phi$ being time-autonomous are imposed, the term t of $\Psi\left(\mathbf{x},{\mathbf{0}},t\right)$ in Lemma 3 cannot be ignored because $\sin\left[\omega_{n}\left(t-\tau\right)-\theta_{n}\right]$ in Eq. (54) in Appendix A.3 contains the time $t$. The detailed explanation is provided in C.
Lemma 4: Under the conditions in Lemma 3 and Assumption 5, along with $\mathrm{\upphi}_{K}(t)=[L_{K}t,L_{K}t,...,L_{K}t]^{\top}$, $v$ satisfying Eq. (14), and an integer $M$ satisfying Eq. (15), the values of all arguments of $\psi_{j}\left(\mathbf{x},\mathbf{u}_{0},t\right)$ in Lemma 3 are within a compact domain $\Omega_{M}\left(TB_{K},B_{K},T\right)=\left[-TB_{K},TB_{K}\right]^{pM}\times\left[-B_{K},B_{K}\right]^{p}\times\left[0,T\right]$ for all $\mathbf{u}(t)\in K$ and j=1,2,,qj=1,2,...,q, the Lipschtiz constant $L_{{\psi_{j}}}\left(M\right)$ of $\psi_{j}\left(\mathbf{x},\mathbf{u}_{0},t\right)$ with respect to its each argument is
$\begin{array}[]{l}{L_{{\psi_{j}}}\left(M\right)}=\max\left\{{{D_{{\Phi_{j}}}}+p{B_{K}}MT^{-1}c_{L}{{\phi^{\prime}}_{{\Phi_{j}}}}(0),\max\left[{\left({1+2M}\right),T^{-1}}\right]{{\phi^{\prime}}_{{\Phi_{j}}}}(0)}\right\},\end{array}$ (17)
and a linear modulus of continuity $\phi_{\psi_{j},M}(x)$ of $\psi_{j}\left(\mathbf{x},\mathbf{u}_{0},t\right)$ under $L^{2}$-norm is
${\phi_{{\psi_{j},M}}}(x)=\sqrt{p(M+1)+1}{L_{{\psi_{j}}}\left(M\right)}x,$ (18)
where $\phi_{\Phi_{j}}(\cdot)$ is the monotonic modulus of continuity of $\Phi_{j}$ in Assumption 6, ${\phi}^{\prime}_{\Phi_{j}}(0)$ is the first-order derivative of $\phi_{\Phi_{j}}(\cdot)$ at zero, $B_{K}$ is the bound of all $\mathbf{u}(t)\in K$ in Assumption 2, $D_{\Phi_{j}}$ is the derivative bound of the image set $\Phi_{j}\left\{C_{0}([0,T];{\mathbb{R}^{p}})\right\}$ in Assumption 5, and $c_{L}$ is an independent constant.
Remark 5: Lemma 4 establishes a compact domain $\Omega_{M}\left(TB_{K},B_{K},T\right)$ for the arguments of $\psi_{j}\left(\mathbf{x},\mathbf{u}_{0},t\right)$ that is valid for all $\mathbf{u}(t)\in K$, every integer $M$ satisfying Eq. (15), and $v$ satisfying Eq. (14), along with the linear modulus of continuity $\phi_{\psi_{j},M}(x)$ in Eq. (18) of $\psi_{j}\left(\mathbf{x},\mathbf{u}_{0},t\right)$. Consequently, $\psi_{j}\left(\mathbf{x},\mathbf{u}_{0},t\right)$ over $\Omega_{M}\left(TB_{K},B_{K},T\right)$ can be approximated using an MLP, and existing MLP approximation bounds can be applied to quantify the approximation error. By combining a utilized MLP approximation error with the term ${\phi_{\Phi}}({\varepsilon_{K}})$ from Eq. (16) in Lemma 3, an approximation bound of the neural oscillator in Eq. (1), for approximating the causal and uniformly continuous operators between continuous temporal function spaces, is derived and presented in following Theorem 1.
Theorem 1: Given a compact subset $K\subset C_{0}([0,T];{\mathbb{R}^{p}})$ and a causal and uniformly continuous operator $\Phi:C_{0}([0,T];{\mathbb{R}^{p}})\to C_{0}([0,T];{\mathbb{R}^{q}})$ satisfying Assumptions 1 to 6 with $\mathrm{\upphi}_{K}(t)=[L_{K}t,L_{K}t,...,L_{K}t]^{\top}$, for every integer $M_{\it{\Gamma}}$ satisfying Eq. (15) along with $v$ satisfying Eq. (14), and every positive integer $H_{\it{\Pi}}$, there exist two MLPs $\it{\Gamma}(\cdot)$ and $\it{\Pi}(\cdot)$ employing $\sigma_{\mathrm{ReLU}}(\cdot)$, such that for every $\mathbf{u}(t)\in K$, the corresponding solution $\mathbf{y}(t)=[y_{1}(t),y_{2}(t),...,y_{q}(t)]^{\top}$ from Eq. (1) subject to $\mathbf{u}(t)$ with the initial conditions $\mathbf{x}(0)={\mathbf{x}}^{\prime}(0)=\mathbf{0}$ satisfies
$\left|{\Phi\left[{{\bf{u}}(\tau)}\right](t)-{\bf{y}}(t)}\right|\leq{\varepsilon_{\bf{y}}}$ (19)
for all $t\in[0,T]$, where
[…]
$B_{\Omega}=2\max\left(TB_{K},B_{K},T\right)$, $B_{K}$ is the bound of all $\mathbf{u}(t)\in K$ in Assumption 2, $L_{{\psi_{j}}}\left(\cdot\right)$} is in Eq. (17), ${\phi_{\Phi}}(\cdot)$ is the monotonic modulus of continuity of $\Phi$ in Assumption 6, and $c_{\rho,2}$ is the independent constant in Lemma 1. $\it{\Gamma}(\cdot)$ has one hidden layer and the widths of its input, hidden, and output layers are $w_{{\it{\Gamma},{\mathrm{in}}}}$, $w_{\it{\Gamma}}$, and $w_{{\it{\Gamma},{\mathrm{out}}}}$, respectively
$\left\{\begin{aligned} &w_{\it{\Gamma},{\rm{in}}}=p\left(2M_{\it{\Gamma}}+1\right)\\ &w_{\it{\Gamma}}=2pM_{\it{\Gamma}}\\ &{w_{\it{\Gamma},{\rm{out}}}}=pM_{\it{\Gamma}}\end{aligned}.\right$. (21)
For ${\it{\Pi}}(\cdot)$, the widths of its input, hidden, and output layers are $w_{{\it{\Pi},{\mathrm{in}}}}$, $w_{\it{\Pi}}$, and $w_{{\it{\Pi},{\mathrm{out}}}}$, respectively
$\left\{\begin{aligned} &{w_{{\it{\Pi}},{\rm{in}}}}=p\left(M_{\it{\Gamma}}+1\right)+1\\ &w_{\it{\Pi}}=q\left[p\left(M_{\it{\Gamma}}+1\right)+4\right]\\ &{w_{{\it{\Pi}},{\rm{out}}}}=q\end{aligned},\right$. (22)
and its depth is $h_{\it{\Pi}}=H_{\it{\Pi}}+1$.
Remark 6: Beyond the theoretical upper approximation bound for the neural oscillator, Lemma 1, Lemma 3, Lemma 4, and Theorem 1 establish a proof method for deriving the upper approximation bounds of the causal operator network architectures consisting of a linear transformation of input functions followed by an MLP. Particularly, Lemma 4 provides a method for characterizing the continuous mapping $\Psi\left({\bf{x}},{\bf{u}}_{0},t\right)$ in Lemma 3. If the parameters of linear time-continuous RNNs are complex (Ran-Milo et al., 2024), the output functions of the linear complex RNNs can be expressed as the convolution of the trigonometric basis functions with the input functions. Consequently, the proof method established in this study can be directly employed to derive the approximation bound for the complex SS models consisting of a linear time-continuous complex RNN followed by an MLP.
Remark 7: An upper approximation bound of ReLU MLPs by Yarotsky (2017) has been employed in Theorem 3 of Liu et al. (2024) to establish an upper generalization bound of neural operators. The result by Yarotsky (2017) is not employed in this study because its approximation bound constant depends on the argument dimension of a target function and this dependence is implicit. This implicit relationship hinders the establishment of the explicit relationship between the final approximation error $\varepsilon_{\bf{y}}$} in Eq. (19) and $M_{\it{\Gamma}}$ controlling the width of the hidden layer of the MLP ${\it{\Gamma}}(\cdot)$.
Remark 8: Theorem 1 indicates that the approximation error of the neural oscillator in Eq. (1) consists of two parts. The first one is from approximating $\mathbf{u}(t)$ using the sine transform and the second one is from approximating the continuous mapping $\Psi\left(\mathbf{x},\mathbf{u}_{0},t\right)$ in Lemma 3 using the MLP ${\it{\Pi}}(\cdot)$. Eq. (20) indicates that the decay rate $1/\left[p\left(M_{\it{\Gamma}}+1\right)+1\right]$ in the second error polynomially diminishes as the width $w_{\it{\Gamma}}=2pM_{\it{\Gamma}}$ of the hidden layer of the first MLP ${\it{\Gamma}}(\cdot)$ increases. Thus, the neural oscillator in Eq. (1) for approximating the causal continuous operators between continuous temporal function spaces also faces the curse of parametric complexity, which is a common problem in neural operators (Lanthaler, 2023; Lanthaler and Stuart, 2025). A recent result on the quantitative estimates of the number of parameters required in neural operators, presented in Proposition 1 of Kratsios et al. (2024), is similar to the second term on the right side of Eq. (20).
One method to mitigate the curse of parametric complexity is applying other super-expressive activation functions (Maiorov and Pinkus, 1999; Shen et al., 2021; Yarotsky, 2021; Zhang et al., 2022). For instance, Theorem 4 in Maiorov and Pinkus (1999) demonstrates the existence of an activation function that enables a two-hidden-layer neural network architecture to approximate continuous functions with arbitrary accuracy, even when the widths of the hidden layers are held constant. Theorem 1.1 and Corollary 1.2 in Shen et al. (2021) provide a three-hidden-layer MLP architecture, called Floor-Exponential-Step (FLES) network, utilizing the floor function $\lfloor x\rfloor$, the exponential function $2^{x}$, and the step function as the activation function in each neuron. FLES is capable of approximating arbitrary continuous functions with the number of parameters scaling logarithmically with the reciprocal of the approximation error. Theorem 1 in Zhang et al. (2022) introduces a neural network architecture that employs an activation function combining a softsign activation function with a periodic triangular-wave function. This neural network architecture with a width $36d(2d+1)$ and a depth 11 can approximate any continuous function on a $d$-dimensional hypercube within an arbitrarily small error. Super-expressive activation functions have been utilized to mitigate the curse of parametric complexity in neural operators (Schwab et al., 2023). Since the explicit expressions for the Lipschitz constant $L_{\it{\psi}_{j}}(M)$ in Eq. (17) and the linear modulus of continuity $\phi_{\it{\psi}_{j},M}(x)$ in Eq. (18) of the continuous mapping $\psi_{j}\left(\mathbf{x},\mathbf{u}_{0},t\right)$, along with its valid compact argument domain $\Omega_{M}\left(TB_{K},B_{K},T\right)$, have been derived, the super-expressive activation functions mentioned above can be directly applied to the MLP ${\it{\Pi}}(\cdot)$ to mitigate the curse of parametric complexity in the neural oscillator. On the other hand, Lanthaler (2024) has demonstrated that decreasing the required number of parameters in neural operators using super-expressive neural networks may come at the expense of requiring larger model parameter values.
3.2 Second-Order Dynamical Systems
In this subsection, an upper approximation bound of the neural oscillator in Eq. (1) for approximating uniformly asymptotically incrementally stable second-order dynamical systems is derived. First, the asymptotically incrementally stable second-order dynamical system is briefly introduced. Subsequently, Lemma 5 and Lemma 6 establish the relationships between the solution error of two ODEs subject to the same input functions and the error of the derivative functions of the two ODEs. Finally, by utilizing the result in Lemma 6 and the error bound for the MLP approximation of Barron class functions (Barron, 1993), the upper approximation bound of the neural oscillator is demonstrated in Theorem 2. This subsection builds on the methodology of Hanson and Raginsky (2020) and offers technical contributions through detailed mathematical derivations, as elaborated in the accompanying remarks. The proofs for all lemmas and Theorem 2 in this subsection are presented in B.
Two second-order ODEs ${\mathbf{x}}_{i}^{\prime\prime}(t)=g_{i}\left[{\mathbf{x}}_{i}(t),{\mathbf{x}}_{i}^{\prime}(t),\mathbf{u}(t)\right]$, where ${\mathbf{x}}_{i}(t)=[x_{i,1}(t)$ $x_{i,2}(t),\dots,x_{i,r}(t)]^{\top},$ i = 1 and 2, can be rewritten in the state function form
$\left\{\begin{aligned} &{{{\bf{z}}}_{i,1}^{\prime}}(t)={{\bf{z}}_{i,2}}(t)\\ &{{{\bf{z}}}_{i,2}^{\prime}}(t)={g_{i}}\left[{{{\bf{z}}_{i,1}}(t),{{\bf{z}}_{i,2}}(t),{\bf{u}}(t)}\right]\end{aligned}\right$., (23)
where ${\mathbf{z}}_{i}(t)=\left[{\mathbf{z}}_{i,1}^{\top}(t),{\mathbf{z}}_{i,2}^{\top}(t)\right]^{\top}=\left[{\mathbf{x}}_{i}^{\top}(t),{\mathbf{x}}_{i}^{\prime\top}(t)\right]^{\top}$.
Definition 1 (Hanson and Raginsky, 2020): The dynamical system in the form of Eq. (23) is uniformly asymptotically incrementally stable for the inputs in a subset $K\subset C_{0}([0,T];{\mathbb{R}^{p}})$ on a domain $\Omega_{\mathbf{z}}\subset\mathbb{R}^{2r}$ if there exists a set of non-negative stability functions $\beta_{k}(h,t)$ over $h\geq 0$ and $t\geq 0$, k=1,2,,2rk=1,2,...,2r, where $\beta_{k}(0,t)=0$$\lim\limits_{t\to+\infty}\beta_{k}(h,t)=0$, $\beta_{k}(h,t)$ is continuous and strictly increasing with respect to h and is continuous and strictly decreasing with respect to t, such that
$\left|{{z_{k}}(t,{{\bf{z}}_{\tau}})-{z_{k}}(t,{{{\hat{\bf z}}}_{\tau}})}\right|\leq{\beta_{k}}\left({{{\left|{{{\bf{z}}_{\tau}}-{{{\hat{\bf z}}}_{\tau}}}\right|}_{2}},t-\tau}\right)$ (24)
holds for all $\mathbf{u}(t)\in K$, all ${\mathbf{z}}_{\tau}$ and ${\hat{\bf z}}_{\tau}$ $\in\Omega_{\mathbf{z}}$, all $t\in[0,T]$, and all $0\leq\tau\leq t$, where $z_{k}\left(t,{\mathbf{z}}_{\tau}\right)$ is the $k^{\text{th}}$ element of $\mathbf{z}\left(t,{\mathbf{z}}_{\tau}\right)$ and $\mathbf{z}\left(t,{\mathbf{z}}_{\tau}\right)$ represents the solution $\mathbf{z}(t)$ at t driven by $\mathbf{u}(t)$ with an initial condition ${\mathbf{z}}_{\tau}$ at an initial time instant $\tau$, $0\leq\tau\leq t$. The corresponding stability bound $B_{\beta}$ is calculated as
${B_{\beta}}=\int_{0}^{+\infty}{{{\left.{\frac{\partial}{{\partial h}}\sum\limits_{k=1}^{2r}{{\beta_{k}}(h,\tau)}}\right|}_{h={0^{+}}}}{\rm{d}}\tau}.$ (25)
Lemma 5: Let K be a compact subset of $C_{0}([0,T];{\mathbb{R}^{p}})$, $g_{1}(\cdot)$ and $g_{2}(\cdot)$ in Eq. (23) be Lipschitz continuous over $\mathbb{R}^{2r+p}$, and $\Omega_{\mathbf{u}}\subset\mathbb{R}^{p}$ and $\Omega_{\mathbf{z}}\subset\mathbb{R}^{2r}$ be two compact domains. Assume that the values of all $\mathbf{u}(t)\in K$ are in $\Omega_{\mathbf{u}}$ and the values of all possible solutions to Eq. (23), governed by $g_{1}(\cdot)$ or $g_{2}(\cdot)$, driven by all $\mathbf{u}(t)\in K$, with a specified initial condition ${\mathbf{z}}_{0}\in\Omega_{\mathbf{z}}$, remain within $\Omega_{\mathbf{z}}$. Then, for the dynamical systems governed by $g_{1}(\cdot)$ or $g_{2}(\cdot)$ in Eq. (23), their corresponding solutions ${\mathbf{z}}_{1}(t)=[{\mathbf{z}}_{1,1}^{\top}(t),{\mathbf{z}}_{1,2}^{\top}(t)]^{\top}$ and ${\mathbf{z}}_{2}(t)=[{\mathbf{z}}_{2,1}^{\top}(t),{\mathbf{z}}_{2,2}^{\top}(t)]^{\top}$, driven by the same arbitrary $\mathbf{u}(t)\in K$ with the same initial conditions $\mathbf{z}_{1}(0)={\mathbf{z}}_{2}(0)={\mathbf{z}}_{0}$}, satisfy
$\begin{array}[]{l}{{\bf{z}}_{2}}(t,{{\bf{z}}_{0}})-{{\bf{z}}_{1}}(t,{{\bf{z}}_{0}})=\int_{0}^{t}{\frac{{\partial{{\bf{z}}_{1}}\left[{t,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}})}}\left\{{{g_{2}}\left[{{{\bf{z}}_{2,1}}(\tau,{{\bf{z}}_{0}}),{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}}),{\bf{u}}(\tau)}\right]-{g_{1}}\left[{{{\bf{z}}_{2,1}}(\tau,{{\bf{z}}_{0}}),{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}}),{\bf{u}}(\tau)}\right]}\right\}{\rm{d}}\tau}\end{array}$ (26)
for all $t\in[0,T]$, where ${\mathbf{z}}_{i}\left[t,{\mathbf{z}}_{j}\left({\hat{\tau}},{\mathbf{z}}_{\tau}\right)\right]$ represents the solution ${\mathbf{z}}_{i}(t)$ at t driven by $\mathbf{u}(t)$ with the initial condition ${\mathbf{z}}_{j}\left({\hat{\tau}},{\mathbf{z}}_{\tau}\right)$ at $^{\hat{\tau}}$ with $0\leq\tau\leq{\hat{\tau}}\leq t$ and $i,j=$ 1 and 2.
Remark 9: Being analogous to the fundamental Lemma 3.1.2 in Van Handel (2007), which has been employed by Hanson and Raginsky (2020) to derive the approximation bound of RNNs for modeling uniformly asymptotically incrementally stable first-order dynamical systems, Lemma 5 in this study provides a more rigorous proof. In the proof of the step involving Eqs. (81) and (82) in B.1, the proof of the original lemma in Van Handel (2007) utilizes ${\mathbf{z}}_{1}(\tau)={\mathbf{z}}_{1}\left(\tau,{\mathbf{z}}_{0}\right)$ but not ${\mathbf{z}}_{1}\left({\hat{\tau}},{\mathbf{z}}_{\tau}\right)={\mathbf{z}}_{\tau}$ with $\hat{\tau}=\tau$ and replaces an arbitrary ${\mathbf{z}}_{1}\left(\tau,{\mathbf{z}}_{0}\right)$ with its corresponding ${\mathbf{z}}_{2}\left(\tau,{\mathbf{z}}_{0}\right)$ under the same $\mathbf{u}(t)$ to prove Eq. (82). However, this proof method is not rigorous. For an arbitrary time instant $\tau$ ($0\leq\tau\leq T$), the range of all possible values of ${\mathbf{z}}_{1}\left(\tau,{\mathbf{z}}_{0}\right)$ depends on the differential equation governed by $g_{1}(\cdot)$ and all $\mathbf{u}(t)\in K$. It is evident that an arbitrary value within the range of all possible values of ${\mathbf{z}}_{1}\left(\tau,{\mathbf{z}}_{0}\right)$ satisfies Eq. (81). However, it remains uncertain whether all possible values of ${\mathbf{z}}_{2}\left(\tau,{\mathbf{z}}_{0}\right)$ at the same time instant $\tau$ driven by $\mathbf{u}(t)\in K$ lie within the range of the values of ${\mathbf{z}}_{1}\left(\tau,{\mathbf{z}}_{0}\right)$. If not, it is unclear whether ${\mathbf{z}}_{2}\left(\tau,{\mathbf{z}}_{0}\right)$ satisfies Eq. (81) for all $0\leq\tau\leq T$ and all $\mathbf{u}(t)\in K$. In this study, employing an arbitrary $\tau$ and an arbitrary ${\mathbf{z}}_{1}\left({\hat{\tau}},{\mathbf{z}}_{\tau}\right)={\mathbf{z}}_{\tau}$ with $\hat{\tau}=\tau$ in Eq. (81) resolves this uncertainty. ${\mathbf{z}}_{1}\left({\hat{\tau}},{\mathbf{z}}_{\tau}\right)$ is the solution at $^{\hat{\tau}}$ caused by the initial condition ${\mathbf{z}}_{\tau}$ at $\tau$. Being an initial condition, the value of ${\mathbf{z}}_{\tau}$ can be an arbitrary vector within $\Omega_{\mathbf{z}}$. ${\mathbf{z}}_{1}\left({\hat{\tau}},{\mathbf{z}}_{\tau}\right)={\mathbf{z}}_{\tau}$ with $\hat{\tau}=\tau$ indicates that the value of ${\mathbf{z}}_{1}\left({\hat{\tau}},{\mathbf{z}}_{\tau}\right)$ can also be an arbitrary vector within $\Omega_{\mathbf{z}}$. This implies that all vectors within $\Omega_{\mathbf{z}}$ satisfy Eq. (81). From the assumption of Lemma 5, the possible values of ${\mathbf{z}}_{2}\left(\tau,{\mathbf{z}}_{0}\right)$ fall within $\Omega_{\mathbf{z}}$ for all $\tau\in[0,T]$ and all $\mathbf{u}(t)\in K$. Consequently, all possible values of ${\mathbf{z}}_{2}\left(\tau,{\mathbf{z}}_{0}\right)$ satisfy Eq. (81).
Lemma 6: Under the conditions in Lemma 5, the difference between solutions ${\mathbf{z}}_{1}(t)$ and ${\mathbf{z}}_{2}(t)$ can be bounded by
$\begin{array}[]{l}\left|{{{\bf{z}}_{2}}(t,{{\bf{z}}_{0}})-{{\bf{z}}_{1}}(t,{{\bf{z}}_{0}})}\right|\leq{e^{T{L_{{g_{1}}}}}}T\mathop{\max}\limits_{\tau\in[0,T]}\left|{{g_{2}}\left[{{{\bf{z}}_{2,1}}(\tau,{{\bf{z}}_{0}}),{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}}),{\bf{u}}(\tau)}\right]-{g_{1}}\left[{{{\bf{z}}_{2,1}}(\tau,{{\bf{z}}_{0}}),{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}}),{\bf{u}}(\tau)}\right]}\right|\end{array}$ (27)
for all $t\in[0,T]$ and all $\mathbf{u}(t)\in K$, where $L_{g_{1}}=\max\left[L_{g_{1},1},\left(L_{g_{1},2}+1\right)\right]$, $L_{g_{1},1}$ and $L_{g_{1},2}$ are respectively the two Lipschitz constants of $g_{1}(\cdot)$ with respect to its first and second arguments. Furthermore, if the dynamical system governed by $g_{1}(\cdot)$ in Eq. (23) is uniformly asymptotically incrementally stable for all $\mathbf{u}(t)\in K$ on the domain $\Omega_{\mathbf{z}}$ in Lemma 5 with a stability bound $B_{\beta}$ in Eq. (26), then, $\left|{\mathbf{z}}_{2}\left(\tau,{\mathbf{z}}_{0}\right)-{\mathbf{z}}_{1}\left(\tau,{\mathbf{z}}_{0}\right)\right|$ can be bounded by
$\begin{array}[]{l}\left|{{{\bf{z}}_{2}}(t,{{\bf{z}}_{0}})-{{\bf{z}}_{1}}(t,{{\bf{z}}_{0}})}\right|\leq\mathop{\max}\limits_{\tau\in[0,T]}{\left|{{g_{2}}\left[{{{\bf{z}}_{2,1}}(\tau,{{\bf{z}}_{0}}),{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}}),{\bf{u}}(\tau)}\right]-{g_{1}}\left[{{{\bf{z}}_{2,1}}(\tau,{{\bf{z}}_{0}}),{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}}),{\bf{u}}(\tau)}\right]}\right|_{2}}{B_{\beta}}\end{array}$ (28)
for all $\mathbf{u}(t)\in K$ and all $t\in[0,T]$.
Remark 10: As shown in D, the result in Eq. (27) can also be proven using the method based on the Grönwall’s inequality (Ames and Pachpatte, 1997), which was frequently employed in deriving approximation bounds of RNNs for modeling first-order dynamical systems (Funahashi and Nakamura, 1993; Chow and Li, 2000). The proof of Lemma 6 shows that the method based on Lemma 5, which enhances the proof of Lemma 3.1.2 in Van Handel (2007), can simultaneously obtain both categories of the bounds in Eq. (27) and Eq. (28) for the difference between the solutions of two different second-order ODEs subject to the same input functions. This advancement enhances the theoretical understanding of the second-order ODEs. Lemma 6 will be utilized in the subsequent Theorem 2 to establish the approximation bound of the neural oscillator for modeling uniformly asymptotically incrementally stable second-order dynamical systems.
Theorem 2: Let K be a compact subset of $C_{0}([0,T];{\mathbb{R}^{p}})$ and a causal continuous operator $\Phi:C_{0}([0,T];{\mathbb{R}^{p}})\to C_{0}([0,T];{\mathbb{R}^{q}})$ be a second-order dynamical system represented by
$\left\{\begin{aligned} &{\hat{\bf x}}^{\prime\prime}(t)=g\left[{{\hat{\bf x}}(t),{\hat{\bf x}}^{\prime}(t),{\bf{u}}(t)}\right]\\ &{\hat{\bf y}}(t)=h\left[{{\hat{\bf x}}(t)}\right]\end{aligned}\right$. (29)
with the initial conditions ${\hat{\mathbf{x}}}(0)={{\hat{\mathbf{x}}}}^{\prime}(0)=\mathbf{0}$, where ${\hat{\mathbf{x}}}(t)=[{\hat{x}}_{1}(t),{\hat{x}}_{2}(t),\dots,$ ${\hat{x}}_{r}(t)]^{\top}$ and ${\hat{\mathbf{y}}}(t)=\left[{\hat{y}}_{1}(t),{\hat{y}}_{2}(t),\dots,{\hat{y}}_{q}(t)\right]^{\top}$. The functions $g(\cdot)=[g_{1}(\cdot),g_{2}(\cdot),$ $\dots,g_{r}(\cdot)]^{\top}$ and $h(\cdot)=\left[h_{1}(\cdot),h_{2}(\cdot),...h_{q}(\cdot)\right]^{\top}$ are Lipschitz continuous over ${\mathbb{R}}^{2r+p}$ and $\mathbb{R}^{r}$, respectively, with arbitrary finite-order derivatives. Assume that the dynamical system in Eq. (29) is uniformly asymptotically incrementally stable for all $\mathbf{u}(t)\in K$ on a domain $\left[-B_{\mathbf{x}},B_{\mathbf{x}}\right]^{r}\times\left[-B_{{\mathbf{x}}^{\prime}},B_{{\mathbf{x}}^{\prime}}\right]^{r}\times\left[-B_{K},B_{K}\right]^{p}\subset\mathbb{R}^{2r+p}$, where $B_{{\mathbf{x}}}=\alpha B_{\beta{g}}+B_{\hat{\mathbf{x}}}$, $B_{{\mathbf{x}}^{\prime}}=\alpha B_{\beta{g}}+B_{\hat{\mathbf{x}}^{\prime}}$, $\alpha$ is a positive coefficient, $B_{\hat{\mathbf{x}}}$ and $B_{\hat{\mathbf{x}}^{\prime}}$ are the bounds of ${\hat{\mathbf{x}}}(t)$ and ${\hat{\mathbf{x}}}^{\prime}(t)$, respectively, $B_{K}$ is the bound of all $\mathbf{u}(t)\in K$ in Assumption 2, and $B_{\beta{g}}$ is the stability bound of the dynamical system in Eq. (25), then, for arbitrary positive errors $\varepsilon_{1}\leq\alpha/r$ and $\varepsilon_{2}$, there exist two one-hidden-layer MLPs ${\it{\Gamma}}(\cdot)$ and ${\it{\Pi}}(\cdot)$ employing $\sigma_{\mathrm{ReLU}}(\cdot)$, such that for all $\mathbf{u}(t)\in K$, the corresponding solutions $\mathbf{y}(t)$ to Eq. (1) and ${\hat{\mathbf{y}}}(t)$ to Eq. (29) satisfy
$\left|{{\bf{y}}(t)-{\hat{\bf y}}(t)}\right|\leq{L_{h}}{B_{{\beta_{g}}}}\sqrt{r}{\varepsilon_{1}}+q{\varepsilon_{2}}$ (30)
for all $t\in[0,T]$, where $L_{h}$ is the Lipschitz constant of $h(\cdot)$ in Eq. (29). For ${\it{\Gamma}}(\cdot)$, the widths of its input, hidden, and output layers are $w_{{\it{\Gamma},{\mathrm{in}}}}$, $w_{\it{\Gamma}}$, and $w_{{\it{\Gamma},{\mathrm{out}}}}$, respectively,
$\left\{\begin{aligned} &{w_{{\it{\Gamma}},{\rm{in}}}}=2r+p\\ &{w_{\it{\Gamma}}}=8r\left\lceil{{C_{\it{\Gamma}}}\varepsilon_{1}^{-2}}\right\rceil\\ &{w_{{\it{\Gamma}},{\rm{out}}}}=r\end{aligned}\right$., (31)
where $C_{\it{\Gamma}}$ is a constant only depending on $g(\cdot)$, $\sigma_{\mathrm{ReLU}}(\cdot)$, $r$, $p$, $B_{\mathbf{x}}$, $B_{{\mathbf{x}}^{\prime}}$, and $B_{K}$. For ${\it{\Pi}}(\cdot)$, the widths of its input, hidden, and output layers are $w_{{\it{\Pi},{\mathrm{in}}}}$, $w_{\it{\Pi}}$, and $w_{{\it{\Pi},{\mathrm{out}}}}$, respectively,
$\left\{\begin{aligned} &{w_{{\it{\Pi}},{\rm{in}}}}=r+p+1\\ &{w_{\it{\Pi}}}=8q\left\lceil{{C_{\it{\Pi}}}\varepsilon_{2}^{-2}}\right\rceil\\ &{w_{{\it{\Pi}},{\rm{out}}}}=q\end{aligned}\right$., (32)
where $C_{\it{\Pi}}$ is a constant only depending on $h(\cdot)$, $\sigma_{\mathrm{ReLU}}(\cdot)$, $r$, and $B_{\mathbf{x}}$
Remark 11: Theorem 2 demonstrates that when the target causal continuous operator can be explicitly represented as a uniformly asymptotically incrementally stable second-order dynamical system, the widths of the hidden layers in the two MLPs utilized in the neural oscillator scale polynomially with the reciprocals of the approximation errors. This implies that the curse of parametric complexity of the neural oscillator in this situation is mitigated. An analogous approximation bound for RNNs modeling the first-order dynamical systems has been obtained in Hanson and Raginsky (2020).
Remark 12: Approximating an ODE system by a neural ODE involves approximating the derivative function of the target ODE using an MLP. The approximation bound for an MLP approximating the derivative function is established over a compact domain of the derivative function’s arguments and the MLP’s arguments. The derivative function’s arguments are input functions and the solutions of the target ODE. The MLP’s arguments are input functions and the solutions of the neural ODE. Therefore, in establishing an approximation bound for a neural ODE approximating an ODE system, where the derivative function of the target ODE is approximated by the MLP of the neural ODE, one critical issue is ensuring that the values of all possible solutions from the neural ODE, driven by a set of considered input functions, can lie within the compact domain, where the error between the MLP of the neural ODE and the target derivative function can be quantified. To meet this requirement, Lemma 7 in Hanson and Raginsky (2020) proposes to approximate the derivative function of a target ODE using a compactly supported function, such that the values of the solutions from the resulting neural ODE, whose derivative function is modeled by this compactly supported function, can be bounded in a compact domain. In this study, the relationships between the solution error of two ODEs subject to the same input functions and the error of the derivative functions of the two ODEs are established in Eqs. (27) and (28) of Lemma 6. Based on the result of Eq. (28), the error between the MLP ${\it{\Gamma}}(\cdot)$ and the target derivative function $g(\cdot)$ in Eq. (29) is quantified over the compact domain $\Omega\left(B_{\mathbf{x}}^{r},B_{{\mathbf{x}}^{\prime}}^{r},B_{K}^{p}\right)=\left[-B_{\mathbf{x}},B_{\mathbf{x}}\right]^{r}\times\left[-B_{{\mathbf{x}}^{\prime}},B_{{\mathbf{x}}^{\prime}}\right]^{r}\times\left[-B_{K},B_{K}\right]^{p}$. And all possible values of the solutions $\mathbf{x}(t)$ and ${\mathbf{x}}^{\prime}(t)$ to the second-order ODE in Eq. (1) with ${\it{\Gamma}}(\cdot)$ driven by all $\mathbf{u}\in K$ are within the domain $\Omega\left(B_{\mathbf{x}}^{r},B_{{\mathbf{x}}^{\prime}}^{r},B_{K}^{p}\right)$, as highlighted in Eq. (99), Eq. (100), and Eq. (106) in the proof of Theorem 2 in B.3.
4 Numerical Study
In this section, a Bouc-Wen nonlinear system driven by stochastic excitations is employed to validate the results in Theorem 1 and Theorem 2. A numerical study in Huang and Xia (2025) has shown that the neural oscillator can effectively learn the mapping relationship between the input excitation and the output response of a Bouc-Wen system exhibiting plastic deformation. The differential equation of a 5-degree of freedom Bouc-Wen system is
$\left\{\begin{aligned} &\mathbf{M}\ddot{\mathbf{X}}(t)+\mathbf{C}\dot{\mathbf{X}}(t)+\lambda{\mathbf{K}\mathbf{X}}(t)+(1-\lambda){\tilde{\mathbf{K}}\mathbf{Z}}(t)=-{\mathbf{M}}{U_{\text{e}}}(t)\\ &{{\dot{Z}}_{i}}(t)={\dot{\tilde{X}}}_{i}(t)-\beta\left|{\dot{\tilde{X}}_{i}}(t)\right|{\left|{Z_{i}}(t)\right|^{s-1}}{Z_{i}}(t)-\gamma{\dot{\tilde{X}}_{i}}(t){\left|{{Z_{i}}(t)}\right|^{s}},\;\;i=1,2,\ldots,5\end{aligned}\right$., (33)
where ${\mathbf{X}}=\left[X_{1}(t),X_{2}(t),\dots,X_{5}(t)\right]^{\top}$, $\tilde{\bf{X}}(t)=[X_{1}(t),X_{2}(t)-X_{1}(t),$ $X_{3}(t)-X_{2}(t),\dots,X_{5}(t)-X_{4}(t)]^{\top}$, ${\mathbf{Z}}=\left[Z_{1}(t),Z_{2}(t),\dots,Z_{5}(t)\right]^{\top}$, ${\mathbf{X}}(0)={\mathbf{X}}^{\prime}(0)={\mathbf{Z}}(0)={\mathbf{0}}$,
${\mathbf{M}}=\begin{bmatrix}m&0&0&0&0\\ 0&m&0&0&0\\ 0&0&m&0&0\\ 0&0&0&m&0\\ 0&0&0&0&m\end{bmatrix},\;\;\;{\mathbf{K}}=\begin{bmatrix}2k&-k&0&0&0\\ -k&2k&-k&0&0\\ 0&-k&2k&-k&0\\ 0&0&-k&2k&-k\\ 0&0&0&-k&k\end{bmatrix},\;\;\;{\tilde{{\mathbf{K}}}}=\begin{bmatrix}k&-k&0&0&0\\ 0&k&-k&0&0\\ 0&0&k&-k&0\\ 0&0&0&k&-k\\ 0&0&0&0&k\end{bmatrix}$,\;\;\; (34)
$m=1382.4$ kilogram, $k=1.7\times 10^{6}$ Newton/meter, $\mathbf{C}$ is determined using $\mathbf{M}$ and $\mathbf{K}$ such that the damping ratios of all response modes are 0.05, $\lambda=0.01$, $\beta=2$, $\gamma=2$, and $s=3$. $U_{\text{e}}(t)$ in Eq. (33) represents different stochastic excitations in the following two cases. The ODE system in Eq. (33) is solved numerically using a fourth-order Runge-Kutta method (Griffiths and Higham, 2010) with a time discretization of $\Delta t=0.01$ seconds over the time interval $[0,10]$. For the numerical implementation of training, the neural oscillator is discretized using a second-order Runge–Kutta scheme (Huang and Xia, 2025). The approach for training the neural oscillator with input-output pair samples is briefly described in E. All code for this numerical study is available at: https://github.com/ZifengH22/Upper-Approximation-Bounds-for-Neural-Oscillators.
4.1 Case 1
In this case, $U_{\text{e}}(t)$ in Eq. (33) is a stochastic excitation
$U_{\text{e}}(t)=c_{1}\sin(0.4{\pi}t)+c_{2}\cos(0.8{\pi}t)+c_{3}\sin(1.2{\pi}t)+c_{4}\cos(1.6{\pi}t)+c_{5}\sin(2{\pi}t),$ (35)
where $c_{j},j=1,2,\ldots,5$ are are independent random variables uniformly distributed on the interval $\left[-35,35\right]$. $5\times 10^{4}$ independent discrete-time samples $U_{\text{e},l}(t_{i})$ of $U_{\text{e}}(t)$, where $i=0,1,2,\dots,999$ and $l=1,2,\dots,5\times 10^{4}$, are simulated. Their induced $5\times 10^{4}$ samples $E_{X_{5},l}(t_{i})$ of the extreme value process $E_{X_{5}}(t)=\max_{\tau\in\left[0,t\right]}\left|X_{5}(\tau)\right|$ are computed by numerically solving Eq. (33). The mapping from $U_{\text{e}}(t)$ to $E_{X_{5}}(t)$ is learned using the neural oscillator with the simulated samples $U_{\text{e},l}(t_{i})$ and $E_{X_{5},l}(t_{i})$. The decay rate in Theorem 1 is validated using this case.
In this case, five neural oscillators ${\it{\Pi}}_{i}\circ{\Phi}_{{\it{\Gamma}}_{i}}\left(\cdot\right)$, $i=1,2,\ldots,5$, are considered. The sizes of the five one-hidden-layer MLPs ${{\it{\Gamma}}}_{i}(\cdot)$ are the same. Their widths of input, hidden, and output layers are 11, 10, and 5, respectively. The activation functions of the five MLPs ${\it{\Gamma}}_{i}(\cdot)$ are $\sigma_{\mathrm{ReLU}}(\cdot)$. The five MLPs ${{\it{\Pi}}}_{i}(\cdot)$ corresponding to the five neural oscillators share the same input, hidden, and output layer widths of 7, 10, and 1, respectively. The numbers of hidden layers in ${{\it{\Pi}}}_{1}(\cdot)$, ${{\it{\Pi}}}_{2}(\cdot)$, ${{\it{\Pi}}}_{3}(\cdot)$, ${{\it{\Pi}}}_{4}(\cdot)$, and ${{\it{\Pi}}}_{5}(\cdot)$ are 1, 2, 3, 4, and 5, respectively. The activation functions of the five MLPs ${{\it{\Pi}}}_{i}(\cdot)$ are the Parametric Rectified Linear Unit (PReLU) $\sigma_{\text{PReLU}}(x)=\max(0,x)+\alpha\min(0,x)$, where $\alpha$ is a parameter learned during training (He et al., 2015). Since $\sigma_{\mathrm{ReLU}}(\cdot)$ is a special case of $\sigma_{\text{PReLU}}(\cdot)$, the result in Theorem 1 also holds for the neural oscillators ${\it{\Pi}}_{i}\circ{\Phi}_{{\it{\Gamma}}_{i}}\left(\cdot\right)$ when ${{\it{\Pi}}}_{i}(\cdot)$ employs $\sigma_{\text{PReLU}}(\cdot)$.
The $5\times 10^{4}$ samples $E_{X_{5},l}(t_{i})$ of output $E_{X_{5}}(t)$ are arranged in a descending order with respect to their maximum values $\hat{E}_{X_{5},l}=\max_{t_{i}\in\left[0,10\right]}E_{X_{5},l}(t_{i})$. Then, 10000 samples $E_{X_{5},l}(t_{i})$, whose maximum values $\hat{E}_{X_{5},l}$ are approximately equally spaced from the largest one to the smallest one of the $5\times 10^{4}$ $\hat{E}_{X_{5},l}$, along with their corresponding input samples $U_{\text{e},l}(t_{i})$, are selected. Among these 10000 sample pairs, the $5^{\text{th}}$, $10^{\text{th}}$, $15^{\text{th}}$$\dots,$ $10000^{\text{th}}$ pairs (total 2000 pairs) are selected as a validation dataset, and the remaining 8000 sample pairs constitute a training dataset. As this study is concerned solely with validating the approximation error bound rather than investigating the generalization performance, this sampling strategy is appropriate to build the training data. Each of the five neural oscillators is trained using the training approach outlined in E with the 8000 sample pairs. The 8-th power loss function $\ell_{8}$ in Eq. (118) is employed to ensure that the neural oscillator adequately learns the target mapping with respect to the $L^{\infty}$-norm. The training phase comprises 10000 epochs, with a batch size of 2000 for each epoch. The initial learn rate is 0.005. The learning rate drop period is 100 epochs. The learning rate drop factor is 0.9772. To mitigate the risk of exploding gradients during training, the gradient norm clipping strategy (Pascanu et al., 2013) with a threshold of 1 is employed for the five MLPs ${\it{\Gamma}}_{i}(\cdot)$. For the five MLPs ${{\it{\Pi}}}_{i}(\cdot)$, their norm clipping thresholds are $1,\sqrt{203/92},\sqrt{314/92},\sqrt{425/92},\sqrt{536/92}$, respectively, where $92,203,314,425,$ and $536$ are the numbers of learnable parameters in ${{\it{\Pi}}}_{1}(\cdot)$, ${{\it{\Pi}}}_{2}(\cdot)$, ${{\it{\Pi}}}_{3}(\cdot)$, ${{\it{\Pi}}}_{4}(\cdot)$, and ${{\it{\Pi}}}_{5}(\cdot)$. Batch normalization (Ioffe and Szegedy, 2015) is applied to the five MLPs ${{\it{\Pi}}}_{i}(\cdot)$ to facilitate training due to the depth of their hidden layers. Since batch normalization only adds an affine transformation to the input of each hidden layer of the five MLPs ${{\it{\Pi}}}_{i}(\cdot)$, the result of Theorem 1 still holds for ${{\it{\Pi}}}_{i}(\cdot)$ with batch normalization. For the first neural oscillator with ${\it{\Gamma}}_{1}(\cdot)$ and ${{\it{\Pi}}}_{1}(\cdot)$, the initial values of the learnable network parameters are randomly selected. For the second neural oscillator with ${\it{\Gamma}}_{2}(\cdot)$ and ${{\it{\Pi}}}_{2}(\cdot)$, the initial parameter values of ${\it{\Gamma}}_{2}(\cdot)$ are taken directly from the parameters of ${\it{\Gamma}}_{1}(\cdot)$ after the last update. The initial parameter values for the first hidden layer and output layer of ${{\it{\Pi}}}_{2}(\cdot)$ are taken directly from the parameters of the corresponding layers in ${{\it{\Pi}}}_{1}(\cdot)$ after the last update. The second hidden layer of ${{\it{\Pi}}}_{2}(\cdot)$ is initialized as an identity mapping. The MLPs in the remaining three neural oscillators are initialized in the same manner. For each neural oscillator, the neural network achieving the smallest validation loss under the $L^{\infty}$-norm is selected as the final model.
For each trained neural oscillator, its $5\times 10^{4}$ prediction samples $\tilde{E}_{X_{5},l}(t_{i})$ driven by the input samples $U_{\text{e},l}(t_{i})$ are computed. The relative approximation errors under $L^{\infty}$-norm and $L^{2}$-norm, which are respectively denoted as $\tilde{\varepsilon}_{E_{X},\infty}$ and $\tilde{\varepsilon}_{E_{X},2}$, are computed as
$\tilde{\varepsilon}_{E_{X},\infty}=\frac{\max_{t_{i}\in\left[0,10\right],1\leq l\leq 50000}\left|\tilde{E}_{X_{5},l}(t_{i})-E_{X_{5},l}(t_{i})\right|}{\max_{t_{i}\in\left[0,10\right],1\leq l\leq 50000}\left|E_{E_{5},l}(t_{i})\right|}$ (36)
and
$\tilde{\varepsilon}_{E_{X},2}=\frac{\sqrt{\sum\limits_{l=1}^{50000}{\sum\limits_{i=0}^{999}\left[{{E}_{X_{5},l}}(t_{i})-\tilde{E}_{X_{5},l}(t_{i})\right]^{2}}}}{\sqrt{\sum\limits_{l=1}^{50000}\sum\limits_{i=0}^{999}E_{X_{5},l}^{2}(t_{i})}}.$ (37)
Refer to caption(a)
Refer to caption(b)
Figure 1: Neural oscillator approximation errors $\tilde{\varepsilon}_{E_{X},\infty}$ and $\tilde{\varepsilon}_{E_{X},2}$ versus the number $H_{{\it{\Pi}}_{i}}$ of hidden layers in ${\it{\Pi}}_{i}(\cdot)$. (a) $\tilde{\varepsilon}_{E_{X},\infty}$ versus $H_{{\it{\Pi}}_{i}}$. (b) $\tilde{\varepsilon}_{E_{X},2}$ versus $H_{{\it{\Pi}}_{i}}$.
For the five trained neural oscillators, the decays of their $\tilde{\varepsilon}_{E_{X},\infty}$ and $\tilde{\varepsilon}_{E_{X},2}$ with respect to the increasing hidden layer numbers $H_{{{\it{\Pi}}}_{i}}$ of their MLPs ${\it{\Pi}}_{i}(\cdot)$ are shown in Figure 1. In this numerical case, each neural oscillator uses its second-order ODE to encode the input $U_{\text{e}}(t)$ and employs its MLP ${\it{\Pi}}_{i}(\cdot)$ to map the encoded features to the output function $\mathbf{y}(t)$. The results in Figure 1 illustrate that the decays of $\tilde{\varepsilon}_{E_{X},\infty}$ and $\tilde{\varepsilon}_{X,2}$ with increasing hidden layer numbers $H_{{{\it{\Pi}}}_{i}}$ of ${\it{\Pi}}_{i}(\cdot)$ can be approximated by two curves with the same decay rate of 1/7-1/7. This value coincides with the theoretical decay rate $1/\left[p(M_{{\it{\Gamma}}}+1+1)\right]=-1/\left[1\times(5+1)+1\right]=-1/7$ in Eq. (20) of Theorem 1. Such agreement confirms the validity of the decay rate established in Theorem 1.
4.2 Case 2
In this case, $U_{\text{e}}(t)$ in Eq. (33) is a stochastic Gaussian harmonizable earthquake ground motion acceleration whose Wigner-Ville spectrum is (Huang and Beer, 2024)
${W_{\text{e}}}(t,f)=2500{t^{2}}{f^{2}}\exp\left[{-0.3(1+{f^{2}})t}\right].$ (38)
$5\times 10^{4}$ independent discrete-time samples $U_{\text{e},l}(t_{i})$ of $U_{\text{e}}(t)$ are simulated. Their induced $5\times 10^{4}$ samples $X_{5,l}(t_{i})$ of the response $X_{5}(t)$ are computed. In this case, the mapping from $U_{\text{e}}(t)$ to $X_{5}(t)$ is learned using the neural oscillator with the simulated samples $U_{\text{e},l}(t_{i})$ and $X_{5,l}(t_{i})$, and the decay rate in Theorem 2 is validated.
In this case, six neural oscillators ${\it{\Pi}}_{i}\circ{\Phi}_{{\it{\Gamma}}_{i}}\left(\cdot\right)$, $i=1,2,\dots,6$, are considered. The activation functions of all MLPs in the six neural oscillators are $\sigma_{\mathrm{ReLU}}(\cdot)$. The sizes of the six one-hidden-layer MLPs ${{\it{\Pi}}}_{i}(\cdot)$ of the six neural oscillators are the same. Their widths of input, hidden, and output layers are 12, 20, and 1, respectively. The six one-hidden-layer MLPs ${{\it{\Gamma}}}_{i}(\cdot)$ of the six neural oscillators share the same input and output layer widths of 21 and 10, respectively. The hidden layer widths of ${{\it{\Gamma}}}_{1}(\cdot)$, ${{\it{\Gamma}}}_{2}(\cdot)$, ${{\it{\Gamma}}}_{3}(\cdot)$, ${{\it{\Gamma}}}_{4}(\cdot)$, ${{\it{\Gamma}}}_{5}(\cdot)$, and ${{\it{\Gamma}}}_{6}(\cdot)$ are 2, 5, 10, 20, 30, and 40, respectively.
The $5\times 10^{4}$ samples $U_{\text{e},l}(t_{i})$ of input $U_{\text{e}}(t)$ are arranged in a descending order with respect to their standard deviations. Then the $1^{\text{st}}$, $26^{\text{th}}$, $51^{\text{st}}$$\dots,$ $49976^{\text{th}}$ input samples and their induced samples $X_{5,l}(t_{i})$ of $X_{5}(t)$ (total 2000 input and response sample pairs) are selected for training. Among these 2000 sample pairs, the $5^{\text{th}}$, $10^{\text{th}}$, $15^{\text{th}}$$\dots,$ $2000^{\text{th}}$ pairs (total 400 pairs) are selected as a validation dataset, and the remaining 1600 sample pairs constitute a training dataset. Each of the six neural oscillators is trained using the training approach outlined in E with these 2000 sample pairs. The mean squared loss function $\ell_{2}$ in Eq. (118) is used. The training phase comprises 5000 epochs, with a batch size of 1600 for each epoch. The initial learn rate is 0.02. The learning rate drop period is 100 epochs. The learning rate drop factor is set to 0.965. A gradient norm clipping strategy (Pascanu et al., 2013) with a threshold of 1 is applied to all MLPs. The parameters of all MLPs are randomly initialized. For each neural oscillator, the neural network after the last update is selected as the final model.
Refer to caption(a)
Refer to caption(b)
Figure 2: Neural oscillator approximation errors $\tilde{\varepsilon}_{X,\infty}$ and $\tilde{\varepsilon}_{X,2}$ versus the hidden layer width $w_{{{\it{\Gamma}}}_{i}}$ of ${\it{\Gamma}}_{i}(\cdot)$. (a) $\tilde{\varepsilon}_{X,\infty}$ versus $w_{{\it{\Gamma}}_{i}}$. (b) $\tilde{\varepsilon}_{X,2}$ versus $w_{{\it{\Gamma}}_{i}}$.
For each trained neural oscillator, its $5\times 10^{4}$ prediction samples $\tilde{X}_{5,l}(t_{i})$ driven by the input samples $U_{\text{e},l}(t_{i})$ are computed. The relative approximation errors $\tilde{\varepsilon}_{X,\infty}$ and $\tilde{\varepsilon}_{X,2}$ are computed by replacing $\tilde{E}_{X_{5},l}(t_{i})$ and $E_{X_{5},l}(t_{i})$ in Eq. (36) and Eq. (37) with $\tilde{X}_{5,l}(t_{i})$ and $X_{5,l}(t_{i})$, respectively. For the six trained neural oscillators, the decays of their $\tilde{\varepsilon}_{X,\infty}$ and $\tilde{\varepsilon}_{X,2}$ with respect to the increase of hidden layer widths $w_{{\it{\Gamma}}_{i}}$ of the MLPs ${\it{\Gamma}}_{i}(\cdot)$ are shown in Figure 2. The ODE system in Eq. (33) consists of five second-order ODEs with respect to $\mathbf{X}(t)$ and five first-order ODEs with respect to $\mathbf{Z}(t)$. In this case, the each neural oscillator uses its second-order ODE to approximate the ODE system in Eq. (33) and employs its ${\it{\Pi}}_{i}(\cdot)$ as a readout function. The result in Figure 2(a) illustrates that the decay of $\tilde{\varepsilon}_{X,\infty}$ with increasing hidden layer widths $w_{{\it{\Gamma}}_{i}}$ of ${\it{\Gamma}}_{i}(\cdot)$ can be well represented by a curve with a decay rate of 0.5-0.5. This value is the same as the decay rate in Eq. (31) of Theorem 2, where the approximation error is under $L^{\infty}$-norm. The decay rate of $\tilde{\varepsilon}_{X,2}$ under $L^{2}$-norm is slightly smaller than 0.5-0.5. These results demonstrate that the approximation error decay rate in Theorem 2 is reasonable.
5 Conclusions
In this study, two upper approximation bounds of the neural oscillator consisting of a second-order ODE followed by an MLP are derived for approximating the causal and uniformly continuous operators between continuous temporal function spaces and for approximating the uniformly asymptotically incrementally stable second-order dynamical systems. The established proof method of the upper approximation bound for approximating the causal continuous operators can also be directly applied to the SS models consisting of a linear time-continuous complex RNN followed by an MLP. In this study, Theorem 1 indicates that for the causal continuous operators, the decay rate in the approximation error polynomially diminishes as the hidden-layer width of the MLP ${\it{\Gamma}}(\cdot)$ increases. Thus, the neural oscillator for approximating the causal continuous operators between continuous temporal function spaces also faces the curse of parametric complexity. Theorem 2 demonstrates that for approximating the uniformly asymptotically incrementally stable second-order dynamical systems, the neural oscillator can provide an approximation error scaling polynomially with the reciprocals of the widths of the utilized MLPs. This implies that the curse of parametric complexity can be mitigated in this situation. The approximation error decay rate with respect to the hidden layer number of the MLP ${\it{\Pi}}(\cdot)$ in Theorem 1 and that with respect to the hidden layer width of the MLP ${\it{\Gamma}}(\cdot)$ in Theorem 2 have been validated through two numerical cases.
This work only focuses on the approximation error bounds for the neural oscillator. Future research will address its generalization bounds to provide a more comprehensive theoretical understanding.
CRediT authorship contribution statement
Zifeng Huang: Conceptualization, Methodology, Software, Writing-reviewing & editing, Project administration. Konstantin M. Zuev: Methodology, Writing-reviewing & editing. Yong Xia: Writing-reviewing & editing, Supervision, Project administration. Michael Beer: Supervision, Project administration.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgments
This work is supported by the Joint Research Centre for Marine Infrastructure at The Hong Kong Polytechnic University and the German Research Foundation (Project No. 554325683 and HU 3418/2-1). The opinions and conclusions presented in this paper are entirely those of the authors.
Appendix A Proofs for Lemmas and Theorem 1 in Section 3.1
A.1 Proof of Lemma 1
Proof: Given an arbitrary $\mathbf{u}(t)=[u_{1}(t),u_{2}(t),...,u_{p}(t)]^{\top}\in K$, since $K$ is compact, based on Assumption 2, there exists a continuous monotonic modulus of continuity $\mathbf{\upphi}_{K}(t)$ such that for arbitrary $0\leq t_{1}\leq t_{2}\leq T$ and all $\mathbf{u}(t)\in K$, it satisfies $|\mathbf{u}(t_{1})-\mathbf{u}(t_{2})|\leq|{\mathbf{\upphi}}_{K}(t_{2}-t_{1})|$. Define $\mathbf{f}_{t}(\tau)=\mathbf{u}(t-\tau)-\mathbf{u}(0)$ for $\tau\in[0,t]$ and
[…]
${\tilde{\mathbf{f}}}_{t}(\tau)$ is odd and compactly supported, and its Fourier transform $\mathcal{H}{{{\tilde{\mathbf{f}}}}_{t}}(\omega)$ is
$\begin{array}[]{l}\mathcal{H}{{{\tilde{\mathbf{f}}}}_{t}}(\omega)=\int_{-\infty}^{+\infty}{{e^{-{\rm{i}}\omega\tau}}{{{\tilde{\mathbf{f}}}}_{t}}(\tau){\rm{d}}\tau}=-{\rm{i}}\int_{-t}^{t}{\sin(\omega\tau){{{\tilde{\mathbf{f}}}}_{t}}(\tau){\rm{d}}\tau}\\ \;\;\;\;\;\;\;\;\;\;\;\;=-2{\rm{i}}\int_{0}^{t}{\sin(\omega\tau){{\bf{f}}_{t}}(\tau){\rm{d}}\tau}=-2{\rm{i}}\int_{0}^{t}{\sin(\omega\tau)\left[{{\bf{u}}(t-\tau)-{\bf{u}}(0)}\right]{\rm{d}}\tau}\\ \;\;\;\;\;\;\;\;\;\;\;\;=-2{\rm{i}}\left\{{{\mathcal{L}_{t}}{\bf{u}}(\omega)+\frac{{{\bf{u}}(0)}}{\omega}\left[{\cos(\omega t)-1}\right]}\right\}.\end{array}$ (40)
Consider a smooth bump mollifier function $\rho_{v}(\tau)$
[…]
where $v$ is an arbitrary coefficient satisfying $0<v<T$ and $c_{v}$ is a normalization constant depending on $v$ such that $\int_{-v}^{0}\rho_{v}(\tau)\mathrm{d}\tau=1$. The mollification $\tilde{\mathbf{f}}_{v}(\tau)$ of $\tilde{\mathbf{f}}_{t}(\tau)$ is
${{\tilde{\mathbf{f}}}_{v}}(\tau)=\int_{-\infty(-v)}^{+\infty(0)}{{{{\tilde{\mathbf{f}}}}_{t}}(\tau-\lambda){\rho_{v}}(\lambda){\rm{d}}\lambda}.$ (42)
$\tilde{\mathbf{f}}_{v}(\tau)$ is employed to approximate $\tilde{\mathbf{f}}_{t}(\tau)=\mathbf{u}(t-\tau)-\mathbf{u}(0)$ over $\tau\in[0,t]$ and their difference is bounded by
$\begin{array}[]{l}\left|{{{{\tilde{\mathbf{f}}}}_{v}}(\tau)-\left[{{\bf{u}}(t-\tau)-{\bf{u}}(0)}\right]}\right|\\ \;\;\;\;\;=\left|{{{{\tilde{\mathbf{f}}}}_{v}}(\tau)-{{{\tilde{\mathbf{f}}}}_{t}}(\tau)}\right|=\left|{\int_{-v}^{0}{\left[{{{{\tilde{\mathbf{f}}}}_{t}}(\tau-\lambda)-{{{\tilde{\mathbf{f}}}}_{t}}(\tau)}\right]{\rho_{v}}(\lambda){\rm{d}}\lambda}}\right|\\ \;\;\;\;\;\leq\int_{-v}^{0}{\left|{{{{\tilde{\mathbf{f}}}}_{t}}(\tau-\lambda)-{{{\tilde{\mathbf{f}}}}_{t}}(\tau)}\right|{\rho_{v}}(\lambda){\rm{d}}\lambda}\mathop{\leq}\limits^{(a)}\int_{-v}^{0}{\left|{\mathbf{\upphi}_{K}(v)}\right|{\rho_{v}}(\lambda){\rm{d}}\lambda}=\left|{\mathbf{\upphi}_{K}(v)}\right|,\end{array}$ (43)
where Step (a) holds from $|\mathbf{u}(t_{1})-\mathbf{u}(t_{2})|\leq\left|{\mathbf{\upphi}}_{K}(t_{2}-t_{1})\right|$ and $\left|{\mathbf{\upphi}}_{K}(t)\right|$ being non-negative and monotonically increasing.
$\tilde{\mathbf{f}}_{v}(\tau)$, $\tilde{\mathbf{f}}_{t}(\tau)$, and $\rho_{v}(\tau)$ are compactly supported over $\tau\in[-2T,2T]$ for all $t\in[0,T]$ and all $v\in\left(0,T\right)$. $\tilde{\mathbf{f}}_{v}(\tau)$ and $\rho_{v}(\tau)$ are continuous over $\tau\in\mathbb{R}$. $\tilde{\mathbf{f}}_{v}(\tau)$ is odd with respect to $\tau=-0.5v$ because $\tilde{\mathbf{f}}_{t}(\tau)$ is an odd function and $\rho_{v}(\tau)$ is symmetric with respect to $\tau=-0.5v$, thereby $\int_{-\infty}^{+\infty}\tilde{\mathbf{f}}_{v}(\tau)\mathrm{d}\tau=\int_{-2T}^{2T}\tilde{\mathbf{f}}_{v}(\tau)\mathrm{d}\tau=\mathbf{0}$. Then, $\tilde{\mathbf{f}}_{v}(\tau)$ over $\tau\in[-2T,2T]$ can be expressed using the Fourier series
${{\tilde{\mathbf{f}}}_{v}}(\tau)=2\sum\limits_{n=1}^{+\infty}{{\mathop{\rm Re}\nolimits}\left[{{e^{{\rm{i}}{\omega_{n}}\tau}}\mathcal{S}{{{\tilde{\mathbf{f}}}}_{v}}({\omega_{n}})}\right]}\mathop{=}\limits^{(a)}\frac{1}{{2T}}\sum\limits_{n=1}^{+\infty}{{\mathop{\rm Re}\nolimits}\left[{{e^{{\rm{i}}{\omega_{n}}\tau}}\mathcal{H}{{{\tilde{\mathbf{f}}}}_{t}}({\omega_{n}})\mathcal{H}{\rho_{v}}({\omega_{n}})}\right]},$ (44)
where $\omega_{n}=\pi n/{(2T)}$, $\mathcal{S}{{\tilde{\mathbf{f}}}}_{v}(\omega_{n})=1/(4T)\int_{-2T}^{2T}e^{-\mathrm{i}\omega_{n}}\tilde{\mathbf{f}}_{v}(\tau){\mathrm{d}\tau}=\mathcal{H}{{{\tilde{\mathbf{f}}}}_{t}}({\omega_{n}})\mathcal{H}{\rho_{v}}({\omega_{n}})/(4T)$ is the Fourier coefficient of $\tilde{\mathbf{f}}_{v}(\tau)$, Step (a) is valid because $\tilde{\mathbf{f}}_{v}(\tau)$ is the convolution between $\tilde{\mathbf{f}}_{t}(\tau)$ and $\rho_{v}(\tau)$, and $\mathcal{H}\rho_{v}(\omega)$ is the Fourier transform of $\rho_{v}(\tau)$. $\mathcal{H}{{{\tilde{\mathbf{f}}}}_{t}}({\omega_{n}})\mathcal{H}{\rho_{v}}({\omega_{n}})/(2T)$ in Eq. (44) can be bounded by
$\begin{array}[]{l}\frac{1}{{{2T}}}\left|{\mathcal{H}{{{\tilde{\mathbf{f}}}}_{t}}({\omega_{n}})\mathcal{H}{\rho_{v}}({\omega_{n}})}\right|\\ \;\;\;\;\;=\frac{1}{{{2T}}}\left|{\int_{-{2T}}^{{2T}}{{e^{-{\rm{i}}{\omega_{n}}\tau}}{{{\tilde{\mathbf{f}}}}_{v}}(\tau){\rm{d}}\tau}}\right|\\ \;\;\;\;\;=\frac{1}{{{2T}}}\left|{\left.{\frac{1}{{-{\rm{i}}{\omega_{n}}}}{e^{-{\rm{i}}{\omega_{n}}\tau}}{{{\tilde{\mathbf{f}}}}_{v}}(\tau)}\right|_{-{2T}}^{{2T}}-\frac{1}{{-{\rm{i}}{\omega_{n}}}}\int_{-{2T}}^{{2T}}{{e^{-{\rm{i}}{\omega_{n}}\tau}}{{{\tilde{\mathbf{f}}^{\prime}}}_{v}}(\tau){\rm{d}}\tau}}\right|\\ \;\;\;\;\;=\frac{1}{{{2T}}}\left|{\left.{-\frac{1}{{-\omega_{n}^{2}}}{e^{-{\rm{i}}{\omega_{n}}\tau}}{{{\tilde{\mathbf{f}}^{\prime}}}_{v}}(\tau)}\right|_{-{2T}}^{{2T}}-\frac{1}{{\omega_{n}^{2}}}\int_{-{2T}}^{{2T}}{{e^{-{\rm{i}}{\omega_{n}}\tau}}{{{\tilde{\mathbf{f}}^{\prime\prime}}}_{v}}(\tau){\rm{d}}\tau}}\right|\\ \;\;\;\;\;\leq\frac{{{2T}}}{{{\pi^{2}}{n^{2}}}}\int_{-{2T}}^{{2T}}{\left|{{{{\tilde{\mathbf{f}}^{\prime\prime}}}_{v}}(\tau)}\right|{\rm{d}}\tau}\leq\frac{{8T^{2}}}{{{\pi^{2}}{n^{2}}}}\mathop{\max}\limits_{\tau\in[-{2T},{2T}]}\left|{{{{\tilde{\mathbf{f}}^{\prime\prime}}}_{v}}(\tau)}\right|.\end{array}$ (45)
In Eq. (45), since $\rho_{v}(\tau)$ has arbitrary finite-order continuous derivatives and $\tilde{\mathbf{f}}_{v}(\tau)$, $\tilde{\mathbf{f}}_{t}(\tau)$, and $\rho_{v}(\tau)$ are compactly supported over $\tau\in[-2T,2T]$ for all $t\in[0,T]$, the boundary conditions ${\tilde{\mathbf{f}}}_{v}(-2T)={\tilde{\mathbf{f}}}_{v}(2T)={\tilde{\mathbf{f}}}_{v}^{\prime}(-2T)={\tilde{\mathbf{f}}}_{v}^{\prime}(2T)=\mathbf{0}$ are employed. ${\tilde{\mathbf{f}}}_{v}^{\prime\prime}(\tau)$ in Eq. (45) over $\tau\in[-2T,2T]$ for all $t\in[0,T]$ is bounded by
$\begin{array}[]{l}\left|{{{\tilde{\mathbf{f}}^{\prime\prime}}_{v}}(\tau)}\right|=\left|{\frac{{{{\rm{d}}^{2}}}}{{{\rm{d}}{\tau^{2}}}}\int_{-\infty}^{+\infty}{{{{\tilde{\mathbf{f}}}}_{t}}(\lambda){\rho_{v}}(\tau-\lambda){\rm{d}}\lambda}}\right|\leq\int_{-\infty}^{+\infty}{\left|{{{{\tilde{\mathbf{f}}}}_{t}}(\lambda)}\right|\left|{\frac{{{{\rm{d}}^{2}}}}{{{\rm{d}}{\tau^{2}}}}{\rho_{v}}(\tau-\lambda)}\right|{\rm{d}}\lambda}\\ \;\;\;\;\;\;\;\;\;\;\leq\int_{\tau}^{\tau+v}{\left|{{{{\tilde{\mathbf{f}}}}_{t}}(\lambda)}\right|\mathop{\max}\limits_{t\in[-v,0]}\left|{{{\rho}_{v}^{\prime\prime}}(t)}\right|{\rm{d}}\lambda}\leq\int_{0}^{v}{\mathop{\max}\limits_{t\in[0,T]}\left|{{\bf{u}}(t)-{\bf{u}}(0)}\right|\mathop{\max}\limits_{t\in[-v,0]}\left|{{{\rho}_{v}^{\prime\prime}}(t)}\right|{\rm{d}}\lambda}\\ \;\;\;\;\;\;\;\;\;\;\mathop{\leq}\limits^{(a)}v\left|{\mathbf{\upphi}_{K}(T)}\right|\mathop{\max}\limits_{t\in[-v,0]}\left|{{{\rho}_{v}^{\prime\prime}}(t)}\right|\mathop{\leq}\limits^{(b)}\left|{\mathbf{\upphi}_{K}(T)}\right|\frac{{{c_{\rho,2}}}}{{{v^{2}}}},\end{array}$ (46)
where Step (a) employs $|\mathbf{u}(t)-\mathbf{u}(0)|\leq\left|{\mathbf{\upphi}}_{K}(t)\right|\leq\left|{\mathbf{\upphi}}_{K}(T)\right|$, Step (b) holds from $\max_{\tau\in[-v,0]}\left|{\rho}_{v}^{\prime\prime}(\tau)\right|$ $\leq c_{\rho,2}/v^{3}$, and $c_{\rho,2}=\max_{\tau\in[-1,0]}\left|{\rho}_{1}^{\prime\prime}(\tau)\right|$ is an independent constant. The bound of $|{\tilde{\mathbf{f}}}_{v}^{\prime\prime}(\tau)|$ in Eq. (46) is independent of t and $\tau$. Then, based on Eqs. (44)-(46), for every positive integer M larger than one, it can be obtained
$\begin{array}[]{l}\left|{{{{\tilde{\bf f}}}_{v}}(\tau)-\frac{1}{{2T}}\sum\limits_{n=1}^{M}{{\mathop{\rm Re}\nolimits}\left[{{e^{{\rm{i}}{\omega_{n}}\tau}}\mathcal{H}{{{\tilde{\bf f}}}_{t}}({\omega_{n}})\mathcal{H}{\rho_{\nu}}({\omega_{n}})}\right]}}\right|\\ \;\;\;=\left|{\frac{1}{{2T}}\sum\limits_{n=M+1}^{+\infty}{{\mathop{\rm Re}\nolimits}\left[{{e^{{\rm{i}}{\omega_{n}}\tau}}\mathcal{H}{{{\tilde{\bf f}}}_{t}}({\omega_{n}})\mathcal{H}{\rho_{\nu}}({\omega_{n}})}\right]}}\right|\\ \;\;\;\leq\frac{1}{{2T}}\sum\limits_{n=M+1}^{+\infty}{\left|{\mathcal{H}{{{\tilde{\bf f}}}_{t}}({\omega_{n}})\mathcal{H}{\rho_{\nu}}({\omega_{n}})}\right|}\mathop{\leq}\limits^{(a)}\sum\limits_{n=M+1}^{+\infty}{\frac{{8{T^{2}}}}{{{\pi^{2}}{n^{2}}}}\mathop{\max}\limits_{\tau\in[-{2T},{2T}]}\left|{{{{\tilde{\mathbf{f}}}^{\prime\prime}}_{v}}(\tau)}\right|}\\ \;\;\;\leq\frac{8T^{2}}{{{\pi^{2}}}}{\mathop{\max}\limits_{\tau\in[-2T,2T]}\left|{{{{\tilde{\mathbf{f}}}^{\prime\prime}}_{v}}(\tau)}\right|}\sum\limits_{n=M+1}^{+\infty}{\frac{1}{{{n^{2}}}}}\mathop{\leq}\limits^{(b)}\frac{{8T^{2}\left|{\mathbf{\upphi}}_{K}(T)\right|{c_{\rho,2}}}}{{{\pi^{2}}{v^{2}}}}\sum\limits_{n=M+1}^{+\infty}{\frac{1}{{{n^{2}}}}}\mathop{\leq}\limits^{(c)}\frac{{8T^{2}\left|{\mathbf{\upphi}}_{K}(T)\right|{c_{\rho,2}}}}{{{\pi^{2}}{v^{2}}M}},\end{array}$ (47)
where Steps (a) and (b) are from Eqs. (45) and (46), respectively, and Step (c) is derived from the integral test for convergence, see Section 3.3 in (Knopp, 1956),
$\sum\limits_{n=M+1}^{+\infty}{\frac{1}{{{n^{2}}}}}\leq\sum\limits_{m=0}^{+\infty}{\int_{M+m}^{M+m+1}{\frac{1}{{{x^{2}}}}{\rm{d}}x}}=\int_{M}^{+\infty}{\frac{1}{{{x^{2}}}}{\rm{d}}x}=\frac{1}{M},\;\;m,n\in\mathbb{Z}.$ (48)
Finally, based on Eqs. (43) and (47), for every positive integer M larger than one, $\mathbf{u}(t-\tau)$ over $\tau\in[0,t]$ satisfies
$\begin{array}[]{l}\left|{{\bf{u}}(t-\tau)-\left\{{\frac{1}{{2T}}\sum\limits_{n=1}^{M}{{\mathop{\rm Re}\nolimits}\left[{{e^{{\rm{i}}{\omega_{n}}\tau}}\mathcal{H}{{{\tilde{\mathbf{f}}}}_{t}}({\omega_{n}})\mathcal{H}{\rho_{v}}({\omega_{n}})}\right]}+{\bf{u}}(0)}\right\}}\right|\\ \;\;\;\;=\left|{{\bf{u}}(t-\tau)-{\bf{u}}(0)-\frac{1}{{2T}}\sum\limits_{n=1}^{M}{{\mathop{\rm Re}\nolimits}\left[{{e^{{\rm{i}}{\omega_{n}}\tau}}\mathcal{H}{{{\tilde{\mathbf{f}}}}_{t}}({\omega_{n}})\mathcal{H}{\rho_{v}}({\omega_{n}})}\right]}}\right|\\ \;\;\;\;\leq\left|{{\bf{u}}(t-\tau)-{\bf{u}}(0)-{{{\tilde{\mathbf{f}}}}_{v}}(\tau)}\right|+\left|{{{{\tilde{\mathbf{f}}}}_{v}}(\tau)-\frac{1}{{2T}}\sum\limits_{n=1}^{M}{{\mathop{\rm Re}\nolimits}\left[{{e^{{\rm{i}}{\omega_{n}}\tau}}\mathcal{H}{{{\tilde{\mathbf{f}}}}_{t}}({\omega_{n}})\mathcal{H}{\rho_{v}}({\omega_{n}})}\right]}}\right|\\ \;\;\;\;\leq\left|{\mathbf{\upphi}_{K}(v)}\right|+\frac{{8T^{2}\left|{\mathbf{\upphi}}_{K}(T)\right|{c_{\rho,2}}}}{{{\pi^{2}}{v^{2}}M}}\end{array}$ (49)
for all $t\in[0,T]$ and all $\mathbf{u}(t)\in K$. Based on Eq. (40), $1/(2T){\mathrm{Re}}[e^{\mathrm{i}\omega_{n}\tau}\mathcal{H}\tilde{\mathbf{f}}_{t}(\omega_{n})\mathcal{H}\rho_{v}(\omega_{n})]$ in Eq. (49) can be expressed as
$\begin{array}[]{l}\frac{1}{{2T}}{\mathop{\rm Re}\nolimits}\left[{{e^{{\rm{i}}{\omega_{n}}\tau}}\mathcal{H}{{{\tilde{\mathbf{f}}}}_{t}}({\omega_{n}})\mathcal{H}{\rho_{v}}({\omega_{n}})}\right]\\ \;\;\;\;\;=\frac{1}{T}\left\{{{\mathcal{L}_{t}}{\bf{u}}({\omega_{n}})+\frac{{{\bf{u}}(0)}}{{{\omega_{n}}}}\left[{\cos({\omega_{n}}t)-1}\right]}\right\}\left\{{{\mathop{\rm Re}\nolimits}\left[{\mathcal{H}{\rho_{v}}({\omega_{n}})}\right]\sin({\omega_{n}}\tau)+{\mathop{\rm Im}\nolimits}\left[{\mathcal{H}{\rho_{v}}({\omega_{n}})}\right]\cos({\omega_{n}}\tau)}\right\}\\ \;\;\;\;\;={\alpha_{n}}\left\{{{\mathcal{L}_{t}}{\bf{u}}({\omega_{n}})+\frac{{{\bf{u}}(0)}}{{{\omega_{n}}}}\left[{\cos({\omega_{n}}t)-1}\right]}\right\}\sin({\omega_{n}}\tau-{\theta_{n}}),\end{array}$ (50)
where $\alpha_{n}$ and $\theta_{n}$ are the scaled Fourier modulus and phase of $\rho_{v}(\tau)$, respectively, and they are independent of the input function $\mathbf{u}(t)$. $\alpha_{n}$ can be bounded by
$\alpha_{n}=\frac{1}{T}\left|\int_{-2T}^{2T}e^{-\text{i}\omega_{n}\tau}\rho_{v}(\tau)\text{d}\tau\right|\leq\frac{1}{T}\int_{-2T}^{2T}\rho_{v}(\tau)\text{d}\tau=\frac{1}{T}.$ (51)
Eq. (9) can be proven by substituting Eq. (50) into Eq. (49). $\blacksquare$
A.2 Proof of Lemma 2
Proof: For every positive integer $M$ larger than one and arbitrary $M$ real frequencies $\omega_{1}$, $\omega_{2}$,…, $\omega_{M}$, with $\sigma_{\mathrm{ReLU}}(x)-\sigma_{\mathrm{ReLU}}(-x)=x$, there exists a set of equations with the initial conditions $\mathbf{x}(0)=\mathbf{x}^{\prime}(0)=\mathbf{0}$
$\left\{\begin{aligned} &{{{\bf{x}}}_{1}^{\prime\prime}}(t)={\sigma_{{\rm{ReLU}}}}\left[{-\omega_{1}^{2}{{\bf{x}}_{1}}(t)-0{{{\bf{x}}}_{1}^{\prime}}(t)+{\omega_{1}}{\bf{u}}(t)}\right]-{\sigma_{{\rm{ReLU}}}}\left[{\omega_{1}^{2}{{\bf{x}}_{1}}(t)+0{{{\bf{x}}}_{1}^{\prime}}(t)-{\omega_{1}}{\bf{u}}(t)}\right]\\ &{{{\bf{x}}}_{2}^{\prime\prime}}(t)={\sigma_{{\rm{ReLU}}}}\left[{-\omega_{2}^{2}{{\bf{x}}_{2}}(t)-0{{{\bf{x}}}_{2}^{\prime}}(t)+{\omega_{2}}{\bf{u}}(t)}\right]-{\sigma_{{\rm{ReLU}}}}\left[{\omega_{2}^{2}{{\bf{x}}_{2}}(t)+0{{{\bf{x}}}_{2}^{\prime}}(t)-{\omega_{2}}{\bf{u}}(t)}\right]\\ &\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\ \vdots\\ &{{{\bf{x}}}_{M}^{\prime\prime}}(t)={\sigma_{{\rm{ReLU}}}}\left[{-\omega_{M}^{2}{{\bf{x}}_{M}}(t)-0{{{\bf{x}}}_{M}^{\prime}}(t)+{\omega_{M}}{\bf{u}}(t)}\right]-{\sigma_{{\rm{ReLU}}}}\left[{\omega_{M}^{2}{{\bf{x}}_{M}}(t)+0{{{\bf{x}}}_{M}^{\prime}}(t)-{\omega_{M}}{\bf{u}}(t)}\right]\\ \end{aligned},\right$. (52)
such that its solution satisfies $\mathbf{x}(t)=[\mathbf{x}_{1}^{\top}(t),\mathbf{x}_{2}^{\top}(t),...,\mathbf{x}_{M}^{\top}(t)]^{\top}=[\mathcal{L}_{t}\mathbf{u}^{\top}(\omega_{1}),\mathcal{L}_{t}\mathbf{u}^{\top}(\omega_{2}),...,\mathcal{L}_{t}\mathbf{u}^{\top}(\omega_{M})]^{\top}$ for all $t\in[0,T]$ and all $\mathbf{u}(t)\in K$. Eq. (52) can be expressed in the form of the second-order ODE in Eq. (1) with a one-hidden-layer MLP ${\it{\Gamma}}(\cdot)$ utilizing $\sigma_{\mathrm{ReLU}}(\cdot)$. The widths of the input, hidden, and output layers of ${\it{\Gamma}}(\cdot)$ are $w_{{\it{\Gamma},{\mathrm{in}}}}=p(2M+1)$, $w_{\it{\Gamma}}=2pM$, and $w_{{\it{\Gamma},{\mathrm{out}}}}=pM$, respectively. $\blacksquare$
A.3 Proof of Lemma 3
Proof: From Lemma 1 and Lemma 2, for an arbitrary positive $v<T$, and every positive integer $M$ larger than one, there exists an MLP ${\it{\Gamma}}(\cdot)$ described in Lemma 2, such that for every $\mathbf{u}(t)\in K$, its corresponding solution $\mathbf{x}(t)=[\mathbf{x}_{1}^{\top}(t),\mathbf{x}_{2}^{\top}(t),...,\mathbf{x}_{M}^{\top}(t)]^{\top}$ from Eq. (1) described in Lemma 2 satisfies
$\sup_{\tau\in[0,t]}\left|{{\bf{u}}(\tau)-{{{\hat{\mathbf{u}}}}_{t}}(\tau)}\right|\leq{\varepsilon_{K}}$ (53)
for all $t\in[0,T]$, where $\varepsilon_{K}$ is in Eq. (11), ${\hat{\mathbf{u}}}_{t}(\tau)$ over $\tau\in[0,t]$ is expressed as
${{\hat{\mathbf{u}}}_{t}}(\tau)={\bf{u}}(0)+\sum\limits_{n=1}^{M}{{\alpha_{n}}\left\{{{{\bf{x}}_{n}}(t)+\frac{{{\bf{u}}(0)}}{{{\omega_{n}}}}\left[{\cos({\omega_{n}}t)-1}\right]}\right\}\sin\left[{{\omega_{n}}(t-\tau)-{\theta_{n}}}\right]},$ (54)
$\mathbf{x}_{n}(t)$ is the $n^{\text{th}}$ vector element of $\mathbf{x}(t)$ that can be calculated as
${{\bf{x}}_{n}}(t)={\mathcal{L}_{t}}{\bf{u}}({\omega_{n}})=\int_{0}^{t}{{\bf{u}}(t-\tau)\sin({\omega_{n}}\tau){\rm{d}}\tau},$ (55)
$\omega_{1}$, $\omega_{2}$,…, $\omega_{M}$ are related to the network parameters of ${\it{\Gamma}}(\cdot)$, and $\alpha_{n}$ and $\theta_{n}$, n=1,2,,Mn=1,2,...,M, are specified in Lemma 1.
Since $\Phi$ is a causal and uniformly continuous operator from $C_{0}([0,T];{\mathbb{R}^{p}})$ to $C_{0}([0,T];{\mathbb{R}^{q}})$, and the parameters $\omega_{n}$, $\alpha_{n}$, and $\theta_{n}$ in Eq. (54) are independent of the input function $\mathbf{u}(t)$ according to the proof of Lemma 1, ${\Phi}[{\hat{\mathbf{u}}}_{t}(\tau)](t)$ can be regarded as a continuous function with respect to the values of $\mathbf{x}(t)$, $\mathbf{u}(0)$, and time $t$. The continuity of ${\Phi}[{\hat{\mathbf{u}}}_{t}(\tau)](t)$ with respect to the values of $\mathbf{x}(t)$ and $\mathbf{u}(0)$ is evident because $\Phi$ is a continuous operator. The continuity with respect to the value of $t$ is analytically established through Eq. (74) in the proof of Lemma 4 in A.4. Thus, for every positive integer M larger than one, a related continuous mapping $\Psi:{\mathbb{R}}^{p(M+1)}\times{[0,T]}\to{\mathbb{R}}^{q}$ can be defined as $\Psi(\mathbf{x},\mathbf{u}_{0},t)=\Psi[\mathbf{x}(t),\mathbf{u}(0),t]={\Phi}[{\hat{\mathbf{u}}}_{t}(\tau)](t)$, where $\mathbf{x}=[\mathbf{x}_{1}^{\top},\mathbf{x}_{2}^{\top},...,\mathbf{x}_{M}^{\top}]^{\top}$, $\mathbf{x}_{n}$ and $\mathbf{u}_{0}$ represent the values of $\mathbf{x}_{n}(t)$ and $\mathbf{u}(0)$ in Eq. (54), respectively.
The value of ${\Phi}[{\hat{\mathbf{u}}}_{t}(\tau)](t)$ at $t$ is independent of the values of ${\hat{\mathbf{u}}}_{t}(\tau)$ over $\tau>t$. Given a time instant $t_{0}$, $0\leq t_{0}\leq T$, it can be obtained
$\begin{array}[]{l}|{\Phi}[\mathbf{u}(\tau)](t_{0})-\Psi(\mathbf{x},\mathbf{u}_{0},t_{0})|=|{\Phi}[\mathbf{u}(\tau)](t_{0})-{\Phi}[{\hat{\mathbf{u}}}_{t_{0}}(\tau)](t_{0})|\leq{\left\lVert{\Phi}[\mathbf{u}(\tau)](t)-{\Phi}[{\hat{\mathbf{u}}}_{t_{0}}(\tau)](t)\right\rVert}_{L^{\infty}[0,t_{0}]}\\ \;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\mathop{\leq}\limits^{(a)}{\phi_{\Phi}}\left[{\left\lVert\mathbf{u}(t)-\hat{\mathbf{u}}_{t_{0}}(t)\right\rVert}_{L^{\infty}[0,t_{0}]}\right]\mathop{\leq}\limits^{(b)}{\phi_{\Phi}}(\varepsilon_{K}),\end{array}$ (56)
where ${\phi_{\Phi}}(\cdot)=\sum_{j=1}^{q}{\phi}_{\Phi_{j}}(\cdot)$ is a monotonic modulus of continuity of $\Phi$ in Assumption 6, $\phi_{\Phi_{j}}(\cdot)$ is a monotonic modulus of continuity of the $j^{\mathrm{th}}$ element $\Phi_{j}$ of $\Phi$, $\varepsilon_{K}$ is in Eq. (11) of Lemma 1, Step (b) follows from Eq. (53), and Step (a) holds from the causality of $\Phi$ and it will be proven subsequently. Since $\varepsilon_{K}$ is independent of t, $|{\Phi}[\mathbf{u}(\tau)](t_{0})-\Psi(\mathbf{x},\mathbf{u}_{0},t_{0})|\leq{\phi_{\Phi}}(\varepsilon_{K})$ is valid for all $t_{0}\in[0,T]$, which proves Eq. (16).
${\left\lVert{\Phi}[\mathbf{u}_{1}(\tau)](t)-{\Phi}[\mathbf{u}_{2}(\tau)](t)\right\rVert}_{L^{\infty}[0,t_{0}]}\leq{\phi_{\Phi}}\left[{\left\lVert\mathbf{u}_{1}(t)-\mathbf{u}_{2}(t)\right\rVert}_{L^{\infty}[0,t_{0}]}\right]$ is satisfied for arbitrary $\mathbf{u}_{1}(t)$ and $\mathbf{u}_{2}(t)$ $\in C_{0}([0,T];{\mathbb{R}^{p}})$. This can be proven by setting
${{\bf{u}}_{i}}(t)=\left\{\begin{aligned} &{{\bf{u}}_{i}}(t),\;\;\;0\leq t\leq{t_{0}}\\ &{{\bf{u}}_{i}}({t_{0}}),\;\;{t_{0}}\leq t\leq T\end{aligned},\right.\;\;i=1\;{\rm{and}}\;2,$ (57)
then, it is satisfied ${\left\lVert{\Phi}[\mathbf{u}_{1}(\tau)](t)-{\Phi}[\mathbf{u}_{2}(\tau)](t)\right\rVert}_{L^{\infty}[0,t_{0}]}\leq{\left\lVert{\Phi}[\mathbf{u}_{1}(\tau)](t)-{\Phi}[\mathbf{u}_{2}(\tau)](t)\right\rVert}_{L^{\infty}}\leq{\phi_{\Phi}}\left[{\left\lVert\mathbf{u}_{1}(t)-\mathbf{u}_{2}(t)\right\rVert}_{L^{\infty}}\right]={\phi_{\Phi}}\left[{\left\lVert\mathbf{u}_{1}(t)-\mathbf{u}_{2}(t)\right\rVert}_{L^{\infty}[0,t_{0}]}\right]$. Since the values of $\lVert{\Phi}[\mathbf{u}_{1}(\tau)](t)$- ${\Phi}[\mathbf{u}_{2}(\tau)](t)\rVert_{L^{\infty}[0,t_{0}]}$ and ${\phi_{\Phi}}\left[{\left\lVert\mathbf{u}_{1}(t)-\mathbf{u}_{2}(t)\right\rVert}_{L^{\infty}[0,t_{0}]}\right]$ are independent of the values of $\mathbf{u}_{1}(t)$ and $\mathbf{u}_{2}(t)$ over $t_{0}\leq t\leq T$, ${\left\lVert{\Phi}[\mathbf{u}_{1}(\tau)](t)-{\Phi}[\mathbf{u}_{2}(\tau)](t)\right\rVert}_{L^{\infty}[0,t_{0}]}\leq{\phi_{\Phi}}\left[{\left\lVert\mathbf{u}_{1}(t)-\mathbf{u}_{2}(t)\right\rVert}_{L^{\infty}[0,t_{0}]}\right]$ is valid for arbitrary $\mathbf{u}_{1}(t)$ and $\mathbf{u}_{2}(t)$ $\in C_{0}([0,T];{\mathbb{R}^{p}})$. Since $\mathbf{u}(t)$ and $\hat{\mathbf{u}}_{t_{0}}(t)$ in Eq. (56) belong to $C_{0}([0,T];{\mathbb{R}^{p}})$, Step (a) in Eq. (56) is valid. $\blacksquare$
A.4 Proof of Lemma 4
Proof: Since K is compact, by Assumption 2, it is closed and there exists a constant $B_{K}$ such that ${\left\lVert u_{i}(t)\right\rVert}_{L^{\infty}}\leq B_{K}$ for all $\mathbf{u}(t)\in K$ and $1\leq i\leq p$, where $u_{i}(t)$ is the $i^{\text{th}}$ element of $\mathbf{u}(t)$. For every $M$ satisfying Eq. (15), $\Psi\left({\bf{x}},{\bf{u}}_{0},t\right)=\Psi\left[{\bf{x}}(t),{\bf{u}}(0),t\right]=\Phi[{\hat{\mathbf{u}}}_{t}(\tau)](t)$ is a continuous function over a compact domain $\Omega_{M}\left(TB_{K},B_{K},T\right)=\left[-TB_{K},TB_{K}\right]^{pM}\times\left[-B_{K},B_{K}\right]^{p}\times\left[0,T\right]$, where $TB_{K}$ is the bound for the elements of $\mathbf{x}_{n}$, n=1,2,,Mn=1,2,...,M, because $\mathbf{x}_{n}=\mathbf{x}_{n}(t)$ is calculated from Eq. (55). In $\Omega_{M}\left(TB_{K},B_{K},T\right)$, $\left[-TB_{K},TB_{K}\right]^{pM}$ is the value range of $\mathbf{x}$, $\left[-B_{K},B_{K}\right]^{p}$ is the value range of $\mathbf{u}_{0}$, and $[0,T]$ is the value range of t.
The $i^{\text{th}}$ element ${\hat{u}_{t,i}}(\tau)$ of ${\hat{\mathbf{u}}_{t}}(\tau)=[{\hat{u}_{t,1}}(\tau),{\hat{u}_{t,2}}(\tau),...,{\hat{u}_{t,p}}(\tau)]^{\top}$ in Eq. (54) can be expressed as
${\hat{u}_{t,i}}(\tau)={u_{i}}(0)+\sum\limits_{n=1}^{{M}}{{\alpha_{n}}\left\{{{x_{n,i}}(t)+\frac{{{u_{i}}(0)}}{{{\omega_{n}}}}\left[{\cos({\omega_{n}}t)-1}\right]}\right\}\sin\left[{{\omega_{n}}(t-\tau)-{\theta_{n}}}\right]},$ (58)
where $x_{n,i}(t)$ and $u_{i}(0)$ are respectively the $i^{\text{th}}$ elements of $\mathbf{x}_{n}(t)$ and $\mathbf{u}(0)$. Consider two approximates $\hat{\mathbf{u}}_{t,1}(\tau)$} and $\hat{\mathbf{u}}_{t,2}(\tau)$} of ${\mathbf{u}}_{1}(t)$ and ${\mathbf{u}}_{2}(t)$ $\in K$, respectively, with $\mathbf{x}_{1,n}(t)=\mathcal{L}_{t}\mathbf{u}_{1}(\omega_{n})$ and $\mathbf{x}_{2,n}(t)=\mathcal{L}_{t}\mathbf{u}_{2}(\omega_{n})$, from Eq. (58), the difference between the $i^{\text{th}}$ elements ${\hat{u}_{t,1,i}}(\tau)$ of $\hat{\mathbf{u}}_{t,1}(\tau)$} and ${\hat{u}_{t,2,i}}(\tau)$ of $\hat{\mathbf{u}}_{t,2}(\tau)$} can be bounded by
$\begin{array}[]{l}\mathop{\sup}\limits_{\tau\in[0,t]}\left|{{{\hat{u}}_{t,1,i}}(\tau)-{{\hat{u}}_{t,2,i}}(\tau)}\right|\\ \;\;\leq\left|{{u_{1,i}}(0)-{u_{2,i}}(0)}\right|\\ \;\;+\mathop{\sup}\limits_{\tau\in[0,t]}\left({\sum\limits_{n=1}^{M}{\left|{{\alpha_{n}}\left\{{\left[{{x_{1,n,i}}(t)-{x_{2,n,i}}(t)}\right]+\frac{{{u_{1,i}}(0)-{u_{2,i}}(0)}}{{{\omega_{n}}}}\left[{\cos({\omega_{n}}t)-1}\right]}\right\}\sin\left[{{\omega_{n}}(t-\tau)-{\theta_{n}}}\right]}\right|}}\right)\\ \;\;\leq\left|{{u_{1,i}}(0)-{u_{2,i}}(0)}\right|+\sum\limits_{n=1}^{M}{\left|{{\alpha_{n}}}\right|\left|{\left[{{x_{1,n,i}}(t)-{x_{2,n,i}}(t)}\right]+\frac{{{u_{1,i}}(0)-{u_{2,i}}(0)}}{{{\omega_{n}}}}\left[{\cos({\omega_{n}}t)-1}\right]}\right|}\\ \;\;\mathop{\leq}\limits^{(a)}\left|{{u_{1,i}}(0)-{u_{2,i}}(0)}\right|+\frac{1}{T}\sum\limits_{n=1}^{M}{\left[{\left|{{x_{1,n,i}}(t)-{x_{2,n,i}}(t)}\right|+2\left|{\frac{{{u_{1,i}}(0)-{u_{2,i}}(0)}}{{{\omega_{n}}}}}\right|}\right]}\\ \;\;\leq\left({1+\frac{{2M}}{{T{\omega_{1}}}}}\right)\left|{{u_{1,i}}(0)-{u_{2,i}}(0)}\right|+\frac{1}{T}\sum\limits_{n=1}^{M}{\left|{{x_{1,n,i}}(t)-{x_{2,n,i}}(t)}\right|}\\ \;\;\mathop{\leq}\limits^{(b)}\max\left[{\left({1+2M}\right),{T^{-1}}}\right]\left[{\left|{{u_{1,i}}(0)-{u_{2,i}}(0)}\right|+\sum\limits_{n=1}^{M}{\left|{{x_{1,n,i}}(t)-{x_{2,n,i}}(t)}\right|}}\right],\end{array}$ (59)
where $u_{1,i}(0)$ and $u_{2,i}(0)$ are respectively the $i^{\text{th}}$ elements of $\mathbf{u}_{1}(0)$ and $\mathbf{u}_{2}(0)$, $x_{1,n,i}(t)$ and $x_{2,n,i}(t)$ are respectively the $i^{\text{th}}$ elements of $\mathbf{x}_{1,n}(t)$ and $\mathbf{x}_{2,n}(t)$, Step ($a$) holds because $\alpha_{n}$ is bounded by $T^{-1}$ in Eq. (51) and Step ($b$) follows because $\omega_{n}=\pi n/(2T)$.
Subsequently, the first-order derivatives of the $j^{\text{th}}$ element $\psi_{j}\left(\mathbf{x},\mathbf{u}_{0},t\right)$ of $\Psi\left(\mathbf{x},\mathbf{u}_{0},t\right)$, j=1,2,,qj=1,2,...,q, are analyzed. With the modulus of continuity $\phi_{\Phi_{j}}(\cdot)$ of $\Phi_{j}$, which is the $j^{\text{th}}$ element the casual operator $\Phi$, and Eq. (59), the difference between $\psi_{j}\left(\mathbf{x},\mathbf{u}_{0},t\right)$ and $\psi_{j}\left(\mathbf{x}+\Delta\mathbf{x},\mathbf{u}_{0}+\Delta\mathbf{u}_{0},t\right)$ can be bounded by
$\begin{array}[]{l}\left|{{\psi_{j}}({\bf{x}}+\Delta{\bf{x}},{{\bf{u}}_{0}}+\Delta{{\bf{u}}_{0}},t)-{\psi_{j}}({\bf{x}},{{\bf{u}}_{0}},t)}\right|\\ \;\;\mathop{\leq}\limits^{(a)}{\phi_{{\Phi_{j}}}}\left\{{\left\lVert{\hat{\mathbf{u}}_{t,1}}(\tau)-{\hat{\mathbf{u}}_{t,2}}(\tau)\right\rVert}_{L^{\infty}[0,t]}\right\}={\phi_{{\Phi_{j}}}}\left({\mathop{\sup}\limits_{\tau\in[0,t]}\left[{\sum\limits_{i=1}^{p}{\left|{{{\hat{u}}_{t,1,i}}(\tau)-{{\hat{u}}_{t,2,i}}(\tau)}\right|}}\right]}\right)\\ \;\;\leq{\phi_{{\Phi_{j}}}}\left[{\sum\limits_{i=1}^{p}{\mathop{\sup}\limits_{\tau\in[0,t]}\left|{{{\hat{u}}_{t,1,i}}(\tau)-{{\hat{u}}_{t,2,i}}(\tau)}\right|}}\right]\leq{\phi_{{\Phi_{j}}}}\left\{{\max\left[{\left({1+2M}\right),{T^{-1}}}\right]\left({\left|{\Delta{\bf{x}}}\right|+\left|{\Delta{{\bf{u}}_{0}}}\right|}\right)}\right\},\end{array}$ (60)
where Step (a) holds from $\psi_{j}\left({\bf{x}},{\bf{u}}_{0},t\right)=\psi_{j}\left[{\bf{x}}(t),{\bf{u}}(0),t\right]=\Phi_{j}[{\hat{\mathbf{u}}}_{t}(\tau)](t)$. Thus, the first-order derivative of $\psi_{j}\left(\mathbf{x},\mathbf{u}_{0},t\right)$ with respect to each element of $\mathbf{x}$ or $\mathbf{u}_{0}$ is bounded by ${\max\left[{\left({1+2M}\right),{T^{-1}}}\right]}{{\phi}^{\prime}_{{\Phi_{j}}}}\left(0\right)$, where ${\phi}^{\prime}_{\Phi_{j}}(0)$ is the first-order derivative of $\phi_{\Phi_{j}}(\cdot)$ at zero.
Next, the first-order derivative of $\psi_{j}\left(\mathbf{x},\mathbf{u}_{0},t\right)$ with respect to $t$ is derived. The difference between $\psi_{j}\left(\mathbf{x},\mathbf{u}_{0},t\right)$ and $\psi_{j}\left(\mathbf{x},\mathbf{u}_{0},t+\Delta t\right)$ can be calculated as
${\psi_{j}}({\bf{x}},{{\bf{u}}_{0}},t+\Delta t)-{\psi_{j}}({\bf{x}},{{\bf{u}}_{0}},t)={\Phi_{j}}\left[{{{{\hat{\mathbf{u}}}}_{t+\Delta t}}(\tau)}\right](t+\Delta t)-{\Phi_{j}}\left[{{{{\hat{\mathbf{u}}}}_{t}}(\tau)}\right](t),$ (61)
where
${{\hat{\mathbf{u}}}_{t}}(\tau)={{\bf{u}}_{0}}+\sum\limits_{n=1}^{M}{{\alpha_{n}}\left\{{{{\bf{x}}_{n}}+\frac{{{{\bf{u}}_{0}}}}{{{\omega_{n}}}}\left[{\cos({\omega_{n}}t)-1}\right]}\right\}\sin\left[{{\omega_{n}}(t-\tau)-{\theta_{n}}}\right]}$ (62)
and
${{\hat{\mathbf{u}}}_{t+\Delta t}}(\tau)={{\bf{u}}_{0}}+\sum\limits_{n=1}^{M}{{\alpha_{n}}\left({{{\bf{x}}_{n}}+\frac{{{{\bf{u}}_{0}}}}{{{\omega_{n}}}}\left\{{\cos\left[{{\omega_{n}}(t+\Delta t)}\right]-1}\right\}}\right)\sin\left[{{\omega_{n}}(t+\Delta t-\tau)-{\theta_{n}}}\right]}.$ (63)
$\left|\psi_{j}\left(\mathbf{x},\mathbf{u}_{0},t+\Delta t\right)-\psi_{j}\left(\mathbf{x},\mathbf{u}_{0},t\right)\right|$ can be bounded by
$\begin{array}[]{l}\left|{{\psi_{j}}({\bf{x}},{{\bf{u}}_{0}},t+\Delta t)-{\psi_{j}}({\bf{x}},{{\bf{u}}_{0}},t)}\right|\\ \;\;\;\leq\left|{{\Phi_{j}}\left[{{{{\hat{\bf u}}}_{t+\Delta t}}(\tau)}\right](t+\Delta t)-{\Phi_{j}}\left[{{{{\hat{\mathbf{u}}}}_{t+\Delta t}}(\tau)}\right](t)}\right|+\left|{{\Phi_{j}}\left[{{{{\hat{\mathbf{u}}}}_{t+\Delta t}}(\tau)}\right](t)-{\Phi_{j}}\left[{{{{\hat{\mathbf{u}}}}_{t}}(\tau)}\right](t)}\right|\\ \;\;\;\leq\left|{\Delta t}\right|{D_{\Phi_{j}}}+\left|{{\Phi_{j}}\left[{{{{\hat{\mathbf{u}}}}_{t+\Delta t}}(\tau)}\right](t)-{\Phi_{j}}\left[{{{{\hat{\mathbf{u}}}}_{t}}(\tau)}\right](t)}\right|\\ \;\;\;\leq\left|{\Delta t}\right|{D_{\Phi_{j}}}+{\phi_{{\Phi_{j}}}}\left[{{{\left\|{{{{\hat{\mathbf{u}}}}_{t+\Delta t}}(\tau)-{{{\hat{\mathbf{u}}}}_{t}}(\tau)}\right\|}_{{L^{\infty}}\left[0,t\right]}}}\right],\end{array}$ (64)
where $D_{\Phi_{j}}$} is the bound of the first-order derivatives of all functions in $\Phi_{j}\left\{C_{0}([0,T];{\mathbb{R}^{p}})\right\}$ in Assumption 5 and it is independent of $M$.
Based Eqs. (62) and (63), the difference between ${\hat{\mathbf{u}}}_{t}(\tau)$ and ${{\hat{\mathbf{u}}}_{t+\Delta t}}(\tau)$, which is in the last term of Eq. (64), can be calculated as
$\begin{array}[]{l}{{{\hat{\mathbf{u}}}}_{t+\Delta t}}(\tau)-{{{\hat{\mathbf{u}}}}_{t}}(\tau)=\sum\limits_{n=1}^{M}{{\alpha_{n}}\left({{{\bf{x}}_{n}}+\frac{{{{\bf{u}}_{0}}}}{{{\omega_{n}}}}\left\{{\cos\left[{{\omega_{n}}(t+\Delta t)}\right]-1}\right\}}\right)\sin\left[{{\omega_{n}}(t+\Delta t-\tau)-{\theta_{n}}}\right]}\\ \;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;-\sum\limits_{n=1}^{M}{{\alpha_{n}}\left\{{{{\bf{x}}_{n}}+\frac{{{{\bf{u}}_{0}}}}{{{\omega_{n}}}}\left[{\cos({\omega_{n}}t)-1}\right]}\right\}\sin\left[{{\omega_{n}}(t-\tau)-{\theta_{n}}}\right]}\\ \;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;=\sum\limits_{n=1}^{M}{{\alpha_{n}}\left({{{\bf{x}}_{n}}-\frac{{{{\bf{u}}_{0}}}}{{{\omega_{n}}}}+\frac{{{{\bf{u}}_{0}}}}{{{\omega_{n}}}}\cos\left[{{\omega_{n}}(t+\Delta t)}\right]}\right)\sin\left[{{\omega_{n}}(t+\Delta t-\tau)-{\theta_{n}}}\right]}\\ \;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;-\sum\limits_{n=1}^{M}{{\alpha_{n}}\left[{{{\bf{x}}_{n}}-\frac{{{{\bf{u}}_{0}}}}{{{\omega_{n}}}}+\frac{{{{\bf{u}}_{0}}}}{{{\omega_{n}}}}\cos({\omega_{n}}t)}\right]\sin\left[{{\omega_{n}}(t-\tau)-{\theta_{n}}}\right]}\\ \;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;=\sum\limits_{n=1}^{M}{{\alpha_{n}}\left({{{\bf{x}}_{n}}-\frac{{{{\bf{u}}_{0}}}}{{{\omega_{n}}}}}\right){A_{n}}(\Delta t)}+\sum\limits_{n=1}^{M}{{\alpha_{n}}\frac{{{{\bf{u}}_{0}}}}{{{\omega_{n}}}}{B_{n}}(\Delta t)},\end{array}$ (65)
where
${A_{n}}(\Delta t)=\sin\left[{{\omega_{n}}(t+\Delta t-\tau)-{\theta_{n}}}\right]-\sin\left[{{\omega_{n}}(t-\tau)-{\theta_{n}}}\right]$ (66)
and
${B_{n}}(\Delta t)=\cos\left[{{\omega_{n}}(t+\Delta t)}\right]\sin\left[{{\omega_{n}}(t+\Delta t-\tau)-{\theta_{n}}}\right]-\cos({\omega_{n}}t)\sin\left[{{\omega_{n}}(t-\tau)-{\theta_{n}}}\right].$ (67)
Then, $\left|{\hat{\mathbf{u}}}_{t+\Delta t}(\tau)-{{\hat{\mathbf{u}}}_{t}}(\tau)\right|$ is bounded by
$\left|{{{{\hat{\mathbf{u}}}}_{t+\Delta t}}(\tau)-{{{\hat{\mathbf{u}}}}_{t}}(\tau)}\right|\leq\sum\limits_{n=1}^{M}{\left|{{\alpha_{n}}}\right|\left|{\left({{{\bf{x}}_{n}}-\frac{{{{\bf{u}}_{0}}}}{{{\omega_{n}}}}}\right)}\right|\left|{{A_{n}}(\Delta t)}\right|}+\sum\limits_{n=1}^{M}{\left|{{\alpha_{n}}}\right|\left|{\frac{{{{\bf{u}}_{0}}}}{{{\omega_{n}}}}}\right|\left|{{B_{n}}(\Delta t)}\right|}.$ (68)
Based on $|\sin(x)-\sin(y)|\leq|x-y|$, $A_{n}(\Delta t)$ in Eq. (66) is bounded by $\left|A_{n}(\Delta t)\right|\leq\left|\omega_{n}\Delta t\right|$ and the first sum on the right side of Eq. (68) can be bounded by
$\begin{array}[]{l}\sum\limits_{n=1}^{M}{\left|{{\alpha_{n}}}\right|\left|{\left({{{\bf{x}}_{n}}-\frac{{{{\bf{u}}_{0}}}}{{{\omega_{n}}}}}\right)}\right|\left|{{A_{n}}(\Delta t)}\right|}\\ \;\;\;\leq\sum\limits_{n=1}^{M}{\left|{{\alpha_{n}}}\right|\left|{\left({{{\bf{x}}_{n}}-\frac{{{{\bf{u}}_{0}}}}{{{\omega_{n}}}}}\right)}\right|\left|{{\omega_{n}}\Delta t}\right|}\\ \;\;\;\leq\left|{\Delta t}\right|\sum\limits_{n=1}^{M}{\left|{{\alpha_{n}}{\omega_{n}}{{\bf{x}}_{n}}-{\alpha_{n}}{{\bf{u}}_{0}}}\right|}\\ \;\;\;\leq\left|{\Delta t}\right|\sum\limits_{n=1}^{M}{\left|{{\alpha_{n}}{\omega_{n}}{{\bf{x}}_{n}}}\right|}+\left|{\Delta t}\right|\sum\limits_{n=1}^{M}{\left|{{\alpha_{n}}{{\bf{u}}_{0}}}\right|}\\ \;\;\;\leq\left|{\Delta t}\right|\sum\limits_{n=1}^{M}{\left|{{\alpha_{n}}}\right|\left|{{\omega_{n}}}\right|\left|{{{\bf{x}}_{n}}}\right|}+\left|{\Delta t}\right|p{B_{K}}M{T^{-1}}\\ \;\;\;\leq\left|{\Delta t}\right|pT{B_{K}}\sum\limits_{n=1}^{+\infty}{\left|{{\alpha_{n}}}\right|\left|{{\omega_{n}}}\right|}+\left|{\Delta t}\right|p{B_{K}}M{T^{-1}}\\ \;\;\;\mathop{\leq}\limits^{(a)}\left|{\Delta t}\right|pT{B_{K}}\frac{{{\pi^{2}}M{c_{\rho,3}}}}{{24T^{2}{c_{\rho,2}}}}+\left|{\Delta t}\right|p{B_{K}}M{T^{-1}}=\left|{\Delta t}\right|p{B_{K}}MT^{-1}\left({\frac{{{\pi^{2}}{c_{\rho,3}}}}{{24{c_{\rho,2}}}}+1}\right),\end{array}$ (69)
where Step (a) is from
$\begin{array}[]{l}\sum\limits_{n=1}^{+\infty}{\left|{{\omega_{n}}}\right|\left|{{\alpha_{n}}}\right|}=\frac{1}{T}\sum\limits_{n=1}^{+\infty}{\left|{\int_{-2T}^{2T}{{\omega_{n}}{e^{-{\rm{i}}{\omega_{n}}\tau}}{\rho_{v}}(\tau){\rm{d}}\tau}}\right|}\mathop{\leq}\limits^{(b)}\frac{1}{T}\sum\limits_{n=1}^{+\infty}{\frac{1}{{\omega_{n}^{2}}}\int_{-v}^{0}{\left|{{{\rho}^{\prime\prime\prime}_{v}}(\tau)}\right|{\rm{d}}\tau}}\\ \;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\leq\frac{1}{T}\sum\limits_{n=1}^{+\infty}{\frac{4{T^{2}}v}{{\pi^{2}}{n^{2}}}}\mathop{\max}\limits_{t\in[-v,0]}\left|{{{\rho}^{\prime\prime\prime}_{v}}(t)}\right|\leq\frac{{4T{c_{\rho,3}}}}{{{\pi^{2}}{v^{3}}}}\sum\limits_{n=1}^{+\infty}{\frac{1}{{{n^{2}}}}}\mathop{=}\limits^{(c)}\frac{{{\pi^{2}}M{c_{\rho,3}}}}{{24T^{2}{c_{\rho,2}}}},\end{array}$ (70)
Step (b) is derived using the same method as that in Eq. (45), Step (c) holds from Eq. (14) and $\sum_{n=1}^{+\infty}{n^{-2}}=\pi^{2}/6$, and $c_{\rho,3}=\max_{\tau\in[-1,0]}\left|{\rho}_{1}^{\prime\prime\prime}(\tau)\right|$ is an independent constant with ${\rho}_{1}(\tau)$ in Eq. (41).
$\left|B_{N}(\Delta t)\right|$ in Eq. (67) can be bounded by
$\begin{array}[]{l}\left|{{B_{n}}(\Delta t)}\right|\\ \;\;\;\leq\left|{\cos\left[{{\omega_{n}}(t+\Delta t)}\right]\sin\left[{{\omega_{n}}(t+\Delta t-\tau)-{\theta_{n}}}\right]-\cos\left[{{\omega_{n}}(t+\Delta t)}\right]\sin\left[{{\omega_{n}}(t-\tau)-{\theta_{n}}}\right]}\right|\\ \;\;\;\;\;\;\;+\left|{\cos\left[{{\omega_{n}}(t+\Delta t)}\right]\sin\left[{{\omega_{n}}(t-\tau)-{\theta_{n}}}\right]-\cos({\omega_{n}}t)\sin\left[{{\omega_{n}}(t-\tau)-{\theta_{n}}}\right]}\right|\leq 2\left|{{\omega_{n}}}\right|\left|{\Delta t}\right|.\end{array}$ (71)
Thus, the second sum on the right side of Eq. (68) can be simplified as
$\sum\limits_{n=1}^{M}{\left|{{\alpha_{n}}}\right|\left|{\frac{{{{\bf{u}}_{0}}}}{{{\omega_{n}}}}}\right|\left|{{B_{n}}(\Delta t)}\right|}\leq 2\sum\limits_{n=1}^{M}{\left|{{\alpha_{n}}}\right|\left|{\frac{{{{\bf{u}}_{0}}}}{{{\omega_{n}}}}}\right|\left|{{\omega_{n}}}\right|\left|{\Delta t}\right|}\leq\frac{{2p{B_{K}}M}}{T}\left|{\Delta t}\right|.$ (72)
By substituting Eqs. (69) and (72) into Eq. (68), $\left|{\hat{\mathbf{u}}}_{t+\Delta t}(\tau)-{{\hat{\mathbf{u}}}_{t}}(\tau)\right|$ is bounded by
$\left|{{{{\bf{\hat{u}}}}_{t+\Delta t}}(\tau)-{{{\bf{\hat{u}}}}_{t}}(\tau)}\right|\leq p\left|{\Delta t}\right|{B_{K}}MT^{-1}\left({\frac{{{\pi^{2}}{c_{\rho,3}}}}{{24{c_{\rho,2}}}}+3}\right).$ (73)
By substituting Eq. (73) into Eq. (64), $\left|\psi_{j}\left(\mathbf{x},\mathbf{u}_{0},t+\Delta t\right)-\psi_{j}\left(\mathbf{x},\mathbf{u}_{0},t\right)\right|$ can be bounded by
$\left|{{\psi_{j}}({\bf{x}},{{\bf{u}}_{0}},t+\Delta t)-{\psi_{j}}({\bf{x}},{{\bf{u}}_{0}},t)}\right|\leq\left|{\Delta t}\right|{D_{{\Phi_{j}}}}+{\phi_{{\Phi_{j}}}}\left(p{\left|{\Delta t}\right|{B_{K}}MT^{-1}c_{L}}\right),$ (74)
where $c_{L}=24^{-1}{\pi^{2}}{c_{\rho,3}}{c_{\rho,2}^{-1}}+3$ is an independent constant. The first-order derivative of $\psi_{j}\left(\mathbf{x},\mathbf{u}_{0},t\right)$ with respect to t is bounded by
$\begin{array}[]{l}\left|{\frac{{{\partial}{\psi_{j}}({\mathbf{x}},{{\bf{u}}_{0}},t)}}{{{\partial}t}}}\right|\leq{D_{{\Phi_{j}}}}+\mathop{\lim}\limits_{\left|{\Delta t}\right|\to 0}{\left|{\Delta t}\right|^{-1}}{\phi_{{\Phi_{j}}}}\left({p\left|{\Delta t}\right|{B_{K}}{M}T^{-1}c_{L}}\right)={D_{{\Phi_{j}}}}+p{B_{K}}MT^{-1}c_{L}{{\phi^{\prime}}_{{\Phi_{j}}}}(0).\end{array}$ (75)
Eqs. (60) and (75) indicate that for every $M$ satisfying Eq. (15) along with $v$ satisfying Eq. (14), over the compact domain $\Omega_{M}\left(TB_{K},B_{K},T\right)=\left[-TB_{K},TB_{K}\right]^{pM}\times\left[-B_{K},B_{K}\right]^{p}\times\left[0,T\right]$, the Lipschitz constant of $\psi_{j}\left(\mathbf{x},\mathbf{u}_{0},t\right)$ with respect to its each argument is $L_{\psi_{j}}\left(M\right)$ in Eq. (17). The linear modulus of continuity $\phi_{\psi_{j},M}(x)$ in Eq. (18) of $\psi_{j}\left(\mathbf{x},\mathbf{u}_{0},t\right)$ under $L^{2}$-norm is proven by
$\begin{array}[]{l}\left|{{\psi_{j}}({\bf{x}}+\Delta{\bf{x}},{{\bf{u}}_{0}}+\Delta{{\bf{u}}_{0}},t+\Delta t)-{\psi_{j}}({\bf{x}},{{\bf{u}}_{0}},t)}\right|\\ \;\;\;\leq\left|{{\psi_{j}}({\bf{x}}+\Delta{\bf{x}},{{\bf{u}}_{0}}+\Delta{{\bf{u}}_{0}},t+\Delta t)-{\psi_{j}}({\bf{x}},{{\bf{u}}_{0}}+\Delta{{\bf{u}}_{0}},t+\Delta t)}\right|\\ \;\;\;\;\;\;\;+\left|{{\psi_{j}}({\bf{x}},{{\bf{u}}_{0}}+\Delta{{\bf{u}}_{0}},t+\Delta t)-{\psi_{j}}({\bf{x}},{{\bf{u}}_{0}},t+\Delta t)}\right|+\left|{{\psi_{j}}({\bf{x}},{{\bf{u}}_{0}},t+\Delta t)-{\psi_{j}}({\bf{x}},{{\bf{u}}_{0}},t)}\right|\\ \;\;\;\leq{L_{{\psi_{j}}}\left(M\right)}\left({\left|{\Delta{\bf{x}}}\right|+\left|{\Delta{{\bf{u}}_{0}}}\right|+\left|{\Delta t}\right|}\right)\mathop{\leq}\limits^{(a)}\sqrt{p(M+1)+1}{L_{{\psi_{j}}}\left(M\right)}\sqrt{\left|{\Delta{\bf{x}}}\right|_{2}^{2}+\left|{\Delta{{\bf{u}}_{0}}}\right|_{2}^{2}+\left|{\Delta t}\right|_{2}^{2}},\end{array}$ (76)
where Step (a) holds from the Cauchy-Schwarz inequality. $\blacksquare$
A.5 Proof of Theorem 1
Proof: From Remark 2 and Lemma 2, for an arbitrary positive integer $M_{\it{\Gamma}}$ satisfying Eq. (15) along with v satisfying Eq. (14), there exists a one-hidden-layer MLP ${\it{\Gamma}}(\cdot)$ specified in Lemma 2, for every $\mathbf{u}(t)\in K$, its corresponding solution $\mathbf{x}(t)=[\mathbf{x}_{1}^{\top}(t),\mathbf{x}_{2}^{\top}(t),...,\mathbf{x}_{M_{\it{\Gamma}}}^{\top}(t)]^{\top}$ calculated from the second-order ODE in Eq. (1) with the initial conditions $\mathbf{x}(0)={\mathbf{x}}^{\prime}(0)=\mathbf{0}$ equal the sine transform coefficients of $\mathbf{u}(t)$, as shown in Eq. (55). From Lemma 3 and Remark 3, there exists a continuous mapping $\Psi=[\psi_{1},\psi_{2},...,\psi_{q}]^{\top}:\mathbb{R}^{p\left(M_{\it{\Gamma}}+1\right)}\times\left[0,T\right]\to\mathbb{R}^{q}$ for all $\mathbf{u}(t)\in K$, the difference between $\Phi[\mathbf{u}(\tau)](t)$ and $\Psi[\mathbf{x}(t),\mathbf{u}(0),t]$ can be bounded by
$\left|{\Phi\left[{{\bf{u}}(\tau)}\right](t)-\Psi\left[{{\bf{x}}(t),{\bf{u}}(0),t}\right]}\right|\leq{\phi_{\Phi}}\left(\frac{{54^{1/3}p{L_{K}}Tc_{\rho,2}^{{1\mathord{\left/{\vphantom{13}}\right.\kern-1.2pt}3}}}}{{{\pi^{{2\mathord{\left/{\vphantom{23}}\right.\kern-1.2pt}3}}}{M_{{\it{\Gamma}}}^{{1\mathord{\left/{\vphantom{13}}\right.\kern-1.2pt}3}}}}}\right).$ (77)
From Lemma 4, the continuous function $\Psi\left({\bf{x}},{\bf{u}}_{0},t\right)$ in Eq. (77) is compactly supported within $\Omega_{M_{\it{\Gamma}}}\left(TB_{K},B_{K},T\right)=\left[-TB_{K},TB_{K}\right]^{pM_{\it{\Gamma}}}\times\left[-B_{K},B_{K}\right]^{p}\times\left[0,T\right]$ for every integer $M_{\it{\Gamma}}$ satisfying Eq. (15) and all $\mathbf{u}(t)\in K$, where $\mathbf{x}$ and $\mathbf{u}_{0}$ represent the values of $\mathbf{x}(t)$ and $\mathbf{u}(0)$, respectively. The Lipschtiz constant $L_{{\psi_{j}}}\left(M_{\it{\Gamma}}\right)$ of $\psi_{j}\left(\mathbf{x},\mathbf{u}_{0},t\right)$ of $\Psi\left({\bf{x}},{\bf{u}}_{0},t\right)$ with respect to its each argument is in Eq. (17) and a linear modulus of continuity $\phi_{\psi_{j},M_{{\it{\Gamma}}}}(x)$ of $\psi_{j}\left(\mathbf{x},\mathbf{u}_{0},t\right)$ under $L^{2}$-norm is in Eq. (18).
Theorem 1 by Hanin (2019) demonstrates that given an arbitrary continuous function $f\left(\mathbf{z}\right)$, $\mathbf{z}=\left[z_{1},z_{2},...,z_{d}\right]^{\top}$, mapping from $0,1]^{d}$ to $[0,1]$, with a modulus of continuity $w_{f}(z)$ of $f\left(\mathbf{z}\right)$, for an arbitrary error $\varepsilon>0$, there exists an MLP $f_{N}\left(\mathbf{z}\right)$ with $\sigma_{\mathrm{ReLU}}(\cdot)$, input dimension $d$, hidden layer width $d+3$, output dimension 1, and depth $2d!\left[\sqrt{d}/w_{f}^{-1}(\varepsilon)\right]^{d}$, where $w_{f}^{-1}(\varepsilon)$ is the inverse function of $w_{f}\left(z\right)$, such that $\sup_{\mathbf{z}\in\left[0,1\right]^{d}}\left|f\left(\mathbf{z}\right)-f_{N}\left(\mathbf{z}\right)\right|\leq\varepsilon$. Theorem 1 by Hanin and Sellke (2017) and Proposition 53 by Kratsios and Papon (2022) provide similar results.
Using Theorem 1 by Hanin (2019) and the linear modulus of continuity $\phi_{\psi_{j},M_{{\it{\Gamma}}}}(x)$ in Eq. (18), for every positive integer $H_{\it{\Pi}}$, there exists an MLP ${\it{\Pi}}_{j}(\cdot)$ with $\sigma_{\mathrm{ReLU}}(\cdot)$, the width and depth of ${\it{\Pi}}_{j}(\cdot)$ are respectively $p\left(M_{\it{\Gamma}}+1\right)+4$ and $H_{\it{\Pi}}+1$, such that the difference between $\psi_{j}\left[{\bf{x}}(t),{\bf{u}}(0),t\right]$ and ${\it{\Pi}}_{j}\left[{\bf{x}}(t),{\bf{u}}(0),t\right]$ over $\Omega_{M_{\it{\Gamma}}}\left(TB_{K},B_{K},T\right)$ is bounded by
[…]
for all $\mathbf{u}(t)\in K$, where $B_{\Omega}=2\max\left(TB_{K},B_{K},T\right)$. By applying Eq. (78) to all $\psi_{j}\left(\mathbf{x},\mathbf{u}_{0},t\right)$, j=1,2,,qj=1,2,...,q, there exists an MLP ${\it{\Pi}}(\cdot)$ consisting of ${\it{\Pi}}_{1}(\cdot),{\it{\Pi}}_{2}(\cdot),...,{\it{\Pi}}_{q}(\cdot)$, the width of ${\it{\Pi}}(\cdot)$ is in Eq. (22) and its depth is $H_{\it{\Pi}}+1$. The difference between $\Psi\left[{\bf{x}}(t),{\bf{u}}(0),t\right]$ and ${\it{\Pi}}\left[{\bf{x}}(t),{\bf{u}}(0),t\right]$ over $\Omega_{M_{\it{\Gamma}}}\left(TB_{K},B_{K},T\right)$ is bounded by
[…]
for all $\mathbf{u}(t)\in K$. Eq. (19) can be proven by combining Eq. (77) and Eq. (79) and noting that $\mathbf{y}(t)={\it{\Pi}}\left[{\bf{x}}(t),{\bf{u}}(0),t\right]$ in Eq. (1). This completes the proof. $\blacksquare$
Appendix B Proofs for Lemmas and Theorem 2 in Section 3.2
B.1 Proof of Lemma 5
Proof: Since $g_{1}(\cdot)$ and $g_{2}(\cdot)$ in Eq. (23) are Lipschitz continuous over $\mathbb{R}^{2r+p}$ and the values of all $\mathbf{u}(t)\in K$ are within $\Omega_{\mathbf{u}}$, e.g., $\Omega_{\mathbf{u}}=\left[-B_{K},B_{K}\right]^{p}$, the solution ${\mathbf{z}}_{i}\left(t,{\mathbf{z}}_{\tau}\right)$ uniquely exists and its value is bounded for all $\mathbf{u}(t)\in K$, all ${\mathbf{z}}_{\tau}\in\Omega_{\mathbf{z}}$, all $t\in[0,T]$, all $0\leq\tau\leq t$, and i = 1 and 2. Then, the ODEs in Eq. (23) governed by $g_{1}(\cdot)$ or $g_{2}(\cdot)$, driven by the same arbitrary $\mathbf{u}(t)\in K$ with the same initial condition ${\mathbf{z}}_{0}$, have unique solutions ${\mathbf{z}}_{1}\left(t,{\mathbf{z}}_{0}\right)$ and ${\mathbf{z}}_{2}\left(t,{\mathbf{z}}_{0}\right)$, respectively. ${\mathbf{z}}_{2}\left(t,{\mathbf{z}}_{0}\right)-{\mathbf{z}}_{1}\left(t,{\mathbf{z}}_{0}\right)$ can be calculated as
$\begin{array}[]{l}{{\bf{z}}_{2}}(t,{{\bf{z}}_{0}})-{{\bf{z}}_{1}}(t,{{\bf{z}}_{0}})\\ \;\;\;\;\;=\int_{0}^{t}{\frac{{{\rm{d}}{{\bf{z}}_{1}}\left[{t,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{{\rm{d}}\tau}}{\rm{d}}\tau}=\int_{0}^{t}{\frac{{\rm{d}}}{{{\rm{d}}\tau}}\left\{{\int_{\tau}^{t}{{{{\bf{z}}}_{1}^{\prime}}\left[{\lambda,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]{\rm{d}}\lambda}+{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right\}{\rm{d}}\tau}\\ \;\;\;\;\;=\int_{0}^{t}{\left\{{\int_{\tau}^{t}{\frac{{\partial{{{\bf{z}}}_{1}^{\prime}}\left[{\lambda,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial\tau}}{\rm{d}}\lambda}+\frac{{\partial{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}}{{\partial\tau}}-{{\left.{\frac{{\partial{{\bf{z}}_{1}}\left[{\lambda,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial\lambda}}}\right|}_{\lambda=\tau}}}\right\}{\rm{d}}\tau}\\ \;\;\;\;\;=\int_{0}^{t}{\left({\left\{{\int_{\tau}^{t}{\frac{{\partial{{{\bf{z}}}_{1}^{\prime}}\left[{\lambda,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}}{\rm{d}}\lambda}+{\bf{I}}}\right\}\frac{{\partial{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}}{{\partial\tau}}-{{\left.{\frac{{\partial{{\bf{z}}_{1}}\left[{\lambda,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial\lambda}}}\right|}_{\lambda=\tau}}}\right){\rm{d}}\tau}\\ \;\;\;\;\;=\int_{0}^{t}{{\left\{{\int_{\tau}^{t}{\frac{{\partial{{{\bf{z}}}_{1}^{\prime}}\left[{\lambda,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}}{\rm{d}}\lambda}+{\bf{I}}}\right\}\par{\begin{Bmatrix}{{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}})}\\ {{g_{2}}\left[{{{\bf{z}}_{2,1}}(\tau,{{\bf{z}}_{0}}),{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}}),{\bf{u}}(\tau)}\right]}\end{Bmatrix}}}{\rm{d}}\tau}-\int_{0}^{t}{{{{\left.{\frac{{\partial{{\bf{z}}_{1}}\left[{\lambda,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial\lambda}}}\right|}_{\lambda=\tau}}}{\rm{d}}\tau},\end{array}$ (80)
where $\mathbf{I}$ is an identity matrix.
Given an arbitrary initial condition ${\mathbf{z}}_{\tau}\in\Omega_{\mathbf{z}}$ at an arbitrary initial time instant $\tau$ ($0\leq\tau\leq t$), it can be obtained
[…]
Since the values of $\tau$ and ${\mathbf{z}}_{\tau}\in{\Omega}_{\mathbf{z}}$ in Eq. (81) are arbitrary, the value of ${\mathbf{z}}_{1}\left({\hat{\tau}},{\mathbf{z}}_{\tau}\right)={\mathbf{z}}_{\tau}$ at the arbitrary $\hat{\tau}=\tau$ is also arbitrary within ${\Omega}_{\mathbf{z}}$. In addition, ${\mathbf{z}}_{2}\left({\hat{\tau}},{\mathbf{z}}_{0}\right)\in{\Omega}_{\mathbf{z}}$ for all $\hat{\tau}\in[0,T]$. It is able to replace ${\mathbf{z}}_{1}\left({\hat{\tau}},{\mathbf{z}}_{\tau}\right)$ in Eq. (81) with ${\mathbf{z}}_{2}\left({\hat{\tau}},{\mathbf{z}}_{0}\right)$ to obtain
${\left.{\frac{{\partial{{\bf{z}}_{1}}\left[{\lambda,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial\lambda}}}\right|_{\lambda=\tau}}=\left\{{\int_{\tau}^{t}{\frac{{\partial{{{\bf{z^{\prime}}}}_{1}}\left[{\lambda,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}}{\rm{d}}\lambda}+{\bf{I}}}\right\}{\begin{Bmatrix}{{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}})}\\ {{g_{1}}\left[{{{\bf{z}}_{2,1}}(\tau,{{\bf{z}}_{0}}),{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}}),{\bf{u}}(\tau)}\right]}\end{Bmatrix}}.$ (82)
Substituting Eq. (82) in Eq. (80), ${\mathbf{z}}_{2}\left(t,{\mathbf{z}}_{0}\right)-{\mathbf{z}}_{1}\left(t,{\mathbf{z}}_{0}\right)$ can be calculated as
$\begin{array}[]{l}{{\bf{z}}_{2}}(t,{{\bf{z}}_{0}})-{{\bf{z}}_{1}}(t,{{\bf{z}}_{0}})\\ \;\;=\int_{0}^{t}{\left\{{\int_{\tau}^{t}{\frac{{\partial{{{\bf{z}}}_{1}^{\prime}}\left[{\lambda,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}}{\rm{d}}\lambda}+{\bf{I}}}\right\}}{{\begin{Bmatrix}{\bf{0}}\\ {{g_{2}}\left[{{{\bf{z}}_{2,1}}(\tau,{{\bf{z}}_{0}}),{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}}),{\bf{u}}(\tau)}\right]-{g_{1}}\left[{{{\bf{z}}_{2,1}}(\tau,{{\bf{z}}_{0}}),{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}}),{\bf{u}}(\tau)}\right]}\end{Bmatrix}}}\rm{d}\tau\\ \;\;\mathop{=}\limits^{(a)}\int_{0}^{t}{\frac{{\partial{{\bf{z}}_{1}}\left[{t,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}}{\begin{Bmatrix}{\bf{0}}\\ {{g_{2}}\left[{{{\bf{z}}_{2,1}}(\tau,{{\bf{z}}_{0}}),{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}}),{\bf{u}}(\tau)}\right]-{g_{1}}\left[{{{\bf{z}}_{2,1}}(\tau,{{\bf{z}}_{0}}),{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}}),{\bf{u}}(\tau)}\right]}\end{Bmatrix}}{\rm{d}}\tau}\\ \;\;=\int_{0}^{t}{\frac{{\partial{{\bf{z}}_{1}}\left[{t,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}})}}\left\{{{g_{2}}\left[{{{\bf{z}}_{2,1}}(\tau,{{\bf{z}}_{0}}),{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}}),{\bf{u}}(\tau)}\right]-{g_{1}}\left[{{{\bf{z}}_{2,1}}(\tau,{{\bf{z}}_{0}}),{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}}),{\bf{u}}(\tau)}\right]}\right\}{\rm{d}}\tau},\end{array}$ (83)
where Step (a) holds because ${\mathbf{z}}_{1}\left[\tau,{\mathbf{z}}_{2}\left(\tau,{\mathbf{z}}_{0}\right)\right]={\mathbf{z}}_{2}\left(\tau,{\mathbf{z}}_{0}\right)$. This completes the proof. $\blacksquare$
B.2 Proof of Lemma 6
Proof: Considering Eq. (26) in Lemma 5 and let $\mathbf{Y}(t)=\partial{\mathbf{z}}_{1}\left[t,{\mathbf{z}}_{2}\left(\tau,{\mathbf{z}}_{0}\right)\right]/\partial{\mathbf{z}}_{2,2}\left(\tau,{\mathbf{z}}_{0}\right)$, $0\leq\tau\leq t$, it is satisfied that
$\frac{{{\rm{d}}{\bf{Y}}(t)}}{{{\rm{d}}t}}=\frac{{\partial{{{\bf{z}}}_{1}^{\prime}}\left[{t,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}})}}=\frac{{\partial{{{\bf{z}}}_{1}^{\prime}}\left[{t,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{1}}\left[{t,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}\frac{{\partial{{\bf{z}}_{1}}\left[{t,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}})}}=\frac{{\partial{{{\bf{z}}}_{1}^{\prime}}\left[{t,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{1}}\left[{t,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{\bf{Y}}(t),$ (84)
with the initial condition $\mathbf{Y}(\tau)=\left[{\mathbf{0}}_{r\times r},{\mathbf{I}}_{r\times r}\right]^{\top}$, where ${\mathbf{0}}_{r\times r}$ and ${\mathbf{I}}_{r\times r}$ denote the zero and identity matrices, respectively, and
[…]
Then, $\mathbf{Y}(t)$ can be expressed as
${\bf{Y}}(t)=\exp\left\{{\int_{\tau}^{t}{\frac{{\partial{{{\bf{z}}}_{1}^{\prime}}\left[{\hat{\tau},{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{1}}\left[{\hat{\tau},{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{\rm{d}}\hat{\tau}}}\right\}{\begin{bmatrix}{{{\bf{0}}_{r\times r}}}\\ {{{\bf{I}}_{r\times r}}}\end{bmatrix}}$, (86)
where
$\exp\left\{{\int_{\tau}^{t}{\frac{{\partial{{{\bf{z}}}_{1}^{\prime}}\left[{\hat{\tau},{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{1}}\left[{\hat{\tau},{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{\rm{d}}\hat{\tau}}}\right\}=\mathop{\lim}\limits_{\scriptstyle\Delta\hat{\tau}\to 0\atop\scriptstyle n\to+\infty}\prod\limits_{i=1}^{n}{\exp\left\{{\frac{{\partial{{{\bf{z}}}_{1}^{\prime}}\left[{{{\hat{\tau}}_{i}},{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{1}}\left[{{{\hat{\tau}}_{i}},{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}\Delta\hat{\tau}}\right\}}$, (87)
$\tau=\hat{\tau}_{0}<\hat{\tau}_{1}<\hat{\tau}_{2}<...<\hat{\tau}_{n}=t$, $\Delta\hat{\tau}=\hat{\tau}_{i}-\hat{\tau}_{i-1}$, i=1,2,,ni=1,2,...,n, and
$\exp\left\{{\frac{{\partial{{{\bf{z}}}_{1}^{\prime}}\left[{{{\hat{\tau}}_{i}},{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{1}}\left[{{{\hat{\tau}}_{i}},{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}\Delta\hat{\tau}}\right\}=\sum\limits_{m=0}^{+\infty}{\frac{1}{{m!}}}{\left\{{\frac{{\partial{{{\bf{z}}}_{1}^{\prime}}\left[{{{\hat{\tau}}_{i}},{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{1}}\left[{{{\hat{\tau}}_{i}},{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}\Delta\hat{\tau}}\right\}^{m}}.$ (88)
Substituting Eq. (86) into Eq. (26), ${\mathbf{z}}_{2}\left(t,{\mathbf{z}}_{0}\right)-{\mathbf{z}}_{1}\left(t,{\mathbf{z}}_{0}\right)$ can be calculated as
${{\bf{z}}_{2}}(t,{{\bf{z}}_{0}})-{{\bf{z}}_{1}}(t,{{\bf{z}}_{0}})=\int_{0}^{t}{\exp\left\{{\int_{\tau}^{t}{\frac{{\partial{{{\bf{z}}}_{1}^{\prime}}\left[{\hat{\tau},{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{1}}\left[{\hat{\tau},{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{\rm{d}}\hat{\tau}}}\right\}{\bf{\Delta}}g(\tau,{{\bf{z}}_{0}}){\rm{d}}\tau},$ (89)
where
${\bf{\Delta}}g(\tau,{{\bf{z}}_{0}})={\begin{Bmatrix}{{{\bf{0}}_{r\times 1}}}\\ {{g_{2}}\left[{{{\bf{z}}_{2,1}}(\tau,{{\bf{z}}_{0}}),{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}}),{\bf{u}}(\tau)}\right]-{g_{1}}\left[{{{\bf{z}}_{2,1}}(\tau,{{\bf{z}}_{0}}),{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}}),{\bf{u}}(\tau)}\right]}\end{Bmatrix}}.$ (90)
Since $g_{1}(\cdot)$ is Lipschitz continuous, from Eq. (85), it can be obtained
[…]
where $L_{g_{1}}=\max\left[L_{g_{1},1},\left(L_{g_{1},2}+1\right)\right]$, $L_{g_{1},1}$ and $L_{g_{1},2}$ are respectively the two Lipschitz constants of $g_{1}(\cdot)$ with respect to its first and second arguments satisfying
$\left|{{g_{1}}\left({{{\bf{b}}_{1}},{{\bf{z}}_{1,2}}\left[{t,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right],{\bf{u}}(t)}\right)-{g_{1}}\left({{{\bf{b}}_{2}},{{\bf{z}}_{1,2}}\left[{t,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right],{\bf{u}}(t)}\right)}\right|\leq{L_{{g_{1}},1}}\left|{{{\bf{b}}_{1}}-{{\bf{b}}_{2}}}\right|,$ (92)
$\left|{{g_{1}}\left({{{\bf{z}}_{1,1}}\left[{t,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right],{{\bf{b}}_{1}},{\bf{u}}(t)}\right)-{g_{1}}\left({{{\bf{z}}_{1,1}}\left[{t,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right],{{\bf{b}}_{2}},{\bf{u}}(t)}\right)}\right|\leq{L_{{g_{1}},2}}\left|{{{\bf{b}}_{1}}-{{\bf{b}}_{2}}}\right|,$ (93)
and $\mathbf{a}_{1}$, $\mathbf{a}_{2}$, $\mathbf{b}_{1}$, and $\mathbf{b}_{2}$ are arbitrary r-dimensional vectors.
Utilizing Eqs. (87)-(91), $\left|{\mathbf{z}}_{2}\left(t,{\mathbf{z}}_{0}\right)-{\mathbf{z}}_{1}\left(t,{\mathbf{z}}_{0}\right)\right|$ can be bounded by
$\begin{array}[]{l}\left|{{{\bf{z}}_{2}}(t,{{\bf{z}}_{0}})-{{\bf{z}}_{1}}(t,{{\bf{z}}_{0}})}\right|\\ \;\;\leq\int_{0}^{t}{\left|{\exp\left\{{\int_{\tau}^{t}{\frac{{\partial{{{\bf{z}}}_{1}^{\prime}}\left[{\hat{\tau},{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{1}}\left[{\hat{\tau},{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{\rm{d}}\hat{\tau}}}\right\}{\bf{\Delta}}g(\tau,{{\bf{z}}_{0}})}\right|{\rm{d}}\tau}=\int_{0}^{t}{\left|{\mathop{\lim}\limits_{\scriptstyle\Delta\hat{\tau}\to 0\atop\scriptstyle n\to+\infty}\prod\limits_{i=1}^{n}{\exp\left\{{\frac{{\partial{{{\bf{z}}}_{1}^{\prime}}\left[{{{\hat{\tau}}_{i}},{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{1}}\left[{{{\hat{\tau}}_{i}},{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}\Delta\hat{\tau}}\right\}}{\bf{\Delta}}g(\tau,{{\bf{z}}_{0}})}\right|{\rm{d}}\tau}\\ \;\;\leq\int_{0}^{t}{\mathop{\lim}\limits_{\scriptstyle\Delta\hat{\tau}\to 0\atop\scriptstyle n\to+\infty}\left|{\prod\limits_{i=1}^{n}{\left({\sum\limits_{m=0}^{+\infty}{\frac{1}{{m!}}}{{\left\{{\frac{{\partial{{{\bf{z}}}_{1}^{\prime}}\left[{{{\hat{\tau}}_{i}},{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{1}}\left[{{{\hat{\tau}}_{i}},{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}\Delta\hat{\tau}}\right\}}^{m}}}\right)}{\bf{\Delta}}g(\tau,{{\bf{z}}_{0}})}\right|{\rm{d}}\tau}\\ \;\;=\int_{0}^{t}{\mathop{\lim}\limits_{\scriptstyle\Delta\hat{\tau}\to 0\atop\scriptstyle n\to+\infty}\left|{\left({\sum\limits_{m=0}^{+\infty}{\frac{1}{{m!}}}{{\left\{{\frac{{\partial{{{\bf{z}}}_{1}^{\prime}}\left[{{{\hat{\tau}}_{n}},{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{1}}\left[{{{\hat{\tau}}_{n}},{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}\Delta\hat{\tau}}\right\}}^{m}}}\right)\prod\limits_{i=1}^{n-1}{\left({\sum\limits_{m=0}^{+\infty}{\frac{1}{{m!}}}{{\left\{{\frac{{\partial{{{\bf{z}}}_{1}^{\prime}}\left[{{{\hat{\tau}}_{i}},{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{1}}\left[{{{\hat{\tau}}_{i}},{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}\Delta\hat{\tau}}\right\}}^{m}}}\right)}{\bf{\Delta}}g(\tau,{{\bf{z}}_{0}})}\right|{\rm{d}}\tau}\\ \;\;\mathop{\leq}\limits^{(a)}\int_{0}^{t}{\mathop{\lim}\limits_{\scriptstyle\Delta\hat{\tau}\to 0\atop\scriptstyle n\to+\infty}\left[{\sum\limits_{m=0}^{+\infty}{\frac{1}{{m!}}}{{\left({{L_{{g_{1}}}}\Delta\hat{\tau}}\right)}^{m}}}\right]\left|{\prod\limits_{i=1}^{n-1}{\left({\sum\limits_{m=0}^{+\infty}{\frac{1}{{m!}}}{{\left\{{\frac{{\partial{{{\bf{z}}}_{1}^{\prime}}\left[{{{\hat{\tau}}_{i}},{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{1}}\left[{{{\hat{\tau}}_{i}},{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}\Delta\hat{\tau}}\right\}}^{m}}}\right)}{\bf{\Delta}}g(\tau,{{\bf{z}}_{0}})}\right|{\rm{d}}\tau}\\ \;\;\leq\int_{0}^{t}{\mathop{\lim}\limits_{\scriptstyle\Delta\hat{\tau}\to 0\atop\scriptstyle n\to+\infty}{{\left[{\sum\limits_{m=0}^{+\infty}{\frac{1}{{m!}}}{{\left({{L_{{g_{1}}}}\Delta\hat{\tau}}\right)}^{m}}}\right]}^{n}}\left|{{\bf{\Delta}}g(\tau,{{\bf{z}}_{0}})}\right|{\rm{d}}\tau}=\int_{0}^{t}{\exp\left({\int_{\tau}^{t}{{L_{{g_{1}}}}{\rm{d}}\hat{\tau}}}\right)\left|{{\bf{\Delta}}g(\tau,{{\bf{z}}_{0}})}\right|{\rm{d}}\tau}\\ \;\;\leq{e^{T{L_{{g_{1}}}}}}T\mathop{\max}\limits_{\tau\in[0,T]}\left|{{\bf{\Delta}}g(\tau,{{\bf{z}}_{0}})}\right|,\end{array}$ (94)
where Step (a) holds from Eq. (91). Substituting Eq. (90) into Eq. (94), Eq. (27) can be proven.
If the dynamical system governed by $g_{1}(\cdot)$ is uniformly asymptotically incrementally stable for all $\mathbf{u}(t)\in K$ on the domain $\Omega_{\mathbf{z}}$ in Lemma 5 with a set of stability functions $\beta_{k}\left(h,t\right)$, k=1,2,,2rk=1,2,...,2r, then, from Eq. (26), $\left|{\mathbf{z}}_{2}\left(t,{\mathbf{z}}_{0}\right)-{\mathbf{z}}_{1}\left(t,{\mathbf{z}}_{0}\right)\right|$ can be bounded by
$\begin{array}[]{l}\left|{{{\bf{z}}_{2}}(t,{{\bf{z}}_{0}})-{{\bf{z}}_{1}}(t,{{\bf{z}}_{0}})}\right|\\ \;\;\leq\int_{0}^{t}{\left|{\frac{{\partial{{\bf{z}}_{1}}\left[{t,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}}{\begin{Bmatrix}{\mathbf{0}}_{r\times 1}\\ {{g_{2}}\left[{{{\bf{z}}_{2,1}}(\tau,{{\bf{z}}_{0}}),{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}}),{\bf{u}}(\tau)}\right]-{g_{1}}\left[{{{\bf{z}}_{2,1}}(\tau,{{\bf{z}}_{0}}),{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}}),{\bf{u}}(\tau)}\right]}\end{Bmatrix}}}\right|{\rm{d}}\tau}\\ \;\;\leq\int_{0}^{t}{{\left|{\left\{{\frac{{\partial{{\bf{z}}_{1}}\left[{t,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}}}\right\}^{\top}}\right|}_{2,1}{{\left|{{g_{2}}\left[{{{\bf{z}}_{2,1}}(\tau,{{\bf{z}}_{0}}),{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}}),{\bf{u}}(\tau)}\right]-{g_{1}}\left[{{{\bf{z}}_{2,1}}(\tau,{{\bf{z}}_{0}}),{{\bf{z}}_{2,2}}(\tau,{{\bf{z}}_{0}}),{\bf{u}}(\tau)}\right]}\right|}_{2}}{\rm{d}}\tau}\\ \;\;\leq\mathop{\max}\limits_{\tau\in[0,T]}{\left|{{\bf{\Delta}}g(\tau,{{\bf{z}}_{0}})}\right|_{2}}\int_{0}^{t}{{{\left|\left\{{\frac{{\partial{{\bf{z}}_{1}}\left[{t,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}}}\right\}^{\top}\right|}_{2,1}}{\rm{d}}\tau},\end{array}$ (95)
where
$\begin{array}[]{l}{\left|\left\{{\frac{{\partial{{\bf{z}}_{1}}\left[{t,{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}\right]}}{{\partial{{\bf{z}}_{2}}(\tau,{{\bf{z}}_{0}})}}}\right\}^{\top}\right|_{2,1}}=\sum\limits_{k=1}^{2r}{{{\left|{\frac{{\partial{z_{1,k}}\left({t,{{\bf{z}}_{2}}}\right)}}{{\partial{{\bf{z}}_{2}}}}}\right|}_{2}}}=\sum\limits_{k=1}^{2r}{\left|{\frac{{\partial{z_{1,k}}\left({t,{{\bf{z}}_{2}}}\right)}}{{\partial{{\bf{z}}_{2}}}}\cdot{{\bf{\Delta}}_{k}}}\right|}\\ \;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;=\sum\limits_{k=1}^{2r}{\mathop{\lim}\limits_{v\to 0}\frac{{\left|{{z_{1,k}}\left({t,{{\bf{z}}_{2}}+v{{\bf{\Delta}}_{k}}}\right)-{z_{1,k}}\left({t,{{\bf{z}}_{2}}}\right)}\right|}}{{{{\left|{v{{\bf{\Delta}}_{k}}}\right|}_{2}}}}}\\ \;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\mathop{\leq}\limits^{(a)}\sum\limits_{k=1}^{2r}{\mathop{\lim}\limits_{v\to 0}\frac{{{\beta_{k}}({{\left|{v{{\bf{\Delta}}_{k}}}\right|}_{2}},t-\tau)}}{{{{\left|{v{{\bf{\Delta}}_{k}}}\right|}_{2}}}}}={\left.{\sum\limits_{k=1}^{2r}{\frac{\partial}{{\partial h}}{\beta_{k}}(h,t-\tau)}}\right|_{h={0^{+}}}},\end{array}$ (96)
$z_{1,k}\left(t,{\mathbf{z}}_{2}\right)$ denotes the $k^{\text{th}}$ element of ${\mathbf{z}}_{1}\left(t,{\mathbf{z}}_{2}\right)$, ${\mathbf{\Delta}}_{k}$ is the unit vector in the direction of the partial derivative of $z_{1,k}\left(t,{\mathbf{z}}_{2}\right)$ with respect to ${\mathbf{z}}_{2}$, and Step (a) holds from Eq. (24) for all ${\mathbf{z}}_{2}\in\Omega_{\mathbf{z}}$. Substituting Eq. (96) into Eq. (95), $\left|{\mathbf{z}}_{2}\left(t,{\mathbf{z}}_{0}\right)-{\mathbf{z}}_{1}\left(t,{\mathbf{z}}_{0}\right)\right|$ can be bounded by
$\begin{array}[]{l}\left|{{{\bf{z}}_{2}}(t,{{\bf{z}}_{0}})-{{\bf{z}}_{1}}(t,{{\bf{z}}_{0}})}\right|\leq\mathop{\max}\limits_{\tau\in[0,T]}{\left|{{\bf{\Delta}}g(\tau,{{\bf{z}}_{0}})}\right|_{2}}\int_{0}^{t}{{\left.{\sum\limits_{k=1}^{2r}{\frac{\partial}{{\partial h}}{\beta_{k}}(h,t-\tau)}}\right|_{h={0^{+}}}}{\rm{d}}\tau}\\ \;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\mathop{\leq}\limits^{(a)}\mathop{\max}\limits_{\tau\in[0,T]}{\left|{{\bf{\Delta}}g(\tau,{{\bf{z}}_{0}})}\right|_{2}}{B_{\beta}},\end{array}$ (97)
for all $\mathbf{u}(t)\in K$ and all $t\in[0,T]$, where Step (a) holds from $\left.\partial\beta_{k}(h,t-\tau)/\partial h\right|_{h=0^{+}}\geq 0$ and $B_{\beta}$ is the stability bound in Eq. (25). Eq. (28) is proven by substituting Eq. (90) into Eq. (97). $\blacksquare$
B.3 Proof of Theorem 2
Proof: The difference between ${\mathbf{y}}(t)$ and ${\hat{\mathbf{y}}}(t)$ can be bounded by
$\left|{{\bf{y}}(t)-{\hat{\bf y}}(t)}\right|\leq\left|{h\left[{{\bf{x}}(t)}\right]-h\left[{{\hat{\bf x}}(t)}\right]}\right|+\left|{{\it{\Pi}}\left[{{\bf{x}}(t),{\bf{u}}(0),t}\right]-h\left[{{\bf{x}}(t)}\right]}\right|,$ (98)
where $\mathbf{x}(t)$ is the solution of the second-order ODE in Eq. (1).
First, the first part on the right side of Eq. (98) is considered. According to the assumptions in Theorem 2, since $g(\cdot)$ is Lipschitz continuous over $\mathbb{R}^{2r+p}$, the solutions ${\hat{\mathbf{x}}}(t)$ and ${\hat{\mathbf{x}}}^{\prime}(t)$ to Eq. (29) driven by all $\mathbf{u}(t)\in K$, with the initial conditions ${\hat{\mathbf{x}}}(0)={\hat{\mathbf{x}}}^{\prime}(0)=\mathbf{0}$, are unique and they can be bounded by two constants $B_{\hat{\mathbf{x}}}$ and $B_{\hat{\mathbf{x}}^{\prime}}$, respectively, that is $\left|{\hat{x}_{j}}(t)\right|\leq B_{\hat{\mathbf{x}}}$ and $\left|{\hat{x}_{j}^{\prime}}(t)\right|\leq B_{\hat{\mathbf{x}}^{\prime}}$ for all $\mathbf{u}(t)\in K$, all $t\in[0,T]$, and j=1,2,,rj=1,2,...,r. Since the dynamical system in Eq. (29) is uniformly asymptotically incrementally stable for all $\mathbf{u}(t)\in K$ on a domain $\Omega\left(B_{\mathbf{x}}^{r},B_{{\mathbf{x}}^{\prime}}^{r},B_{K}^{p}\right)=\left[-B_{\mathbf{x}},B_{\mathbf{x}}\right]^{r}\times\left[-B_{{\mathbf{x}}^{\prime}},B_{{\mathbf{x}}^{\prime}}\right]^{r}\times\left[-B_{K},B_{K}\right]^{p}$, where $B_{K}$ is the bound of all $\mathbf{u}(t)\in K$ in Assumption 2, $B_{{\mathbf{x}}}=\alpha B_{\beta{g}}+B_{\hat{\mathbf{x}}}$, $B_{{\mathbf{x}}^{\prime}}=\alpha B_{\beta{g}}+B_{\hat{\mathbf{x}}^{\prime}}$, $\alpha$ is a positive coefficient, and $B_{\beta{g}}$ is the stability bound of the dynamical system in Eq. (29), which is calculated by Eq. (25), if the difference between the MLP ${\it{\Gamma}}(\cdot)$ in Eq. (1) and $g(\cdot)$ in Eq. (29) over $\Omega\left(B_{\mathbf{x}}^{r},B_{{\mathbf{x}}^{\prime}}^{r},B_{K}^{p}\right)$ is bounded by
${\left|{{\it{\Gamma}}({\bf{x}},{\bf{x^{\prime}}},{\bf{u}})-g({\bf{x}},{\bf{x^{\prime}}},{\bf{u}})}\right|_{2}}\leq\alpha,$ (99)
then, according to Eq. (28), ${{\mathbf{x}}}(t)$ and ${{\mathbf{x}}}^{\prime}(t)$ from Eq. (1) and ${\hat{\mathbf{x}}}(t)$ and ${\hat{\mathbf{x}}}^{\prime}(t)$ from Eq. (29), driven by the same arbitrary $\mathbf{u}(t)\in K$, with the same initial conditions ${{\mathbf{x}}}(0)=\mathbf{x}^{\prime}(0)={\hat{\mathbf{x}}}(0)={\hat{\mathbf{x}}}^{\prime}(0)=\mathbf{0}$, satisfy
$\begin{array}[]{l}\left|{{\bf{x}}(t)-{\hat{\bf x}}(t)}\right|+\left|{{\bf{x^{\prime}}}(t)-{\hat{\bf x}^{\prime}}(t)}\right|\leq\mathop{\max}\limits_{\tau\in[0,T]}{\left|{{\it{\Gamma}}\left[{{\bf{x}}(\tau),{\bf{x^{\prime}}}(\tau),{\bf{u}}(\tau)}\right]-g\left[{{\bf{x}}(\tau),{\bf{x^{\prime}}}(\tau),{\bf{u}}(\tau)}\right]}\right|_{2}}{B_{{\beta_{g}}}}\leq\alpha{B_{{\beta_{g}}}},\end{array}$ (100)
and the values of ${{\mathbf{x}}}(t)$ and ${{\mathbf{x}}}^{\prime}(t)$ are contained in $\Omega\left(B_{\mathbf{x}}^{r},B_{{\mathbf{x}}^{\prime}}^{r},B_{K}^{p}\right)$ for all $\mathbf{u}(t)\in K$ and all $t\in[0,T]$.
Since each $g_{j}(\cdot)$ of $g(\cdot)$ has arbitrary finite-order derivatives over $\mathbb{R}^{2r+p}$, there exists a corresponding ${\tilde{g}}_{j}(\cdot)$ equaling $g_{j}(\cdot)$ over $\Omega\left(B_{\mathbf{x}}^{r},B_{{\mathbf{x}}^{\prime}}^{r},B_{K}^{p}\right)$ and the first moment of the Fourier transform $\mathcal{H}_{{\tilde{g}}_{j}}\left(\boldsymbol{\upomega}\right)$ of ${\tilde{g}}_{j}(\cdot)$ is finite (Barron, 1993), that is
$\int_{{\mathbb{R}^{2r+p}}}{\left|\boldsymbol{\upomega}\right|_{2}}\left|{{\mathcal{H}_{{{\tilde{g}}_{j}}}}(\boldsymbol{\upomega})}\right|{\rm{d}}\boldsymbol{\upomega}<+\infty,\;\;j=1,2,\ldots,r,$ (101)
where $\boldsymbol{\upomega}=\left[\omega_{1},\omega_{2},...,\omega_{2r+p}\right]^{\top}$. Based on Theorem 2.2 in Yukich et al. (1995) and the fact that four ReLU activation functions can form a total variation activation function that has a bounded support, e.g., $\left[\sigma_{\mathrm{ReLU}}(x)-\sigma_{\mathrm{ReLU}}(x-1)\right]-\left[\sigma_{\mathrm{ReLU}}(x-2)-\sigma_{\mathrm{ReLU}}(x-3)\right]$, for an arbitrary $\varepsilon_{1}$ satisfying $0<\varepsilon_{1}<\alpha/r$, there exist $r$ one-hidden-layer MLPs ${\it{\Gamma}}_{j}(\cdot)$ with $\sigma_{\mathrm{ReLU}}(\cdot)$, j=1,2,,rj=1,2,...,r, the widths of whose input and output layers are respectively $2r+p$ and 1, and the width of the hidden layer for each ${\it{\Gamma}}_{j}(\cdot)$ is at most
${w_{{{\it{\Gamma}}_{j}}}}\leq 8\left\lceil{C_{{{\it{\Gamma}}_{j}}}^{2}\varepsilon_{1}^{-2}}\right\rceil,$ (102)
such that the difference between ${\it{\Gamma}}_{j}(\cdot)$ and $g_{j}(\cdot)$ over $\Omega\left(B_{\mathbf{x}}^{r},B_{{\mathbf{x}}^{\prime}}^{r},B_{K}^{p}\right)$ is bounded by
$\left|{{{\it{\Gamma}}_{j}}({\bf{x}},{\bf{x^{\prime}}},{\bf{u}})-{g_{j}}({\bf{x}},{\bf{x^{\prime}}},{\bf{u}})}\right|\leq{\varepsilon_{1}}.$ (103)
In Eq. (102), $C_{{\it{\Gamma}}_{j}}=C_{{\it{\Gamma}}_{j}}\left[\mathcal{H}_{{\tilde{g}}_{j}}\left(\boldsymbol{\upomega}\right),\sigma_{\mathrm{ReLU}}(\cdot),2r+p,\Omega\left(B_{\mathbf{x}}^{r},B_{{\mathbf{x}}^{\prime}}^{r},B_{K}^{p}\right)\right]$ is a constant only depending on the Fourier transform $\mathcal{H}_{{\tilde{g}}_{j}}\left(\boldsymbol{\upomega}\right)$ of ${\tilde{g}}_{j}(\cdot)$ corresponding to $g_{j}(\cdot)$, the domain $\Omega\left(B_{\mathbf{x}}^{r},B_{{\mathbf{x}}^{\prime}}^{r},B_{K}^{p}\right)$, the activation function $\sigma_{\mathrm{ReLU}}(\cdot)$, and the number $2r+p$ of the arguments of $g_{j}(\cdot)$.
Then, from Eqs. (99), (100), and (102), for an arbitrary $\varepsilon_{1}$ satisfying $0<\varepsilon_{1}<\alpha/r$, there exists a one-hidden-layer MLP ${\it{\Gamma}}(\cdot)$ consisting of above ${\it{\Gamma}}_{1}(\cdot),{\it{\Gamma}}_{2}(\cdot),...,{\it{\Gamma}}_{r}(\cdot)$, the widths of the input and output layers of ${\it{\Gamma}}(\cdot)$ are respectively $2r+p$ and $r$, and the width of its hidden layer is at most
${w_{{{\it{\Gamma}}}}}\leq 8r\left\lceil{C_{{{\it{\Gamma}}}}^{2}\varepsilon_{1}^{-2}}\right\rceil,$ (104)
where ${C_{\it{\Gamma}}}=\max\left({{C_{{{\it{\Gamma}}_{1}}}},{C_{{{\it{\Gamma}}_{2}}}},\ldots,{C_{{{\it{\Gamma}}_{r}}}}}\right)$, such that for all $\mathbf{u}(t)\in K$, the corresponding solutions $\mathbf{x}(t)$ from Eq. (1) and ${\hat{\mathbf{x}}}(t)$ from Eq. (29), with the initial conditions $\mathbf{x}(0)={\mathbf{x}}^{\prime}(0)={\hat{\mathbf{x}}}(0)={{\hat{\mathbf{x}}}}^{\prime}(0)=\mathbf{0}$, satisfy
$\begin{array}[]{l}\left|{{\bf{x}}(t)-{\hat{\bf x}}(t)}\right|\leq\mathop{\max}\limits_{\tau\in[0,T]}{\left|{{\it{\Gamma}}\left[{{\bf{x}}(\tau),{\bf{x}}(\tau),{\bf{u}}(\tau)}\right]-g\left[{{\bf{x}}(\tau),{\bf{x}}(\tau),{\bf{u}}(\tau)}\right]}\right|_{2}}{B_{{\beta_{g}}}}\\ \;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\leq{B_{{\beta_{g}}}}\sqrt{\sum\limits_{j=1}^{r}{\mathop{\max}\limits_{\tau\in[0,T]}\left\{{{{\it{\Gamma}}_{j}}\left[{{\bf{x}}(\tau),{\bf{x}}(\tau),{\bf{u}}(\tau)}\right]-{g_{j}}\left[{{\bf{x}}(\tau),{\bf{x}}(\tau),{\bf{u}}(\tau)}\right]}\right\}^{2}}}\leq{B_{{\beta_{g}}}}\sqrt{r}{\varepsilon_{1}}\end{array}$ (105)
for all $t\in[0,T]$, where Eq. (99) is satisfied by
$\begin{array}[]{l}\mathop{\max}\limits_{\tau\in[0,T]}{\left|{{\it{\Gamma}}\left[{{\bf{x}}(\tau),{\bf{x}}(\tau),{\bf{u}}(\tau)}\right]-g\left[{{\bf{x}}(\tau),{\bf{x}}(\tau),{\bf{u}}(\tau)}\right]}\right|_{2}}\\ \;\;\;\;\leq\sqrt{\sum\limits_{j=1}^{r}{\mathop{\max}\limits_{\tau\in[0,T]}\left\{{{{\it{\Gamma}}_{j}}\left[{{\bf{x}}(\tau),{\bf{x}}(\tau),{\bf{u}}(\tau)}\right]-{g_{j}}\left[{{\bf{x}}(\tau),{\bf{x}}(\tau),{\bf{u}}(\tau)}\right]}\right\}^{2}}}\leq\sqrt{r}{\varepsilon_{1}}\leq\alpha{r^{-0.5}}\leq\alpha.\end{array}$ (106)
The first term on the right side of Eq. (98) can be bounded by
$\left|{h\left[{{\hat{\bf x}}(t)}\right]-h\left[{{\bf{x}}(t)}\right]}\right|\leq{L_{h}}\left|{{\hat{\bf x}}(t)-{\bf{x}}(t)}\right|\leq{L_{h}}{B_{{\beta_{g}}}}\sqrt{r}{\varepsilon_{1}}$ (107)
for all $t\in[0,T]$, where $L_{h}$ is the Lipschitz constant of the function $h(\cdot)$.
Subsequently, the second term on the right side of Eq. (98) is taken into account. Using the same method based on the Barron class (Barron, 1993), since $h(\cdot)$ has arbitrary finite-order derivatives over $\mathbb{R}^{r}$, for each $j^{\text{th}}$ element $h_{j}(\cdot)$ of $h(\cdot)$, j=1,2,,qj=1,2,...,q, there exists a corresponding ${\tilde{h}}_{j}(\cdot)$ equaling $h_{j}(\cdot)$ over $\Omega\left(B_{\mathbf{x}}^{r}\right)=\left[-B_{\mathbf{x}},B_{\mathbf{x}}\right]^{r}$, the first moment of the Fourier transform $\mathcal{H}_{{\tilde{h}}_{j}}\left(\boldsymbol{\upomega}\right)$ of ${\tilde{h}}_{j}(\cdot)$ is finite (Barron, 1993). Based on Theorem 2.2 in Yukich et al. (1995) and the fact that four ReLU activation functions can form a total variation activation function that has a bounded support, for an arbitrary positive $\varepsilon_{2}$, there exists q one-hidden-layer MLPs ${\it{\Pi}}_{j}(\cdot)$ with $\sigma_{\mathrm{ReLU}}(\cdot)$, the parameters in each ${\it{\Pi}}_{j}(\cdot)$ with respect to $\mathbf{u}(0)$ and $t$ are zero, the widths of the input and output layers of each ${\it{\Pi}}_{j}(\cdot)$ are respectively $r+p+1$ and 1, the width of the hidden layer for each ${\it{\Pi}}_{j}(\cdot)$ is at most
${w_{{{\it{\Pi}}_{j}}}}\leq 8\left\lceil{C_{{{\it{\Pi}}_{j}}}^{2}\varepsilon_{2}^{-2}}\right\rceil,$ (108)
such that the difference between ${\it{\Pi}}_{j}\left[\mathbf{x},\mathbf{u}(0),t\right]$ and $h_{j}\left(\mathbf{x}\right)$ over $\mathbf{x}\in\Omega\left(B_{\mathbf{x}}^{r}\right)$ is bounded by
$\left|{{{\it{\Pi}}_{j}}\left[{{\bf{x}},{\bf{u}}(0),t}\right]-{h_{j}}({\bf{x}})}\right|\leq{\varepsilon_{2}}.$ (109)
In Eq. (108), $C_{{{\it{\Pi}}_{j}}}=\left[\mathcal{H}_{{\tilde{h}}_{j}}\left(\boldsymbol{\upomega}\right),\sigma_{\mathrm{ReLU}}(\cdot),r,\Omega\left(B_{\mathbf{x}}^{r}\right)\right]$ is a constant only depending on the Fourier transform $\mathcal{H}_{{\tilde{h}}_{j}}\left(\boldsymbol{\upomega}\right)$ of ${\tilde{h}}_{j}(\cdot)$ corresponding to ${{h}}_{j}(\cdot)$, the domain $\Omega\left(B_{\mathbf{x}}^{r}\right)$, the activation function $\sigma_{\mathrm{ReLU}}(\cdot)$, and the number r.
Then, for an arbitrary positive $\varepsilon_{2}$, there exists a one-hidden-layer MLP ${\it{\Pi}}(\cdot)$ consisting of above q MLPs, ${\it{\Pi}}_{1}(\cdot),{\it{\Pi}}_{2}(\cdot)$ $\dots,{\it{\Pi}}_{q}(\cdot),$ the widths of the input and output layers of ${\it{\Pi}}(\cdot)$ are respectively $r+p+1$ and $q$, and the width of its hidden layer is at most
${w_{{{\it{\Pi}}}}}\leq 8q\left\lceil{C_{{{\it{\Pi}}}}^{2}\varepsilon_{2}^{-2}}\right\rceil,$ (110)
where ${C_{\it{\Pi}}}=\max\left({{C_{{{\it{\Pi}}_{1}}}},{C_{{{\it{\Pi}}_{2}}}},\ldots,{C_{{{\it{\Pi}}_{q}}}}}\right)$, such that for all $\mathbf{u}(t)\in K$, the corresponding solutions $\mathbf{x}(t)$ from Eq. (1), whose ${\it{\Gamma}}(\cdot)$ satisfying Eqs. (104) to (106), with the initial conditions $\mathbf{x}(0)={\mathbf{x}}^{\prime}(0)=\mathbf{0}$, satisfy
$\left|{{\it{\Pi}}\left[{{\bf{x}}(t),{\bf{u}}(0),t}\right]-h\left[{{\bf{x}}(t)}\right]}\right|\leq q{\varepsilon_{2}}$ (111)
for all $t\in[0,T]$. Substituting Eqs. (107) and (111) into Eq. (98), Eq. (30) is proven. $\blacksquare$
Appendix C Explanation of the time dependence of $\Psi\left({\mathbf{x}},{\mathbf{0}},t\right)$
Assume that $\mathbf{u}(0)=\mathbf{0}$ holds for all $\mathbf{u}(t)\in K$, then, the continuous mapping $\Psi\left({\bf{x}},{\bf{0}},t\right)$ in Lemma 3 is now defined as $\Psi\left({\mathbf{x}},{\mathbf{0}},t\right)=\Psi\left[{\mathbf{x}}(t),{\bf{0}},t\right]=\Phi[{\hat{\mathbf{u}}}_{t}(\tau)](t)$, where the approximate ${\hat{\mathbf{u}}}_{t}(\tau)$ of $\mathbf{u}(t)$ is calculated as $\hat{\mathbf{u}}_{t}(\tau)=\sum_{n=1}^{M}{{\alpha_{n}}{{\bf{x}}_{n}}(t)\sin\left[{{\omega_{n}}(t-\tau)-{\theta_{n}}}\right]}$ and $\mathbf{x}_{n}(t)=\mathcal{L}_{t}\mathbf{u}(\omega_{n})$ in Eq. (55). For $\forall\mathbf{u}_{1}(t)\in K$ with a positive $\Delta t$, its shifted version $\mathbf{u}_{2}(t)$ is
$\mathbf{u}_{2}(t)=\left\{\begin{aligned} &\mathbf{u}_{1}(t-\Delta t),\;\;\;t\geq\Delta t\\ &\mathbf{0},\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\text{otherwise}\end{aligned}.\right$. (112)
For $\forall t_{1}\geq 0$ and $t_{2}=t_{1}+\Delta t\leq T$, it is satisfied $\mathbf{x}_{1,n}(t_{1})=\mathcal{L}_{t_{1}}\mathbf{u}_{1}(\omega_{n})=\mathcal{L}_{t_{2}}\mathbf{u}_{2}(\omega_{n})=\mathbf{x}_{2,n}(t_{2})$, which is denoted as $\mathbf{x}_{1,n}(t_{1})=\mathbf{x}_{2,n}(t_{2})=\mathbf{x}_{n}$. Then, the approximates ${\hat{\mathbf{u}}_{1,t_{1}}}(\tau)$ and ${\hat{\mathbf{u}}_{2,t_{2}}}(\tau)$ of $\mathbf{u}_{1}(t)$ and $\mathbf{u}_{2}(t)$ are respectively ${\hat{\mathbf{u}}_{1,t_{1}}}(\tau)=\sum_{n=1}^{M}{{\alpha_{n}}{{\mathbf{x}}_{n}}\sin\left[{{\omega_{n}}(t_{1}-\tau)-{\theta_{n}}}\right]}$ and ${\hat{\mathbf{u}}_{2,t_{2}}}(\tau)=\sum_{n=1}^{M}{{\alpha_{n}}{{\bf{x}}_{n}}\sin\left[{{\omega_{n}}(t_{2}-\tau)-{\theta_{n}}}\right]}$. The related $\Psi\left({\mathbf{x}}_{n},{\mathbf{0}},t_{1}\right)$ is $\Psi\left({\mathbf{x}}_{n},{\mathbf{0}},t_{1}\right)$ $\Phi\left[\sum_{n=1}^{M}{{\alpha_{n}}{{\mathbf{x}}_{n}}\sin\left[{{\omega_{n}}(t_{1}-\tau)-{\theta_{n}}}\right]}\right](t_{1})$ and $\Psi\left({\mathbf{x}}_{n},{\mathbf{0}},t_{2}\right)$ is $\Psi\left({\mathbf{x}}_{n},{\mathbf{0}},t_{2}\right)=\Phi\left[\sum_{n=1}^{M}{{\alpha_{n}}{{\mathbf{x}}_{n}}\sin\left[{{\omega_{n}}(t_{2}-\tau)-{\theta_{n}}}\right]}\right](t_{2})$. Although $\Phi$ is assumed to be time-autonomous, thereby $\Phi\left[\mathbf{u}_{1}(\tau)\right](t_{1})=\Phi\left[\mathbf{u}_{2}(\tau)\right](t_{2})$, $\Psi({\mathbf{x}}_{n},{\mathbf{0}},$ $t_{1})$ does not equal $\Psi\left({\mathbf{x}}_{n},{\mathbf{0}},t_{2}\right)$. This is because the valid range of $\tau$ in $\sin\left[{{\omega_{n}}(t_{1}-\tau)-{\theta_{n}}}\right]$ is $0,t_{1}]$ but that of $\sin\left[{{\omega_{n}}(t_{2}-\tau)-{\theta_{n}}}\right]$ is $0,t_{2}]$. Thus, the time dependence of $\Psi\left({\mathbf{x}},{\mathbf{0}},t\right)$ cannot be ignored even if $\Phi$ is time-autonomous.
An alternative approximate ${\hat{\mathbf{u}}_{2,t_{2},\Delta t}}(\tau)$ of $\mathbf{u}_{2}(t)$ is
${\hat{\mathbf{u}}_{2,t_{2},\Delta t}}(\tau)=\left\{\begin{aligned} &\sum\limits_{n=1}^{M}{{\alpha_{n}}{{\bf{x}}_{n}}\sin\left[{{\omega_{n}}(t_{2}-\tau)-{\theta_{n}}}\right]},\;\;\;\tau\in[\Delta t,t_{2}]\;\text{and}\;\Delta t\leq t_{2}\leq T\\ &\mathbf{0},\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\tau\in\left[0,\Delta t\right)\end{aligned}.\right$. (113)
It is satisfied $\Phi\left[\hat{\mathbf{u}}_{1,t_{1}}(\tau)\right](t_{1})=\Phi\left[\hat{\mathbf{u}}_{2,t_{2},\Delta t}(\tau)\right](t_{2})$. However, the approximate in Eq. (113) cannot apply to the input functions with non-zero values over $0,\Delta t)$, and thus it is not applicable to all functions in K.
Appendix D Proof of Eq. (27) with the Grönwall’s inequality
Proof: $\left|{\mathbf{z}}_{2}\left(t,{\mathbf{z}}_{0}\right)-{\mathbf{z}}_{1}\left(t,{\mathbf{z}}_{0}\right)\right|$, which are driven by the same $\mathbf{u}(t)\in K$ with the same initial condition ${\mathbf{z}}_{0}$, can be bounded by
$\begin{array}[]{l}\left|{\mathbf{z}}_{2}\left(t,{\mathbf{z}}_{0}\right)-{\mathbf{z}}_{1}\left(t,{\mathbf{z}}_{0}\right)\right|\\ \;\;\;=\left|{\int_{0}^{t}{{{{\bf{z}}}_{2}^{\prime}}(\tau){\rm{d}}\tau}-\int_{0}^{t}{{{{\bf{z}}}_{1}^{\prime}}(\tau){\rm{d}}\tau}}\right|\\ \;\;\;=\left|{\int_{0}^{t}{{{\bf{z}}_{2,2}}(\tau){\rm{d}}\tau}-\int_{0}^{t}{{{\bf{z}}_{1,2}}(\tau){\rm{d}}\tau}}\right|+\left|{\int_{0}^{t}{{g_{1}}\left[{{{\bf{z}}_{1,1}}(\tau),{{\bf{z}}_{1,2}}(\tau),{\bf{u}}(\tau)}\right]}{\rm{d}}\tau-\int_{0}^{t}{{g_{2}}\left[{{{\bf{z}}_{2,1}}(\tau),{{\bf{z}}_{2,2}}(\tau),{\bf{u}}(\tau)}\right]{\rm{d}}\tau}}\right|\\ \;\;\;\leq\int_{0}^{t}{\left|{{{\bf{z}}_{2,2}}(\tau)-{{\bf{z}}_{1,2}}(\tau)}\right|{\rm{d}}\tau}+\int_{0}^{t}{\left|{{g_{1}}\left[{{{\bf{z}}_{1,1}}(\tau),{{\bf{z}}_{1,2}}(\tau),{\bf{u}}(\tau)}\right]-{g_{2}}\left[{{{\bf{z}}_{2,1}}(\tau),{{\bf{z}}_{2,2}}(\tau),{\bf{u}}(\tau)}\right]}\right|}{\rm{d}}\tau\\ \;\;\;\leq\int_{0}^{t}{\left|{{{\bf{z}}_{2,2}}(\tau)-{{\bf{z}}_{1,2}}(\tau)}\right|{\rm{d}}\tau}+\int_{0}^{t}{\left|{{g_{1}}\left[{{{\bf{z}}_{1,1}}(\tau),{{\bf{z}}_{1,2}}(\tau),{\bf{u}}(\tau)}\right]-{g_{1}}\left[{{{\bf{z}}_{2,1}}(\tau),{{\bf{z}}_{1,2}}(\tau),{\bf{u}}(\tau)}\right]}\right|}{\rm{d}}\tau\\ \;\;\;\;\;\;\;+\int_{0}^{t}{\left|{{g_{1}}\left[{{{\bf{z}}_{2,1}}(\tau),{{\bf{z}}_{1,2}}(\tau),{\bf{u}}(\tau)}\right]-{g_{1}}\left[{{{\bf{z}}_{2,1}}(\tau),{{\bf{z}}_{2,2}}(\tau),{\bf{u}}(\tau)}\right]}\right|}{\rm{d}}\tau\\ \;\;\;\;\;\;\;+\int_{0}^{t}{\left|{{g_{1}}\left[{{{\bf{z}}_{2,1}}(\tau),{{\bf{z}}_{2,2}}(\tau),{\bf{u}}(\tau)}\right]-{g_{2}}\left[{{{\bf{z}}_{2,1}}(\tau),{{\bf{z}}_{2,2}}(\tau),{\bf{u}}(\tau)}\right]}\right|}{\rm{d}}\tau.\end{array}$ (114)
Since $g_{1}(\cdot)$ is Lipschitz continuous, $\left|{\mathbf{z}}_{2}\left(t,{\mathbf{z}}_{0}\right)-{\mathbf{z}}_{1}\left(t,{\mathbf{z}}_{0}\right)\right|$ can be bounded by
$\begin{array}[]{l}\left|{\mathbf{z}}_{2}\left(t,{\mathbf{z}}_{0}\right)-{\mathbf{z}}_{1}\left(t,{\mathbf{z}}_{0}\right)\right|\\ \;\;\;\leq\int_{0}^{t}{\left|{{{\bf{z}}_{2,2}}(\tau)-{{\bf{z}}_{1,2}}(\tau)}\right|{\rm{d}}\tau}+{L_{{g_{1}},1}}\int_{0}^{t}{\left|{{{\bf{z}}_{2,1}}(\tau)-{{\bf{z}}_{1,1}}(\tau)}\right|}{\rm{d}}\tau+{L_{{g_{1}},2}}\int_{0}^{t}{\left|{{{\bf{z}}_{1,2}}(\tau)-{{\bf{z}}_{2,2}}(\tau)}\right|}{\rm{d}}\tau\\ \;\;\;\;\;\;\;+\int_{0}^{t}{\left|{{g_{1}}\left[{{{\bf{z}}_{2,1}}(\tau),{{\bf{z}}_{2,2}}(\tau),{\bf{u}}(\tau)}\right]-{g_{2}}\left[{{{\bf{z}}_{2,1}}(\tau),{{\bf{z}}_{2,2}}(\tau),{\bf{u}}(\tau)}\right]}\right|}{\rm{d}}\tau\\ \;\;\;\leq{L_{{g_{1}}}}\int_{0}^{t}{\left|{{{\bf{z}}_{1}}(\tau)-{{\bf{z}}_{2}}(\tau)}\right|}{\rm{d}}\tau+\mathop{\max}\limits_{\tau\in[0,T]}\left|{{g_{1}}\left[{{{\bf{z}}_{2,1}}(\tau),{{\bf{z}}_{2,2}}(\tau),{\bf{u}}(\tau)}\right]-{g_{2}}\left[{{{\bf{z}}_{2,1}}(\tau),{{\bf{z}}_{2,2}}(\tau),{\bf{u}}(\tau)}\right]}\right|T,\\ \;\;\;\mathop{\leq}\limits^{(a)}{e^{T{L_{{g_{1}}}}}}T\mathop{\max}\limits_{\tau\in[0,T]}\left|{{g_{1}}\left[{{{\bf{z}}_{2,1}}(\tau),{{\bf{z}}_{2,2}}(\tau),{\bf{u}}(\tau)}\right]-{g_{2}}\left[{{{\bf{z}}_{2,1}}(\tau),{{\bf{z}}_{2,2}}(\tau),{\bf{u}}(\tau)}\right]}\right|,\end{array}$ (115)
where $L_{g_{1}}=\max\left[L_{g_{1},1},\left(L_{g_{1},2}+1\right)\right]$, $L_{g_{1},1}$ and $L_{g_{1},2}$ are respectively the two Lipschitz constants of $g_{1}(\cdot)$ with respect to its first and second arguments, and Step (a) holds from the Grönwall’s inequality (Ames and Pachpatte, 1997). Eq. (115) is the same as Eq. (27).
Appendix E Approach for Training the Neural Oscillator
By introducing the state space vector $\mathbf{z}(t)=\left[\mathbf{z}_{1}^{\top}(t),\mathbf{z}_{2}^{\top}(t)\right]^{\top}=\left[\mathbf{x}^{\top}(t),\mathbf{x}^{\prime\top}(t)\right]^{\top}$, Eq. (1) over the time interval $\left[0,T\right]$ can be rewritten in the state function form
$\left\{\begin{aligned} &\mathbf{z}_{1}^{\prime}(t)=\mathbf{z}_{2}(t)\\ &\mathbf{z}_{2}^{\prime}(t)={\it{\Gamma}}\left[\mathbf{z}_{1}(t),\mathbf{z}_{2}(t),\mathbf{u}(t)\right]\\ &\mathbf{y}(t)={\it{\Pi}}\left[\mathbf{z}_{1}(t),\mathbf{u}(0),t\right]\end{aligned}.\right$. (116)
Utilizing a second-order Runge-Kutta scheme (Griffiths and Higham, 2010), a time-discretization form of Eq. (116) is
$\left\{\begin{aligned} &\mathbf{y}(t_{i+1})={\it{\Pi}}\left[\mathbf{z}_{1}(t_{i+1}),\mathbf{u}(0),t_{i+1}\right]\\ &\mathbf{z}(t_{i+1})=0.5\Delta{t_{i}}\left(\mathbf{k}_{2}+\mathbf{k}_{1}\right)\\ &\mathbf{k}_{2}=\begin{Bmatrix}\mathbf{z}_{2}(t_{i})+\Delta{t_{i}}\mathbf{k}_{12}\\ {\it{\Gamma}}\left[\mathbf{z}_{1}(t_{i})+\Delta{t_{i}}\mathbf{k}_{11},\mathbf{z}_{2}(t_{i})+\Delta{t_{i}}\mathbf{k}_{12},\mathbf{u}(t_{i+1})\right]\end{Bmatrix}\\ &\mathbf{k}_{1}=\begin{bmatrix}\mathbf{k}_{11}\\ \mathbf{k}_{12}\end{bmatrix}=\begin{Bmatrix}\mathbf{z}_{2}(t_{i})\\ {\it{\Gamma}}\left[\mathbf{z}_{1}(t_{i}),\mathbf{z}_{2}(t_{i}),\mathbf{u}(t_{i})\right]\end{Bmatrix}\end{aligned},\right$. (117)
where $\mathbf{z}(0)=\mathbf{0}$, $\mathbf{y}(0)={\it{\Pi}}\left[\mathbf{z}_{1}(0),\mathbf{u}(0),t\right]$, $\Delta{t_{i}}=t_{i+1}-t_{i}$, $0=t_{0}<t_{1}<t_{2}<\dots<t_{I}=T$. Given available output function samples $\hat{\mathbf{y}}_{l}(t_{i})=\left[\hat{y}_{l,1}(t_{i}),\hat{y}_{l,2}(t_{i}),\dots,\hat{y}_{l,q}\right]^{\top}$ caused by the input samples $\mathbf{u}_{l}(t_{i})$, $l=1,2,\dots,L$, the $r$-th power loss function $\ell_{r}$ is used to learn the trainable parameters of the MLPs ${\it{\Gamma}}(\cdot)$ and ${\it{\Pi}}(\cdot)$ in Eq. (116),
$\ell_{r}=\frac{1}{LIq}{\sum\limits_{l=1}^{L}{\sum\limits_{i=0}^{I}\sum\limits_{j=1}^{q}{{\left|{y_{l,j}}(t_{i})-\hat{y}_{l,j}(t_{i})\right|}^{r}}}},$ (118)
where $\mathbf{y}_{l}(t_{i})=\left[y_{l,1}(t_{i}),y_{l,2}(t_{i}),\dots,y_{l,q}(t_{i})\right]^{\top}$ represents the output computed from the neural oscillator driven by $\mathbf{u}_{l}(t_{i})$ using Eq. (117). In the numerical cases of this study, the Adam method (Kingma and Ba, 2014) implemented using the PyTorch (Paszke et al., 2019) is utilized to train the neural oscillator using the Runge-Kutta scheme in Eq. (117) and the loss function $\ell_{r}$ in Eq. (118).